### v1

# What’s the Weight?

Abstract
A basic descriptive question in statistics often asks whether there are differences in mean outcomes between groups based on levels of a discrete covariate (e.g., racial disparities in health outcomes). However, when this categorical covariate of interest is correlated with other factors related to the outcome, direct comparisons may lead to biased estimates and invalid inferential conclusions without appropriate adjustment. Propensity score methods are broadly employed with observational data as a tool to achieve covariate balance, but how to implement them in complex surveys is less studied – in particular, when the survey weights depend on the group variable under comparison. In this work, we focus on a specific example when sample selection depends on race. We propose identification formulas to properly estimate the average controlled difference (ACD) in outcomes between Black and White individuals, with appropriate weighting for covariate imbalance across the two racial groups and generalizability. Via extensive simulation, we show that our proposed methods outperform traditional analytic approaches in terms of bias, mean squared error, and coverage when estimating the ACD for our setting of interest. We are motivated by the interplay between race and social determinants of health when estimating racial differences in telomere length using data from the National Health and Nutrition Examination Survey (NHANES). The NHANES sampling scheme depends on self-reported race. We build a propensity for race to properly adjust for other social determinants while characterizing the controlled effect of race on telomere length. We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for confounding by socioeconomic factors and after utilizing appropriate propensity score and survey weighting techniques. Software to implement these methods and code to reproduce our results can be found in the R package svycdiff at https://github.com/salernos/svycdiff.
Keywords Controlled Outcome Differences  $\cdot$ Propensity Scores  $\cdot$ Complex Surveys  $\cdot$ NHANES  $\cdot$ Racial Disparities
1 Introduction
There is a deep and broad literature in the United States (US) documenting racial disparities along multiple socioeconomic axes (e.g., income, education; see Needham et al., 2013; Rehkopf and Needham, 2019; Zota and VanNoy, 2021). These dimensions are, of course, not independent, making it difficult to separate the variation in outcomes associated with race from differences attributable to other sources. Self-identified race, as measured in many studies, is a complex composite of biology and lived experiences, making it impossible to disentangle the two without additional data (e.g., genetic ancestry). Propensity scores are commonly used in causal inference, but they are also applicable under weaker assumptions to study such differences among groups by balancing complex correlates across these groups (Stuart et al., 2011; Lim et al., 2018; Zeng et al., 2021). While previous work has shown that propensity scores can be used to study racial disparities (see Li and Li, 2023), a gap remains in determining the ideal practice when the sampling design is complex. Specifically, surveys are the backbone of such observational research, but, especially as response rates decline and costs rise, they almost always involve complex designs and non-response patterns. Of particular interest is when the design probability of selection into the survey sample depends on the comparison variable of interest, such as race. In more traditional causal inference terminology, these sensitive variables may be analogous to ‘treatment’ or ‘exposure’ variables. However, unlike in randomized trials, which are often viewed as the gold standard to estimate controlled outcome differences (Zabor et al., 2020), certain observational scenarios preclude randomizing individuals to a particular experimental condition due to moral or biological restrictions. These include practical examples such as the timing of transplant waitlisting or the inability to ‘assign’ race to an individual (Peng et al., 2017; Li and Li, 2023). Our paper addresses this knowledge gap by proposing a strategy for incorporating survey weights from complex survey designs with propensity score methods to achieve covariate balance among heterogeneous groups, such as race.
Given the complex relationship between race, socioeconomic status (SES), and the porous collection of such data in the United States, the ability to make definitive statements on race differences in health outcomes is immensely challenging (Li and Li, 2023). Observational data can offer a wealth of information, however, many observational studies have their limitations in terms of confounding, covariate imbalance, and a lack of representation for generalizing results (Imai and Ratkovic, 2013; Pattanayak et al., 2011). Nationally representative surveys provide the opportunity to study individuals from a range of socioeconomic and cultural backgrounds using techniques such as stratified and clustered sampling methods, however researchers must consider the additional challenge of generalizing results outside the context of the sample. Such generalizations of group differences falter unless we adjust our comparison groups so that we have have similar levels of other covariates that are correlated with the comparison group and outcome of interest. This adjustment, which we term covariate balance can be accomplished a number of ways. In this paper, we use propensity scores. Rosenbaum and Rubin (1983) originally proposed propensity scores as a method of contrasting the conditional probability of experimental group assignment while adjusting for baseline covariates (Rosenbaum and Rubin, 1983). In a typical causal framework, the ability to reduce biases due to confounding by balancing covariate differences is a task that relies on several assumptions, namely (1) strong exchangeability of the treatment/exposure, (2) positivity of the treatment/exposure, and (3) stable unit treatment value. When transporting to a different sample or generalizing to a larger target population, we also require (4) strong selection exchangeability and (5) selection positivity. Rather than interpreting the difference in the observed outcomes as a causal effect, we seek to estimate a controlled descriptive comparison to adjust for differences in covariates between groups. Thus, we leverage propensity scores to compute what Li et al. (2013) call an average controlled difference (ACD). This quantity accounts for associations between variables of interest under weaker assumptions – positivity (1) and (5) and weak selection exchangeability (4) – but does not render a causal interpretation to the estimated group differences. While we later show that our proposed methods also target a causal quantity such as the population average treatment effect (PATE) under stronger assumptions, we do not require this for validity of our proposed estimation procedure, as one or more of these assumptions may be violated in observational data settings – particularly when measuring racial differences in health outcomes. We will later present an auxiliary example where we illustrate PATE estimation for an environmental exposure.
To further complicate things in the study of health disparities, it is often the case that the population characteristics of interest, such as race, are factors that influence the probability of being selected into the sample. For instance, the sampling mechanism for the National Health and Nutrition Examination Survey (NHANES) is partially determined by race and indicators of socioeconomic status. For example, NHANES oversamples non-Hispanic Black individuals and those individuals at or below 130% of the poverty level (Curtin et al., 2012). Previous studies have discussed using propensity score methods for analyzing complex survey data (Austin and Stuart, 2015; Lee et al., 2011; DuGoff et al., 2014; Fuentes et al., 2022), however, not when sample selection depends on race (or other variables commonly used for comparing outcome differences). For example, Rossen and Talih (2014) used propensity scores to assess disparities in weight among US children and adolescents from 2001 to 2010 with data from NHANES. However, there was not consensus on how to appropriately incorporate survey weights at the propensity score estimation stage.
There are several challenges in the problem we consider that are distinct from a traditional PATE estimation problem. A key assumption and function of the propensity score, traditionally, is conditional independence of the outcome and comparison group of interest conditional on the propensity score or any balancing score. Due to the complex nature of these relationships, it is not immediately clear which variables should be conditioned on in the propensity score versus the downstream outcome model. Second, it is often useful to express the research question and corresponding dependencies in terms of conceptual diagrams or directed acyclic graphs (such as Figure 1). In estimating the ACD, we include diagrams to demonstrate associations between the variables of interest. However, given the nature of the variables considered, we caution against interpreting these relationships as causal. If considering these diagrams broadly and even causally, in some instances it would be possible to open backdoor pathways when conditioning on selection into the survey, allowing for the possibility of violating assumptions that causal conclusions rely on. Addressing this in practice often relies on multi-stage modeling techniques such as inverse probability weighting (IPW) or direct standardization via ‘g-methods’ (see Richardson and Rotnitzky, 2014, for a more detailed history).
In this paper, we introduce a strategy for incorporating sampling mechanisms, such as survey weights, and propensity scores to make statements on the controlled difference in outcomes when an individual’s probability of selection depends on the group membership variable of interest (herein, race; see Figure 1). We explore how the proposed methods compare to current analytic strategies. Our motivation comes from a research question addressing the relationship between race, SES, and telomere length, a biomarker for health measured in a particular nationally representative survey (NHANES). The remainder of this manuscript is structured as follows. Section 2 reviews related work in this area. Section 3 introduces and justifies our proposed methods for estimating the quantity of interest, which we will refer to as the ACD in the spirit of Li and Li (2023), and discusses differences with other common approaches. Section 4 provides simulation studies to support our recommendations. Section 5 introduces the data motivating this work and provides results for this example. We close with a summary discussion of the concepts outlined here and some final recommendations to researchers who may be working under similar considerations.
Figure 1: Conceptual diagrams. (A) Assumed variable relationships: $S$ = sample selection indicator; $A$ = group membership variable of interest; $X$ = covariates; $Y$ = observed outcome. (B) Conceptual diagram for our motivating data example: SES = Socioeconomic Status. (C) Directed acyclic graph for our auxiliary data example.
2 Related Work
To our knowledge, Zanutto (2006) and DuGoff et al. (2014) were among the first to discuss propensity score methods as an important tool for addressing confounding in observational studies that may include survey weights. Both authors suggest that, intuitively, since the purpose of incorporating propensity scores is to achieve in-sample balance rather than in the population, weighting the propensity score by the survey weight is unnecessary (Zanutto, 2006; DuGoff et al., 2014). The authors also recommend including survey weights as covariates in the propensity model, drawing a parallel between this intuition to the recommendation by Lumley (2011) of including these covariates to account for non-response in surveys (Lumley, 2011). However, this simple inclusion is often problematic in practice, as it may lead to separation in the data or large biases (Ridgeway et al., 2015). DuGoff et al. (2014) also addressed limitations regarding how to handle extreme weights, model misspecification, nonlinear propensity score estimation, and additional covariates in the models. Unfortunately, these recommendations are largely based on intuition, necessitating more rigorous arguments of consistency and conditional independence.
Subsequently, Ridgeway et al. (2015) began to formalize the discussion of propensity score analysis for complex survey designs. The authors demonstrated through proof and simulation that not appropriately accounting for the selection mechanism can potentially lead to biased estimates. A natural question is in what stages of the analysis (i.e., the propensity score and/or outcome model stage) should the survey weights be included? The authors recommended that researchers use survey weights in both stages of the based on their results. However, the simulation results pertaining to our scenario are inconclusive as to which approach achieves the lowest bias, and the derivations of consistency are not specific to this scenario. Further, they consider situations with only the population average treatment effect on the treated (PATT) as their target estimand (Ridgeway et al., 2015).
Similarly, others have addressed the task of combining causal inference and survey methodology. Austin et al. (2018), for example, focused primarily on propensity score matching techniques in simulation. However, these results did not provide a clear answer with respect to which method of estimating the propensity score model, survey-weighted or not, was best. Dong et al. (2020) explored the performance of different propensity and outcome weighted models and suggested the use of survey weights at both stages, though they concluded the estimates from different strategies were not statistically different in their data example. Lenis et al. (2018) further studied the causal inference framework for estimating population versus sample effects and explored the impact of non-response mechanisms with matching techniques. Their simulation setting followed that of Austin et al. (2018) and conclude that the degree of misspecification impacted their analysis results and population balance. In our setting, the sample selection weights are supplied by NHANES, however the true sample selection mechanism is unknown. Further, there are many SES variables collected in the survey, which are potential covariates in selection and propensity score models. Here, considerations as to the limited degrees of freedom in the modeling process with the addition of survey weights, strata, and PSU’s may be especially important. Lastly, race is part of the selection mechanism in the NHANES design, which complicates the relationships underlying our motivating research question.
3 Methods
3.1 Notation
Let $Y$ denote our observed outcome, here telomere length. Our goal is to estimate the de-confounded, or controlled difference in mean telomere length between non-Hispanic White and non-Hispanic Black-identifying individuals in the United States. To that end, let $\mathcal{A}$ be the discrete set of potential comparison groups of interest, such that $a\in\mathcal{A}$ denotes a specific group. We are interested in comparing the values of $Y$ across levels of $A$. Our motivating analysis departs from the traditional causal literature in that $\mathcal{A}$ represents race, with $a=1$ denoting Black race and $a=0$ denoting White race, rather than a potential ‘treatment,’ per se (see Li and Li, 2023, for further discussion on using propensity scores for racial disparities analysis). Given the nature of our variables, and race in particular, this contrast of interest will primarily be viewed as a controlled comparison of the observed outcomes after balancing the two groups with respect to potential confounding variables (Li et al., 2013; Li and Li, 2023). As the lived experience of Black and White individuals differs significantly in terms of social determinants of health, we believe a controlled descriptive comparison can help inform the dynamics of these complex relationships (Li and Li, 2023). While race may not be interpreted as a modifiable treatment, we conceptualize race as an important construct worthy of study. Moreover, the methods described herein are equally applicable in settings where a participant’s sample selection is dependent on a treatment condition or exposure status in a more traditional causal inference framework (see Section 3.2 for more details).
Our observed data, $\mathcal{D}=\{(Y_{i},A_{i},X_{i},S_{i}=1);i=1,\ldots,n\}$, consist of $n$ independently and identically distributed copies of $(Y,A,X,S)$ sampled from a population distribution, $\mathcal{P}$, of $N$ individuals where $Y$ is the observed outcome, $A$ is a participant’s group membership, $X$ is a $p$-dimensional vector of covariates, and $S$ is an indicator for being selected into the sample. In our motivating example, $\Pr(S_{i}=1)$ is known for each individual $i$ in the super population through the explicit NHANES sampling design. There is also the possibility of having missing outcome values among sampled individuals, but this situation will not be explored here and is instead left to future work. Further, our assumed variable relationships are given in Figure 1. Here, we can see that our outcome ($Y$), the variable of interest ($A$), and the sample selection indicator ($S$) all depend on a particular set of covariates ($X$), and sample selection also depends on race.
3.2 Proposed Estimators
We are interested in drawing inference about the super population of individuals from which we have selected our study participants. Here, the targets of inference are the conditional outcome means, i.e., $\mathbb{E}\left[Y\mid A=a,X\right]$ for $a\in\mathcal{A}$ and their expected difference over $X$, i.e.,
${\rm ACD}=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=1,X\right]-\mathbb{E}\left[Y\mid A=0,X\right]\right].$ (1)
Our goal is to estimate the individual group specific conditional means, of which ACD is the difference. We make the following assumptions:
  1. (i)
    Positivity: $\Pr(A=a\mid X=x)>0\ \forall a\in\mathcal{A}$ and every $x$ such that $f_{X}(x)>0$
  2. (ii)
    Positivity of Sample Selection: $\Pr(S=1\mid A=a,X=x)>0$ for every $a,x$ such that $f_{A,X}(a,x)>0$
  3. (iii)
    Weak Selection Exchangeability: $\mathbb{E}[Y\mid A=a,X]=\mathbb{E}[Y\mid A=a,S=1,X]$
We require (i) and (ii) to ensure that our observed data functional, denoted in the next section by $\mu(a)$, is well defined. Condition (iii) allows us to generalize the conditional mean of the outcome from the study sample to the target population. This assumption is less stringent than exchangeability in distribution (i.e, $Y\perp S\mid A,X$), as it does not require the entire conditional distribution of the outcomes to be independent of selection, only their conditional expectations. Depending on the factorization of the joint probability of being sampled $(S=1)$ and belonging to a particular group $(A=a)$, we propose three formulas, which can be estimated using either outcome model–based estimators (i.e., g-computation) or inverse probability weighted estimators:
$\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right],$ (2)
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right],$ (3)
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right].$ (4)
We will refer to the methods for estimating Equations (2), (3), and (4) as Outcome Modeling and Direct Standardization (OM), Inverse Probability Weighting 1 (IPW1), and Inverse Probability Weighting 2 (IPW2), respectively. For additional details on this derivation, see Appendix A.1.
The assumptions required to estimate the ACD are relatively weaker than those required to identify a causal quantity. Figure 1, Panel (C) gives an example directed acyclic graph (DAG) where a causal target may be warranted, such as the case where $A$ is an environmental exposure, like lead, which can also alter telomere length. We provide a supplementary analysis in which we estimate the causal effect of lead exposure on telomere length for illustration, as a practitioner may also be interested in such causally interpretable estimated differences. If the group variable under study is a treatment or exposure, we show under stronger assumptions that Formulas (2) - (4) can equivalently target the population potential outcome means, denoted by $\mathbb{E}[Y^{a}]\ \forall a\in\mathcal{A}$. Along with the previous assumptions, the following conditions,
  1. (iii*)
    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}[Y^{a}\mid A=a,X]=\mathbb{E}[Y^{a}\mid A=a,S=1,X]$
  2. (iv)
    Stable Unit Treatment Value Assumption (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
  3. (v)
    Weak Exchangeability: $\mathbb{E}[Y^{a}\mid X]=\mathbb{E}[Y^{a}\mid A=a,X]$,
are sufficient to identify $\mathbb{E}[Y^{a}]$. The additional SUTVA assumption underpins the existence of the counterfactual outcomes, $Y^{a}$. The compact representation in (iv) captures two aspects of SUTVA – no interference, which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and consistency, which states that the potential outcome is well-defined for each treatment level, $a$ (Rubin, 1980). The weak exchangeability assumption allows us to represent the conditional distribution of the unobserved potential outcome using that of the observed potential outcome. As before, we only require a weaker version of this assumption, though strong exchangeability, i.e., $Y^{a}\perp A\mid X$, is often assumed. For additional details and the full identification proof, see Appendix A.2. This is thematically similar to Dahabreh et al. (2020), who study the generalization of causal inferences from individuals in randomized trials. The authors also differentiate between mean exchangeability in $A$ for the randomized population and mean exchangeability in $S$ for the target population in the context of randomized trials.
3.3 Comparisons to Existing Approaches
The three formulas derived above differ from standard, existing approaches that one might take in practice, which warrants some brief discussion here. Table 1 compares our proposed approaches to such existing methods. As shown, there are two sources of bias if our assumed conceptual diagram holds – bias from covariate imbalance between comparison groups (referred to herein as confounding bias) and unequal selection induced by the design of the complex survey (referred to herein as selection bias), which impacts the generalizability of our within-sample results to a larger, target population. Existing approaches may incorporate either or both balancing and/or generalizability weights to account for these two sources of bias, however if either are foregone, then we expect to obtain biased estimates of our target quantity of interest. For instance, the estimated (unadjusted) effect of race can be either exaggerated or attenuated depending on the direction of the true association between the confounders (e.g., SES), race, and the outcome in simple regression (Approach 1 in Table 1; Williams et al., 2019). Multiple regression (Approach 2) further adjusts for covariates, but unless the outcome model includes the appropriate interaction terms, this approach will suffer from confounding bias as well (Rao, 1971). With inverse probability weighting (Approach 3), the race-covariate relationship is modeled, rather than the covariate–outcome relationship, foregoing the need to specify any race $\times$ covariate interactions (Mansournia and Altman, 2016). However, with all three of these approaches, the target of inference is the within-sample difference in outcomes, which cannot be generalized to a larger population if a probability sample was taken (Shi et al., 2023; Degtiar and Rose, 2023). Survey-weighted regression (Approach 4) accounts for selection bias by accounting for how many individuals in the population that one individual in the sample represents (Lumley and Scott, 2017), allowing for generalizability, but these results will be biased if there is covariate imbalance in the sample (DuGoff et al., 2014; Ridgeway et al., 2015).
Table 1: Comparison of existing approaches for analysis to our proposed approach. Dashes denote that the model/weight is not used.
IPTW = Inverse Probability of Treatment Weighting
$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X)$
A more principled approach would be one that accounts for both confounding and selection bias, as given by Approaches 6 and 7, which utilize a combination of propensity score and survey weighting. These approaches are advocated for by Zanutto (2006); DuGoff et al. (2014) and Ridgeway et al. (2015), respectively, and may be viewed as current analytic standard since they suggest simple survey-weighted generalizations of the usual use of IPTW, either at the first (propensity model) and/or second (outcome model) stage. Though it may seem that either simple extension addresses both sources of bias, there has been little theoretical justification of how weighting one or multiple models within IPTW accomplishes both tasks. This is also true of the ‘naive g-computation’ approach (Approach 8) we study, which weights the predicted conditional expectation for each individual’s outcome and comparison group/race by their survey contribution. Additionally, propensity scores are often seen as in-sample balancing scores (Zanutto, 2006; DuGoff et al., 2014). From this perspective, their generalizability (and corresponding selection bias correction based on survey weights) is often ignored. Survey-weighting both the propensity score and outcome model is reasonable (Ridgeway et al., 2015), however directly combining survey-weighted IPWs and survey weights in a second stage outcome regression may also lead to biased results if additional correction is not made. In our research question, we have an intersecting consideration of both selection and confounding bias; thus, neither source of bias should be disregarded.
From a practical standpoint, one question at hand is: “should we implement survey-weighted propensity score estimation in IPTW or not”. Our analytic results show that for estimation of the ACD, the propensity score does not have to be survey-weighted. We see this in Equations (3) and (4), which differ with respect to the factorization of the joint distribution into propensity and survey weight models. If the propensity score is not survey-weighted, based on Equation (4), we need $\Pr(S=1\mid X)$, i.e., one may need to marginalize $\Pr(S=1\mid A,X)$ over the group membership variable, $A$. If survey weights are included in the propensity score estimation, then we must know $\Pr(S=1\mid A,X)$, i.e., the probability of selection conditional on each level of the group membership variable and the covariates.
Moreover, our proposed approaches carry additional differences amongst themselves, and prove to be useful in different settings. Formula (2), outcome modeling and direct standardization (OM), requires individual-level data on $X$, or knowledge of $f_{X}(x)$, as an outcome regression, $g_{a}(X)$, is fit in practice. Formulas (3) and (4), IPW1 and IPW2, do not require estimation of an outcome model and may be the best options depending on how much information is available in practice. We view IPW1 and IPW2 as simple corrections to what one might do in practice, namely the traditional inverse probability of treatment weighted (IPTW) estimator, and thus may be a better starting point for an analysis of this type as they additionally account for selection bias. Further, IPW1 and IPW2 do not assume additivity, and thus may be more robust if the outcome model for (2) is misspecified. However, while OM, IPW1, and IPW2 are all consistent, OM is most efficient if correctly specified. Thus, depending on the availability of the data or the ability to estimate certain key quantities, a researcher may prefer one approach over the others.
Regarding the incorporation of propensity and sample selection weights into the analysis, our weights are not the same as those in the traditional g-computation or IPTW frameworks. We derived a new estimator with weights both for within-sample inference and for generalizability to our target population. Specifically, note the term $\Pr(S=1\mid X)$ in the denominators of Formulas (2) and (4). If selection does not depend on $A$, then identification reduces to the standard proof for the IPTW identification formula. However, in our setting, selection depends on the participant’s self-reported race. Thus, we need the following decomposition to calculate $\Pr(S=1\mid X)$:
$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X).$
As with OM, IPW2 requires that the probability of selection be marginalized over $A$, i.e., $\Pr(S=1\mid X)$, and that the propensity model be fit without the incorporation of selection weights, i.e., $\Pr(A=a\mid S=1,X)$. IPW1 requires the selection mechanism be known (e.g., in the case of survey weights), conditional on $A$, i.e., $\Pr(S=1\mid A=a,X)$, and that a survey-weighted propensity model be fit, i.e., $\Pr(A=a\mid X)$. The key difference here comes with respect to how the joint propensity and selection mechanisms are factorized.
3.4 Estimation
To estimate the ACD based on the three formulas presented above (OM, IPW1, and IPW2), we propose the following. For OM, define the functional
$\mu(a)=\mathbb{E}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right].$
We denote the estimator for the expectation of the outcome, $\mathbb{E}[Y\mid A=a,S=1,X]$, by $\hat{g}_{a}(X)$. In practice, $\hat{g}_{a}(X)$ can be estimated, for example, using a standard parametric outcome model with a finite-dimensional parameter vector. Our outcome model-based estimator for $\mu(a)$ is given by
$\hat{\mu}_{\rm OM}(a)=\frac{1}{n}\sum_{i=1}^{n}\hat{g}_{a}(X_{i})\frac{\Pr(S_{i}=1)}{\Pr(S_{i}=1\mid X_{i})}.$ (5)
[…]
$\hat{\mu}_{\rm IPW1}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(A_{i}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}^{w}_{a}(X_{i})Pr(S_{i}=1\mid A_{i}=a,X_{i})}\quad{\rm or}$ (6)
$\hat{\mu}_{\rm IPW2}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(\mathcal{A}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}_{a}(X_{i})Pr(S_{i}=1\mid X_{i})}.$ (7)
[…]
3.4.1 Estimation of Uncertain Sample Selection or Non-Probability Samples
[…]
3.5 Inference
[…]
$\phi(Y,X,\theta,\mathcal{P})=M^{-1}\psi[Y,X,\theta(\mathcal{P})],$
where
[…]
is the Jacobian matrix of the estimating equations (Huber, 1996). Using numerical optimization, the covariance matrix of the estimated parameters, $\hat{V}(\hat{\theta})$, including the ACD estimator, is given by
$\hat{V}(\hat{\theta})=\sum_{i=1}^{n}\phi_{i}(Y,X,\hat{\theta},\mathcal{P})\phi_{i}(Y,X,\hat{\theta},\mathcal{P})^{\prime},$ (8)
where $\phi_{i}(Y,X,\hat{\theta},\mathcal{P})$ is the empirical evaluation of the $m\times 1$ vector of the $i$th observation’s values of the influence function for $\hat{\theta}$ by replacing the unknown parameters with sample estimates.
3.5.1 Inference in Stratified, Clustered Multi-Stage Survey Samples
NHANES employs a complex stratified, clustered sampling design. While bias incurred from the differential sample selection probabilities (encoded in the survey weights) is accounted for in the estimation procedure, inference is affected by the fact that individuals within strata and primary sampling units (PSUs) are correlated. To account for this, we modify this linearized variance estimator, known as the robust or sandwich estimator outside of survey-based analysis, by summing the contributions of the $i$th individual $(i=1,\ldots,n_{j})$, nested in the $j$th primary sampling unit (PSU; $j=1,\ldots,J_{k}$), nested in the $k$th sampling stratum $(k=1,\ldots,K)$. Thus, the estimated covariance matrix for our parameters is expressed in terms of the variability of the between PSU-level sums of the first order Taylor approximations within the sampling strata. This becomes
$\sum_{k=1}^{K}\frac{J_{k}}{J_{k}-1}\sum_{j=1}^{J_{k}}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}^{\prime}$ (9)
where $K$ and $J_{k}$ are the number of strata and number of PSUs within the $k$th stratum, respectively, $\phi_{\cdot jk}(y,\hat{\theta},\mathcal{P})=\sum_{i=1}^{n_{j}}\phi_{ijk}(y,\hat{\theta},\mathcal{P})$ is the sum of the influence function contributions among individuals in the $j$th cluster, and $\bar{\phi}_{\cdot\cdot k}(y,\hat{\theta},\mathcal{P})=\frac{1}{J_{k}}\sum_{j=1}^{J_{k}}\phi_{\cdot jk}$ is the “average” influence function value within the $k$th stratum. Software for these methods can be found in the R package svycdiff at https://github.com/salernos/svycdiff, in addition to the code and data necessary to reproduce the simulation studies and data analysis described below.
4 Simulation Studies
We conduct a series of simulation studies to better understand the finite sample performance of our proposed estimators, as compared to approaches typically used in practice. Based on the simulation studies conducted in DuGoff et al. (2014) and Ridgeway et al. (2015), we extend this work in several important ways. First, we have our proposed estimators, which represent novel alternatives to various combinations of weighting schemes outlined in Table 1. Second, we have varied the set of associations in our data to systematically include or exclude associations between $A$, $X$, and $S$ in our data generating mechanism to understand how certain relationships impact our estimation and inference. The purpose of the overall simulation study is to quantify the bias and variance associated with the use of our proposed approach, as compared to the various weighting schemes which may be currently employed.
4.1 Data Generation
As in DuGoff et al. (2014) and Ridgeway et al. (2015), let $X$ denote a normally-distributed covariate which relates to our outcome, $Y$, race, $A$, and sample selection indicator, $S$. We generate population-level data with $X\sim\mathcal{N}(1,1)$ for a super population of $N=100,000$ individuals from which we repeatedly take samples. We then generate the remaining data from three models:
  1. 1.
    Propensity Model: $\Pr(A\mid X)={\rm logit}^{-1}(\tau_{0}+\tau_{X}X)$,
  2. 2.
    Sample Selection Model: $\Pr(S\mid A,X)={\rm logit}^{-1}(\beta_{0}+\beta_{A}A+\beta_{X}X+\varepsilon_{S});\ \varepsilon_{S}\sim\mathcal{N}(0,0.1)$,
  3. 3.
    Outcome Model: $f(Y\mid A,X)$,
where ${\rm logit}^{-1}(u)=[1+\exp\{-u\}]^{-1}$ is the inverse logit function and we add a small amount of noise to the generation of the sample selection indicator to model this variable in our later procedure. From these models, we simulate the binary variable of interest, $A\mid X\sim\text{Bin}[N,\Pr(A\mid X)]$, the sample selection indicator, $S\mid A,X\sim\text{Bin}[N,\Pr(S\mid A,X)]$, and sample selection weights, $\omega_{S}=\Pr^{-1}(S\mid A,X)$. In simulating the outcome $Y$, we consider a heterogeneous treatment effect which varies with $X$, such that
$Y=\gamma_{0}+\gamma_{X}X+\gamma_{A}A+\gamma_{AX}AX+\varepsilon_{O}\$
where $\varepsilon_{O}\sim\mathcal{N}(0,1)$. We consider eight simulation settings, over which we vary the coefficients $\tau_{X},\beta_{A}$, and $\beta_{X}$, corresponding to the parameters in the true propensity and selection models (see Figure 2a-b), where each setting corresponds to a progressively more complex relationship among the variables. Across all simulation settings, we fixed $\tau_{0}=-1$, $\beta_{0}=-4.5$, $\gamma_{0}=1$, $\gamma_{A}=1$, $\gamma_{X}=1$, and $\gamma_{AX}=0.1$. Within each setting, we drew $200$ sub-samples based on our simulated sample selection indicator, $S$. We fit the various combinations of approaches listed in Figure 2. Our target quantity of interest was the ACD. We compared the percent bias, mean squared error (MSE), and coverage probability at a significance level of 0.05 for each method.
4.2 Results
Results from our simulation studies are given in Figure 2c-d and Supplemental Table LABEL:tab:simulation_results. Here, we discuss these results in the context of each simulation setting, where the data are generated to vary what sources of bias we expect to encounter when estimating the ACD.
4.2.1 Settings 1 and 3: No Bias Expected
In Settings 1 and 3, there is neither confounding nor selection bias in the data generating mechanism. Setting 1 provides a fair comparison across all the methods considered, as both the group propensity, $\Pr(A\mid X)$, and sample selection probabilities, $\Pr(S\mid A,X)$, were generated independent of any covariates. As shown, all of the methods except for the naive g-computation approach perform comparably in this setting with respect to the percent bias of the ACD estimator. Specifically, the percent bias among the three proposed estimators ranged from 0.499-0.647%, as compared to 0.166-0.538% among the existing methods (excluding the naive g-computation), suggesting that all of the methods considered accurately estimate the average controlled difference. This is expected, as the methods which do not account for confounding or selection bias are correctly specified, and the methods that do are unencumbered by the additional weights. The proposed approaches also had among the lowest MSEs, ranging from 0.475-0.480, as compared to 0.482-1.103 among the existing approaches. Lastly, across 200 independent replications, the proposed methods had coverage rates of 92.0-92.5%, as compared to 87.5-93.0% for the existing methods. Similarly, in Setting 3, where selection depends on $A$ but neither selection nor $A$ depend on $X$, all methods tend to perform well. Here, we see that the percent bias for the proposed methods ranged from 0.217-1.497%, on par with the other methods (0.138-0.417%). Mean squared errors ranged from 0.219-0.243 for the proposed methods, similar to MSEs of 0.222-0.547 among the existing methods. Further, all of the methods had nominal coverage rates of 94.5-99.0%.
4.2.2 Setting 2 and 4: Confounding Bias Expected
In Settings 2 and 4, we generate $A$ conditional on $X$, with selection further being dependent on $A$ in Setting 4. In these settings, we now expect there to be confounding bias, as $\Pr(A\mid X)$ is no longer uniform over the range of $X$ values. As such, we see that all methods which account for both $A$ and $X$ tend to perform comparatively well, whereas the simple linear regression (SLR) of $Y$ on $A$ performs poorly due to confounding bias in $X$. Specifically, we see that the percent bias in the SLR estimator has increased from 0.138-0.166% in Settings 1 and 3 to 78.249-79.480% in Settings 2 and 4, whereas the other estimators (including the proposed methods) still see biases ranging in magnitude from 0.977-2.778% in Setting 2 and 0.210-1.526% in Setting 4. We further see an increase in MSE for the SLR estimator, with values ranging from 75.408-75.638, while the MSE for the other existing methods (0.404-1.176) and the proposed estimators (0.362-1.220) remained low. Once again, the coverage probabilities for the proposed (92.5-94.5%) and other existing (88.5-94.0%) remained at the nominal level, whereas the SLR did not cover the true ACD (0%).
4.2.3 Settings 5 and 7: Selection Bias Expected
In Settings 5 and 7, we now generate $S$ conditional on $X$, with selection further being dependent on $A$ in Setting 7. Here, we now expect there to be selection bias induced by the $S\mid X$ relationship. In Setting 5, the percent bias among the existing approaches ranges from 9.182-9.544%, while the bias among the proposed methods ranges from 0.302-1.252%, demonstrating empirically that the proposed estimators are unbiased for the target ACD. Moreover, these results persist in Setting 7, with the percent bias among the existing approaches (3.992-6.809%) being higher than the proposed approaches (0.172-2.322%). Similar patterns can be seen when comparing the MSEs of the existing approaches (1.135-1.311 in Setting 5 and 0.587-0.663 in Setting 7) to those proposed in this work (0.236-0.325 in Setting 5 and 0.135-0.265 in Setting 7). Further, only the proposed methods achieved the nominal coverage rate (93.5-95.5% in Setting 5 and 91.0-97.0% in Setting 7), while the other methods show coverage ranging from 9.0-36.5% and 19.0-80.5% in Settings 5 and 7, respectively.
4.2.4 Settings 6 and 8: Both Confounding and Selection Bias Expected
Lastly, Settings 6 and 8 represent the most complex scenarios of confounding and selection bias, with Setting 8 mirroring the conceptual diagram for our motivating data. In this setting, our proposed estimators outperform all current approaches with respect to the percent bias, MSE, and coverage for estimating the ACD. In particular, percent bias among existing approaches ranges from 7.847-82.816% in Setting 6 and 6.226-73.914% in Setting 8, whereas the bias among the proposed estimators ranges from 0.633-1.514% and 1.083-1.115%, respectively. This is expected as, the proposed methods have the correct balancing and generalizability weights under these settings, whereas the existing approaches fail to incorporate one or both weights, or fail to incorporate them correctly into the form of the estimator. The proposed approaches again had the lowest MSEs, ranging from 0.131-0.540 across both settings, as compared to 0.927-82.577 for the existing approaches. Lastly, the proposed approaches achieve nominal coverage (93.5-97.0% across both settings), whereas the existing methods do not (0-83.5%). For full results, see Supplemental Table LABEL:tab:simulation_results.
4.3 Sensitivity Analysis
In a sensitivity analysis, we study our proposed approaches under increasing model misspecification by introducing effect heterogeneity. Specifically, we systematically included larger interactions of race and covariates, where we varied $\gamma_{AX}\in\{0,0.01,0.05,0.1,0.5\}$. We estimate IPW1 and IPW2 as is, and we fit two versions of OM where we (1) correctly specify versus (2) misspecify the outcome model by not including the interaction term present in the data generating mechanism. Here we seek to assess the robustness of these approaches against the oracle estimator. The results from this analysis show that overall, all methods tend to perform well in terms of the bias, MSE, and coverage of the estimator, with OM performing acceptably under moderate misspecification. In particular, for $\gamma_{AX}\in\{0,0.01,0.05,0.1\}$, the method with the misspecified outcome model attained nominal coverage (ranging from 93.5-96.5%), with biases below 0.05% (0.008-0.028) and comparable MSEs to the other methods (0.001-0.003). However this approach incurs bias as treatment heterogeneity becomes more evident. Specifically, for the setting in which $\gamma_{AX}=0.5$, we see that the bias increases to 0.14% and the method does not cover the true ACD. These results suggest that the IPTW-type estimators (IPW1 and IPW2), which do not assume additivity, may be more robust. Full simulation results are presented in Supplemental Table LABEL:tab:sim_sensitivity.
4.4 Summary Takeaways
In summary, the simulation settings described above are designed to vary the associations between $Y$, $A$, $X$, and $S$ with increasing complexity, in order to introduce either/both confounding and/or selection bias into the data generating mechanism. In settings with neither confounding nor selection bias, all existing and proposed methods perform comparably well, estimating the ACD with low bias and MSE, and nominal coverage. In settings of confounding bias (without selection bias), the methods which account for both $A$ and $X$, namely all of the multiple regression and IPTW-based estimators, tend to perform comparatively well, whereas methods that only model $A$ suffer from high bias and negligible coverage. When $X$ is introduced into the generation of either/both $S$ and $A$, that is, when selection or selection + confounding bias exists, we see that standard analytic approaches lead to biased estimates and do not attain nominal coverage, whereas the proposed methods perform empirically well, supporting theoretical results derived in the previous section. Lastly, our sensitivity analysis reveals that the proposed methods which rely on contrasts and inverse probability weighting, rather than outcome modeling and direct standardization, are more robust to misspecification, particularly as the degree of effect heterogeneity between the groups of interest increases.
5 NHANES: Race, SES, and Telomere Length
We now turn our attention to our motivating example, which comes from the National Health and Nutrition Examination Survey (NHANES). NHANES is a nationally representative survey sample collected by the Centers for Disease Control, which contains a wealth of demographic, dietary, physical examination, laboratory, and questionnaire data on individuals. Our variable of interest is the survey participant’s self-reported race. Racial groups are categories of people treated as a social entity by virtue of physical characteristics. While an individual’s race is often self-reported, racial groups may represent commonalities such as geographic origin, genetic ancestry, cultural norms and traditions, social history, or current social circumstances (Williams et al., 2010). From a statistical perspective, race is often a proxy for other sources of shared variation which must be accounted for when making inferential statements. We focus on non-Hispanic participants who self-identify as Black or White as, particularly in the US, it is known that Black individuals tend to have poorer health outcomes than White individuals due to structural inequalities (Arias and Xu, 2022).
5.1 Motivating Question
Telomeres are regions of DNA at the ends of chromosomes that protect against cellular aging and senescence (Lin et al., 2016). Telomere shortening is associated with disease incidence and a variety of health outcomes (Shammas, 2011). Variation in telomere lengths is attributable to age and shortening during mitosis, in addition to genetics, social history, exposure to environmental contaminants, and psycho-social stressors (Jiang et al., 2007; Blackburn et al., 2015). Paradoxically, Black individuals typically have longer telomeres than White individuals (Needham et al., 2013; Venkateswaran et al., 2024), with plausible explanations citing leukocyte cell composition, population stratification, or genetic differences (Freedman et al., 1997; Lin et al., 2016; Hansen et al., 2016). Others posit that environmental exposures associated with longer telomere length underlie this so-called ‘race effect,’ as these exposures differ significantly between Blacks and Whites (Needham et al., 2020; Roberts et al., 2022). Further, studies have found telomere length to be comparable between Black and White individuals in socio-economically homogeneous populations (Geronimus et al., 2015), calling into question purely genetic explanations. Our motivating question is this: if we could hypothetically balance socioeconomic status (SES; $X$) between Black and White individuals ($A$) in a nationally representative sample ($S$), would we still see significant Black/White differences in telomere length ($Y$)?
Figure 2: Simulation setup and results. Columns denote each of eight simulation settings. The top row gives the assumed relationships between variables for each setting. The second row gives the parameter settings for each data generating mechanism. The third row gives the average percent bias and Monte Carlo standard errors for each method across 200 simulation replicates. The fourth row gives the coverage probabilities for each method.
5.2 NHANES Sampling Design
The stratified, clustered sampling design of the NHANES survey is designed to over-represent Black individuals and underrepresented groups in particular, to achieve a more nationally representative sample. Namely, NHANES samples in four stages. In the first stage, primary sampling units (PSUs; i.e., counties) are selected. Counties are stratified and selected with probability proportional to their weighted average population counts, where the weights give a higher probability of selection to counties with a higher proportion of underrepresented subgroups (e.g., certain age groups, marginalized racial groups, lower-income households). In the second stage, the PSUs are further divided (e.g., into city blocks) and sub-sampled, and in the third stage, households are sampled from these sample segments. Again, households within segments that have a greater proportion of underrepresented age, ethnic, or income groups are given additional weight for being selected. Lastly, individuals within households are sampled according to designated age-sex-race/ethnicity screening subdomains. This complex sampling scheme motivates our proposed methods, as sample selection depends on self-reported race and potential confounders (i.e., socioeconomic status).
5.3 Study Sample
The 1999-2002 NHANES waves consisted of 7,839 participants with recorded laboratory measures, representing our survey sample $(S=1)$. Our primary endpoint, $Y$, was the log-transformed mean ratio of an individual’s telomere length to a standard reference DNA sample across all leukocyte cell types (mean T/S). Telomere length was assayed via quantitative polymerase chain reaction (PCR; Cawthon, 2002; Needham et al., 2013). We focus on the 1999-2002 NHANES waves, as they featured 4-year adjusted survey weights designed for aggregating data across cohorts. Among the initial 7,839 participants, 5,308 (67.7%) self-identified as either non-Hispanic White or non-Hispanic Black, the levels of our group variable $A$. Excluding those participants without our outcome of interest, our final analytic sample contained 5,298 Non-Hispanic White or Non-Hispanic Black identifying participants with measured telomere length. As discussed previously, race was our variable of interest. We further adjusted our models for study participant age, sex, and blood cell composition, including complete blood count and five part differential consisting of the proportion of neutrophils, eosinophils, basophils, lymphocytes, and monocytes, to account for known differences in these factors (Freedman et al., 1997), as well as twelve indicators of SES, collectively representing our confounding variables $X$. Ten of these, namely marital status, education level, household income, insurance status, Special Supplemental Nutrition Program for Women, Infants, and Children (WIC) usage, household size, home ownership, home type, food security status, and an individual’s poverty income ratio (PIR), were extracted directly from the NHANES demographic and occupation questionnaires. Occupation category was constructed by mapping occupation group codes in the NHANES 1999-2002 occupation questionnaire to the national statistics socioeconomic job classifications, and employment status was derived from three occupational measures: type of work done last week, hours worked last week at all jobs, and main reason for not working last week (Rehkopf et al., 2008; Rose et al., 2005). In later modeling, we perform complete case analysis on $n=5,270$ participants after removing 28 individuals without measured blood composition.
Descriptive statistics for our analytic sample are presented in Table LABEL:tab:nhanesdesc, both overall and stratified by self-reported race. As shown, non-Hispanic Black participants had a median telomere length (mean T/S ratio) of 1.05 (i.e., 0.05 times longer than the standard reference DNA), as compared to non-Hispanic White participants with a median telomere length of 0.98. With respect to the identified socioeconomic characteristics, non-Hispanic Black versus White participants differed significantly across all 12 SES indicators, with Black participants having lower educational attainment, lower household income, and lower home ownership rates, as well as higher food insecurity and WIC utilization rates, larger household sizes, and a higher proportion uninsured, unemployed, and below the federal poverty line. Further, a higher proportion of Black participants reported not working or working in low blue collar occupations, a higher proportion reported being widowed, divorced, or separated, and a higher proportion reported residing in apartments, all of which suggest lower socio-economic status (Rose et al., 2005). Based on these factors, we then calculated propensity scores with race as the outcome using unweighted and survey-weighted logistic regressions. The results from these models fit are given in supplementary Table LABEL:tab:nhanespropscore. The propensity scores from these models were then carried forward to our later outcome models in order to estimate the ACD of telomere length by race.
5.4 Results
We compared our proposed approaches to six methods that one might consider in practice, each of which successively increase in their respective adjustments for potential confounding and selection bias. Namely, we fit a multiple linear regression (Approach 2 in Table 1) of log-transformed telomere length on race, age, sex, the blood composition variables, and the socioeconomic indicators, not accounting for the potential confounding or selection bias in the study. We then fit a survey-weighted multiple regression ((Approach 4), and inverse probability of treatment weighted (IPTW) multiple regression (Approach 5), a regression weighted by the product of the within sample propensity score weights and the survey weights (Approach 6), and a regression weighted by the product of the survey-weighted propensity score weights and the survey weights (Approach 7), not unlike our simulation study. We omitted the oracle regression, simple linear regression, and the naive g-computation methods, as they were not applicable (as with the oracle regression) or inadvisable (as with the simple linear regression and naive g-computation). For the proposed methods, standard errors for the ACD were calculated via Equation (9), whereas for the existing methods, we calculate either model-based or design-based standard errors where appropriate.
ACD estimates and corresponding 95% confidence intervals across the various analytic approaches are presented in Table 3 below. In general, the ACD estimates across different analytic strategies tend to attenuate in magnitude as we made more appropriate adjustments for confounding and selection bias. For example, the multiple linear regression had an ACD estimate of 0.0265, corresponding to approximately a 2.7% longer telomere length, on average, among non-Hispanic Black individuals as compared to non-Hispanic White individuals, adjusting for all other factors (95% CI: 0.0106-0.0423). In contrast the ACD point estimates from our proposed approaches varied from 0.0132 (IPW2) to 0.0176 (OM), though these average controlled differences were not statistically different from zero. These results suggest that there is a strong confounding relationship between SES and race. Moreover, the methods which properly incorporated the NHANES stratified, clustered design tended to have more conservative standard error estimates than those which did not, as evidenced by the wider confidence intervals (Table 3).
These results align with our analytic and simulation results in two important ways. Firstly, we show analytically that if our conceptual diagram holds, the proposed approaches will lead to consistent estimates of the ACD, whereas methods which do not account for confounding and/or the sampling design appropriately will incur bias in estimating the ACD. This is further evidenced in our simulations, which show that the ACD is overestimated, on average, in the settings which closely reflect our assumed conceptual diagram for these data. The ACD estimates for the proposed approaches are roughly 50% those of the existing approaches. Secondly, the methods which respect the complex survey design tend to have wider confidence intervals, appropriately, as individuals within sampling units are correlated.
5.5 Supplementary Analysis: Lead Exposure, SES and Telomere Length in NHANES
In a supplementary analysis, we present a different example from NHANES which may be amenable to a causal interpretation (in the sense that counterfactual/potential outcomes and plausibly changing exposures are more meaningful). To illustrate how the proposed methods can, under stronger assumptions, identify the average difference in potential outcome means (population average treatment effect; PATE), we study the effect of exposure to a particular environmental contaminant, lead, on telomere length (Zota et al., 2015). Lead exposure is associated with negative health outcomes, such as cardiovascular and kidney diseases (Navas-Acien et al., 2007, 2009), and biological markers of oxidative stress and inflammation (Lin et al., 2016). Recent works suggest that lead exposure may be associated with telomere shortening (Lin et al., 2016; Wu et al., 2012; Zota et al., 2015), however, evidence from these studies is mixed and depends on the analytic strategy. Thus, further exploring the estimated effect of lead exposure on telomere length with our proposed approaches may be of interest.
Blood lead levels were measured at the CDC’s National Center for Environmental Health (Atlanta, Georgia) using a simultaneous multi-element atomic absorption spectrometer (SIMAA 6000; PerkinElmer, Norwalk, Connecticut) with Zeeman background correction (Centers for Disease Control and Prevention, 2001a, b). The limits of detection (LOD) were 0.3 $\mu$g/dL for blood lead, with < 1% of study participants having blood lead concentrations below the LOD. Among these patients, we imputed their blood lead concentrations by LOD divided by the square root of 2 (Hornung and Reed, 1990). To define our exposure, we dichotomized LOD-imputed blood lead concentrations at the median $\mu$g/dL to study the difference in potential outcome means between ‘higher’ and ‘lower’ exposure groups.
We repeated our analysis above, treating the dichotomized lead variable as our exposure $(A)$. As before, we fit a propensity score for lead exposure conditional on the twelve socioeconomic indicators $(X)$ described above. We further included self-reported race using two strategies – first as a precision covariate along with age and blood composition (i.e., not in $X$) and second, in a sensitivity analysis, as a part of our confounders, $X$. The results of this analysis are presented in Supplementary Table C2. While the point estimates and findings differed across methods, we found a significant exposure effect for both the traditional and proposed IPW estimators, with an estimated 3.3-5.0% reduction in telomere length. For the remaining methods, including the proposed outcome modeling estimator, we failed to reject the null hypothesis. Further, there were no appreciable differences in these estimates or the corresponding conclusions regarding significance when comparing our results with versus without including self-reported race in the propensity model (where applicable).
Table 2: Characteristics of the $n=5,270$ non-Hispanic Black and non-Hispanic White National Health and Nutrition Examination Survey (NHANES) participants in our analytic sample; overall and stratified by self-reported race. Descriptive statistics are reported as median (IQR) for continuous measures or number (%) for categorical measures. $P$-values denote the univariate tests of significance (Wilcoxon rank sum test, Pearson’s Chi-squared test, or Fisher’s exact test) comparing the distribution of each measure among non-Hispanic Black versus non-Hispanic White participants.
Characteristic Overall, N = 5,2981 Non-Hispanic White, N = 3,9651 Non-Hispanic Black, N = 1,3331 p-value2
Telomere Length, Mean T/S 1.00 (0.84, 1.18) 0.98 (0.83, 1.16) 1.05 (0.88, 1.25) $<$0.001
Age, Years 50 (35, 67) 52 (35, 70) 45 (34, 62) $<$0.001
Sex 0.4
   Male 2,574 (49%) 1,939 (49%) 635 (48%)
   Female 2,724 (51%) 2,026 (51%) 698 (52%)
Education
   High School or GED 2,621 (49%) 1,790 (45%) 831 (62%)
   Some College 1,448 (27%) 1,102 (28%) 346 (26%)
   College Graduate 1,222 (23%) 1,068 (27%) 154 (12%)
   Refused/Unknown 7 (0.1%) 5 (0.1%) 2 (0.2%)
Marital Status $<$0.001
   Never Married 754 (14%) 436 (11%) 318 (24%)
   Widowed/Divorced/Separated 1,157 (22%) 775 (20%) 382 (29%)
   Married/Living with Partner 3,144 (59%) 2,574 (65%) 570 (43%)
   Refused/Unknown 243 (4.6%) 180 (4.5%) 63 (4.7%)
Household Size $<$0.001
   1 Person 882 (17%) 663 (17%) 219 (16%)
   2 People 1,903 (36%) 1,587 (40%) 316 (24%)
   3 People 997 (19%) 720 (18%) 277 (21%)
   4 People 775 (15%) 551 (14%) 224 (17%)
   5+ People 741 (14%) 444 (11%) 297 (22%)
Home Ownership $<$0.001
   Does Not Own Home 1,551 (29%) 924 (23%) 627 (47%)
   Owns Home 3,706 (70%) 3,011 (76%) 695 (52%)
   Refused/Unknown 41 (0.8%) 30 (0.8%) 11 (0.8%)
Home Type $<$0.001
   Family Home Detached 3,361 (63%) 2,681 (68%) 680 (51%)
   Family Home Attached 464 (8.8%) 332 (8.4%) 132 (9.9%)
   Apartment 938 (18%) 527 (13%) 411 (31%)
   Other 493 (9.3%) 395 (10.0%) 98 (7.4%)
   Refused/Unknown 42 (0.8%) 30 (0.8%) 12 (0.9%)
Household Income $<$0.001
   \$0 - \$20,000 1,076 (20%) 715 (18%) 361 (27%)
   \$20,000 - \$35,000 971 (18%) 682 (17%) 289 (22%)
   \$35,000 - \$55,000 935 (18%) 723 (18%) 212 (16%)
   \$55,000 - \$75,000 609 (11%) 506 (13%) 103 (7.7%)
   \$75,000+ 1,140 (22%) 979 (25%) 161 (12%)
   Refused/Unknown 567 (11%) 360 (9.1%) 207 (16%)
Poverty-Income Ratio $<$0.001
   $<$ 1.3 1,055 (20%) 643 (16%) 412 (31%)
   1.3 $\leq$ PIR $<$ 3.5 1,841 (35%) 1,348 (34%) 493 (37%)
   $\geq$ 3.5 1,954 (37%) 1,672 (42%) 282 (21%)
   Refused/Unknown 448 (8.5%) 302 (7.6%) 146 (11%)
Employment Status  0.001
   Full-Time 932 (18%) 677 (17%) 255 (19%)
   Part-Time 272 (5.1%) 206 (5.2%) 66 (5.0%)
   Retired 602 (11%) 485 (12%) 117 (8.8%)
   Not Working 398 (7.5%) 262 (6.6%) 136 (10%)
   Refused/Unknown 3,094 (58%) 2,335 (59%) 759 (57%)
Occupation Category $<$0.001
   No Work 53 (1.0%) 30 (0.8%) 23 (1.7%)
   Low Blue Collar 690 (13%) 451 (11%) 239 (18%)
   High Blue Collar 250 (4.7%) 196 (4.9%) 54 (4.1%)
   Low White Collar 490 (9.2%) 362 (9.1%) 128 (9.6%)
   High White Collar 644 (12%) 543 (14%) 101 (7.6%)
   Refused/Unknown 3,171 (60%) 2,383 (60%) 788 (59%)
Insurance Status $<$0.001
   Not Insured 681 (13%) 430 (11%) 251 (19%)
   Insured 4,569 (86%) 3,503 (88%) 1,066 (80%)
   Refused/Unknown 48 (0.9%) 32 (0.8%) 16 (1.2%)
Food Security Status $<$0.001
   Food Secure 4,397 (83%) 3,443 (87%) 954 (72%)
   Marginally Food Secure 272 (5.1%) 144 (3.6%) 128 (9.6%)
   Food Insecure 428 (8.1%) 238 (6.0%) 190 (14%)
   Refused/Unknown 201 (3.8%) 140 (3.5%) 61 (4.6%)
WIC Utilization $<$0.001
   No WIC 4,919 (93%) 3,775 (95%) 1,144 (86%)
   Received WIC 366 (6.9%) 185 (4.7%) 181 (14%)
   Refused/Unknown 13 (0.2%) 5 (0.1%) 8 (0.6%)
White Blood Cell Count, SI 6.90 (5.60, 8.40) 7.10 (5.90, 8.60) 6.30 (5.10, 7.60) $<$0.001
   Unknown 2 1 1
Lymphocytes, % 29 (23, 35) 28 (23, 33) 34 (28, 41) $<$0.001
   Unknown 28 19 9
Monocytes, % 8.10 (6.80, 9.58) 8.00 (6.80, 9.40) 8.40 (6.90, 9.90) $<$0.001
   Unknown 28 19 9
Segmented Neutrophils, % 59 (53, 65) 61 (55, 66) 54 (46, 61) $<$0.001
   Unknown 28 19 9
Eosinophils, % 2.30 (1.50, 3.50) 2.30 (1.50, 3.50) 2.30 (1.40, 3.50) 0.14
   Unknown 28 19 9
Basophils, % 0.60 (0.40, 0.80) 0.60 (0.40, 0.90) 0.60 (0.40, 0.80) 0.2
   Unknown 28 19 9
1Median (IQR); n (%)
2Wilcoxon rank sum test; Pearson’s Chi-squared test; Fisher’s exact test
Table 3: Average controlled difference (ACD) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of race (non-Hispanic Black versus non-Hispanic White participants) on log-transformed telomere length among $n=5,270$ participants in the National Health and Nutrition Examination Survey (NHANES). IPTW: Inverse Probability of Treatment Weighting.
Method ACD Estimate 95% Confidence Interval
Multiple Regression 0.0265 0.0106, 0.0423
IPTW Estimator 0.0263 0.0080, 0.0446
Survey-Weighted Multiple Regression 0.0298 -0.0012, 0.0607
IPTW Multiple Regression 0.0181 0.0052, 0.0311
IPTW + Survey-Weighted Multiple Regression 0.0219 -0.0090, 0.0527
Weighted IPTW + Survey-Weighted Multiple Regression 0.0186 -0.0127, 0.0499
Outcome Modeling and Direct Standardization 0.0176 -0.0030, 0.0381
Inverse Probability Weighting 1 0.0150 -0.0151, 0.0451
Inverse Probability Weighting 2 0.0132 -0.0081, 0.0345
6 Discussion
There is a growing literature considering the use of propensity score weighting techniques in complex survey design settings. However, a consensus has not been reached for the setting we consider in this manuscript. Notably, propensity score methods have been developed and used primarily for simple random samples. When analyzing results from more complicated sampling schemes, such as the NHANES clustered, stratified multi-stage sample, one must make the distinction between the population and the sample as the target of inference (Heckman and Vytlacil, 2001). This distinction between the target populations and target estimands are important as researchers must be clear in their inferential statements at both the propensity score model and outcome model stages in their analysis. Further, this article examines a particular consideration for drawing inference in which sample selection depends on the group variable of interest, motivated from a social epidemiology question regarding racial disparity and socioeconomic confounding in health outcomes. Our analytic and numeric results suggest that, in settings where our assumed structure among the variables holds, our proposed methods minimize the bias in estimating the target average controlled difference and achieve nominal coverage rates, as compared to standard analysis methods.
We believe that the context of studying racial disparities presents these particularities in such a way that should be rigorously studied for best practice recommendations, as racial disparities are important and of increasing relevance in scientific research. Moreover the assumed notion of ‘race’ itself warrants discussion. If the scientific question under study seeks to understand the biological underpinnings of a so-called ‘race effect,’ say due to differences in genetic ancestry, then we cannot posit a causal counterfactual framework, as ‘race’ cannot be manipulated or thought of as a ‘cause.’ However, if race is measured as a social construct, we can then discuss altering racialization as a lived experience (Marcellesi, 2013). In our motivating example, we are interested in studying differences in telomere length between non-Hispanic Black and non-Hispanic White individuals. We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for available socioeconomic status variables in NHANES and after utilizing appropriate propensity score and survey weighting techniques. Among our proposed methods, we estimated the ACD in telomere length to be between 1.32% and 1.76%, though we failed to detect a statistically significant difference. This contrasts with the results from other potential approaches, which varied from 1.81% to 3.0% in this study. Needham et al. (2013) estimated 4.8% longer telomeres in Black individuals, on average, adjusting for education, income, health behavior, and body mass index. Needham et al. (2020) and Roberts et al. (2022) also suggested evidence of social and environmental factors playing a role in racial disparities in telomere length. Taken together, our proposed method contributes to the collection of strategies that can investigate the role of social determinants of health on disparities, and in particular through the use of nationally representative survey data.
Though we focus on a particular research question and methods for propensity scores within the frame of complex survey designs, these concepts readily extend to other settings where observational data are collected via an underlying sample selection mechanism. More broadly, this type of problem is important to study to ensure research does not exacerbate health inequities through inadequate consideration of confounding and biases due to non-representative selection into databases. In particular, other large-scale databases, such as electronic medical records, provide valuable insights into real-world data but may be subject to their own selection biases. Areas of interest for future work include electronic health records data with unknown sample selection probabilities, non-Gaussian outcomes, and more complex questions that may include affect heterogeneity. Several conceptual questions about the in-sample or population generalizable propensity score have not yet been reconciled in the literature. Further, causal principles related to this work may be extended to more complex relational diagrams or sampling designs. For instance, as certain biomedical data become inexpensive and easy to collect, but other endpoints are not, a natural question is whether this framework can be extended to two-stage sampling or sequential designs. Lastly, we can formally expand on the utility of the proposed method by proposing a doubly robust, augmented inverse probability weighted estimator.
In conclusion, this work outlines approaches and novel estimands for using propensity score and survey weights in tandem to study controlled outcome differences when the groups under comparison influence sample selection. Though the conceptual diagram for this study is rather specific, this setup is common to many settings of health disparities research. That is, our proposed methods and investigation to estimate the ACD follow a motivating health disparities research question in which the sampling scheme and corresponding survey weights depend on the target of the propensity score model. We fill this gap in the literature by providing recommendation on the form of the appropriate estimators in this setting and provide analytic and numeric evidence of their properties. We have shown that when sample selection depends on the comparison groups of interest and/or potential confounders, we must use the direct standardization or propensity weighting strategies outlined here, rather than standard analytic approaches. It is our hope that this work will not only further our understanding of appropriate survey design considerations in this framework, but also make the task of analyzing such data in similar settings more accessible to researchers.
Acknowledgments
We thank Dr. Yajuan Si for helpful discussions on this topic. This work was supported in part by NIH/NIGMS R35 GM144128 and the Fred Hutchinson Cancer Center J. Orin Edson Foundation Endowed Chair (SS), NIH/NIMH DP2MH122405, R01HD107015, and P2C HD042828 (THM), NSF DMS 1712933 and NIH/NCI UG3-CA267907 (BM), and NIH/NIGMS R01GM139926 (XS).
Appendix A Additional Analytic Results
A.1 Average Controlled Difference
The assumptions required to estimate the averaged controlled difference (ACD) are relatively weaker than those required to identify the causal potential outcome means (see below). In particular, the following assumptions are sufficient for showing that our functionals (2), (3), and (4) in the main text estimate the ACD:
  1. (i)
    Positivity: $\Pr(A=a\mid X=x)>0\ \forall a\in\mathcal{A}$ and every $x$ such that $f_{X}(x)>0$
  2. (ii)
    Positivity of Sample Selection: $\Pr(S=1\mid A=a,X=x)>0$ for every $a,x$ such that $f_{A,X}(a,x)>0$
  3. (iii)
    Mean (Weak) Generalizability (Selection Exchangeability): $\mathbb{E}[Y\mid A=a,X]=\mathbb{E}[Y\mid A=a,S=1,X]$, that is, selection bias comes from $A,X$. This is implied by, but does not imply, strong generalizability (in distribution), $Y\perp S\mid A,X$.
We proceed by deriving the estimators in Equations (2), (3), and (4) from the main text below. We define the ACD as
${\rm ACD}=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=1,X\right]-\mathbb{E}\left[Y\mid A=0,X\right]\right].$
Outcome Modeling and Direct Standardization
Under assumptions (i) - (iii), we show that the observed data functional, $\mu(a)$, given by
$\mu(a)=\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right],$
can be used to estimate the ACD. Assumptions (i) and (ii) come in to ensure that $\mu(a)$ is well defined. Focusing on $\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$, we have that
$\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$ =$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
[…]
$=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right]$ (Main Eq. 2)
Inverse Probability Weighting
Under assumptions (i) - (iii), we further show that the observed data functionals for our inverse probability weighted estimators, given by
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$
also estimate $\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$. Let $C=A\times S=\mathbb{I}(A=1,S=1)$, that is,
$C=\begin{cases}1;&A=a,S=1\\ 0;&S=0\text{ or }A\neq a,S=1\end{cases}$
so that
$\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$ =$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
$=\mathbb{E}_{X}\left[\frac{C\times Y}{\Pr(C=1\mid X)}\right];\quad\text{following the traditional IPW proof}$
$=\mathbb{E}_{X}\left[\frac{A\times S\times Y}{\Pr(A=a,S=1\mid X)}\right];\quad\text{by definition of }C$
$=\mathbb{E}_{X}\left[\frac{A\times 1\times Y}{\Pr(A=a,S=1\mid X)}\mid S=1\right]\times\Pr(S=1)+0$
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a,S=1\mid X)}\mid S=1\right]$
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$ (Main Eq. 3)
or
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$ (Main Eq. 4)
where again, the observed data functionals are well defined under (i) and (ii).
A.2 Identification of the Potential Outcome Means
Though not applicable to our motivating research question, in many scenarios the group membership variable of interest is a ‘treatment’ or ‘exposure’ in the more traditional sense, and we seek to target a causal quantity. To that end, we can adopt a counterfactual framework and, under stronger assumptions, show that the proposed estimators equivalently identify such a parameter. We introduce some additional notation below.
Take $Y^{a}$ to be the random variable representing the potential counterfactual outcome within each group (Rubin, 2005). That is, despite only observing one outcome per individual, we know that each study participant had two potential outcomes – the outcome that would be observed if the individual was in comparison group $a=0$, denoted by $Y^{0}$, and the potential outcome that would be observed if the participant instead was in comparison group $a=1$, denoted by $Y^{1}$. By considering an individual’s potential outcome values based on their observed and counterfactual group membership, an inherent missing data problem arises. Importantly, there are two layers to the missingness mechanism in our setup. Firstly, we observe no data when $S=0$. Secondly, due to the missingness of counterfactual values, we cannot fully observe all potential outcomes, $Y^{a}$. Further, as before, $\mathbb{E}\left[Y^{a}\right]\neq\mathbb{E}\left[Y^{a}\mid S=1\right]$ in general, as the sampled study participants may not be representative of the larger target population. We can, however, make assumptions to estimate the unobserved, counterfactual outcome in the population. With the assumptions of the previous section, the following additional conditions are sufficient to identify the population potential outcome means, $\mathbb{E}[Y^{a}]\ \forall a\in\mathcal{A}$:
  1. (iii*)
    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}[Y^{a}\mid A=a,X]=\mathbb{E}[Y^{a}\mid A=a,S=1,X]$, that is, selection bias comes from $A,X$. This is implied by, but does not imply, strong generalizability (in distribution), $Y^{a}\perp S\mid A,X$.
  2. (iv)
    Stable Unit Treatment Value (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
  3. (v)
    Mean (Weak) Exchangeability: $\mathbb{E}[Y^{a}\mid X]=\mathbb{E}[Y^{a}\mid A=a,X]$, that is, confounding bias comes from $X$. This is implied by, but does not imply, strong exchangeability, i.e., $Y^{a}\perp A\mid X$.
This is thematically similar to Dahabreh et al. (2020), who study the generalization of causal inferences from individuals in randomized trials. The authors also differentiate between conditional exchangeability in the mean over $A$ in a randomized population and conditional exchangeability in the mean over $S$ in the context of randomized trials.
Outcome Modeling and Direct Standardization
Under assumptions (i) - (v), we show that the observed data functional, $\mu(a)$, identifies the potential outcome mean in the target population, $\mathbb{E}[Y^{a}]$, where
$\mu(a)=\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right].$
As before, $\mu(a)$ is well defined under (i) and (ii). The SUTVA, or consistency, assumption underpins the existence of the counterfactual outcomes, $Y^{a}$. The compact representation in (iv) captures two aspects of SUTVA – ‘no interference,’ which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and ‘no hidden variation,’ which states that the potential outcome is well-defined for each treatment level, $a$. To identify $\mathbb{E}\left[Y^{a}\right]$, we have that
$\mathbb{E}\left[Y^{a}\right]$ $=\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid X\right]\right];\quad\text{by law of total expectation}$
=$\stackrel{{\scriptstyle(v)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,X\right]\right]$
=$\stackrel{{\scriptstyle(iii*)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,S=1,X\right]\right]$
=$\stackrel{{\scriptstyle(iv)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
[…]
$=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right]$ (Main Eq. 2)
Inverse Probability Weighting
Under assumptions (i) - (v), we further show that the observed data functionals for our inverse probability weighted estimators, given by
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$
also identify the potential outcome mean in the target population, $\mathbb{E}[Y^{a}]$. Let $C=A\times S=\mathbb{I}(A=1,S=1)$, that is,
$C=\begin{cases}1;&A=a,S=1\\ 0;&S=0\text{ or }A\neq a,S=1\end{cases}$
so that
$\mathbb{E}\left[Y^{a}\right]$ $=\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid X\right]\right];\quad\text{by law of total expectation}$
=$\stackrel{{\scriptstyle(v)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,X\right]\right]$
=$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,S=1,X\right]\right]$
=$\stackrel{{\scriptstyle(iv)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
$=\mathbb{E}_{X}\left[\frac{C\times Y}{\Pr(C=1\mid X)}\right];\quad\text{following the traditional IPW proof}$
$=\mathbb{E}_{X}\left[\frac{A\times S\times Y}{\Pr(A=a,S=1\mid X)}\right];\quad\text{by definition of }C$
$=\mathbb{E}_{X}\left[\frac{A\times 1\times Y}{\Pr(A=a,S=1\mid X)}\mid S=1\right]\times\Pr(S=1)+0$
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a,S=1\mid X)}\mid S=1\right]$
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$ (Main Eq. 3)
or
$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$ (Main Eq. 4)
where again, the observed data functionals are well defined under (i) and (ii).
Appendix B Additional Simulation Results
Table B1: Simulation results comparing different analytic methods for calculating the population average controlled difference (ACD) for $A$ on $Y$ across different data generating mechanism settings. Results presented are the estimated ACD, percent bias, analytic standard error (SE), Monte Carlo standard error, mean squared error (MSE) and coverage probability over 200 subsamples for each setting
Method True ACD Estimated ACD Percent Bias Analytic SE Monte Carlo SE MSE Coverage
Setting 1: $\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=0$
Oracle Estimator 1.096 1.101 0.471 0.068 0.069 0.477 0.925
Simple Regression 1.096 1.098 0.166 0.097 0.105 1.103 0.930
Multiple Regression 1.096 1.101 0.450 0.068 0.070 0.484 0.915
IPTW Estimator 1.096 1.101 0.475 0.068 0.070 0.484 0.920
Survey-Weighted Multiple Regression 1.096 1.102 0.536 0.068 0.069 0.481 0.930
IPTW Multiple Regression 1.096 1.101 0.453 0.060 0.070 0.484 0.880
IPTW + Survey-Weighted Multiple Regression 1.096 1.102 0.538 0.060 0.069 0.482 0.875
Weighted IPTW + Survey-Weighted Multiple Regression 1.096 1.102 0.538 0.060 0.069 0.482 0.875
Naive G-Computation 1.096 0.992 -9.478 0.030 0.102 2.111 0.290
Outcome Modeling and Direct Standardization 1.096 1.090 -0.523 0.067 0.069 0.476 0.920
Inverse Probability Weighting 1 1.096 1.089 -0.647 0.068 0.069 0.480 0.925
Inverse Probability Weighting 2 1.096 1.091 -0.499 0.067 0.069 0.475 0.920
Setting 2: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=0$
Oracle Estimator 1.090 1.079 -0.989 0.066 0.067 0.452 0.930
Simple Regression 1.090 1.957 79.480 0.083 0.076 75.638 0.000
Multiple Regression 1.090 1.079 -0.977 0.066 0.067 0.455 0.925
IPTW Estimator 1.090 1.071 -1.793 0.094 0.107 1.176 0.930
Survey-Weighted Multiple Regression 1.090 1.080 -0.886 0.066 0.068 0.468 0.925
IPTW Multiple Regression 1.090 1.077 -1.246 0.060 0.069 0.490 0.885
IPTW + Survey-Weighted Multiple Regression 1.090 1.078 -1.131 0.060 0.070 0.506 0.890
Weighted IPTW + Survey-Weighted Multiple Regression 1.090 1.078 -1.132 0.060 0.070 0.506 0.885
Naive G-Computation 1.090 0.969 -11.125 0.029 0.089 2.268 0.180
Outcome Modeling and Direct Standardization 1.090 1.069 -1.935 0.065 0.066 0.480 0.930
Inverse Probability Weighting 1 1.090 1.060 -2.778 0.093 0.106 1.220 0.925
Inverse Probability Weighting 2 1.090 1.060 -2.742 0.093 0.106 1.204 0.925
Setting 3: $\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=0$
Oracle Estimator 1.100 1.096 -0.422 0.050 0.047 0.220 0.960
Simple Regression 1.100 1.099 -0.138 0.073 0.074 0.547 0.945
Multiple Regression 1.100 1.096 -0.418 0.050 0.047 0.223 0.965
IPTW Estimator 1.100 1.096 -0.421 0.050 0.047 0.223 0.965
Survey-Weighted Multiple Regression 1.100 1.096 -0.417 0.056 0.047 0.222 0.990
IPTW Multiple Regression 1.100 1.096 -0.417 0.050 0.047 0.223 0.965
IPTW + Survey-Weighted Multiple Regression 1.100 1.096 -0.416 0.056 0.047 0.222 0.990
Weighted IPTW + Survey-Weighted Multiple Regression 1.100 1.096 -0.417 0.050 0.047 0.222 0.970
Naive G-Computation 1.100 0.998 -9.341 0.025 0.073 1.589 0.230
Outcome Modeling and Direct Standardization 1.100 1.084 -1.495 0.051 0.047 0.242 0.955
Inverse Probability Weighting 1 1.100 1.098 -0.217 0.051 0.047 0.219 0.965
Inverse Probability Weighting 2 1.100 1.084 -1.497 0.051 0.047 0.243 0.955
Setting 4: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$
Oracle Estimator 1.106 1.102 -0.419 0.055 0.058 0.342 0.905
Simple Regression 1.106 1.972 78.249 0.070 0.069 75.408 0.000
Multiple Regression 1.106 1.121 1.377 0.058 0.062 0.404 0.915
IPTW Estimator 1.106 1.123 1.487 0.103 0.105 1.129 0.940
Survey-Weighted Multiple Regression 1.106 1.123 1.499 0.050 0.062 0.411 0.890
IPTW Multiple Regression 1.106 1.116 0.923 0.044 0.065 0.426 0.830
IPTW + Survey-Weighted Multiple Regression 1.106 1.118 1.045 0.050 0.065 0.430 0.870
Weighted IPTW + Survey-Weighted Multiple Regression 1.106 1.119 1.116 0.045 0.065 0.434 0.820
Naive G-Computation 1.106 0.997 -9.867 0.022 0.070 1.680 0.175
Outcome Modeling and Direct Standardization 1.106 1.090 -1.442 0.055 0.058 0.362 0.925
Inverse Probability Weighting 1 1.106 1.104 -0.210 0.076 0.081 0.648 0.940
Inverse Probability Weighting 2 1.106 1.089 -1.526 0.081 0.083 0.706 0.945
Setting 5: $\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=1$
Oracle Estimator 1.104 1.108 0.389 0.046 0.048 0.227 0.940
Simple Regression 1.104 1.209 9.544 0.048 0.045 1.311 0.365
Multiple Regression 1.104 1.205 9.214 0.034 0.032 1.135 0.110
IPTW Estimator 1.104 1.205 9.222 0.034 0.032 1.136 0.110
Survey-Weighted Multiple Regression 1.104 1.205 9.236 0.046 0.045 1.239 0.355
IPTW Multiple Regression 1.104 1.205 9.214 0.030 0.032 1.135 0.090
IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.209 0.041 0.045 1.235 0.270
Weighted IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.182 0.040 0.045 1.225 0.265
Naive G-Computation 1.104 1.003 -9.119 0.015 0.074 1.559 0.135
Outcome Modeling and Direct Standardization 1.104 1.090 -1.252 0.048 0.047 0.236 0.935
Inverse Probability Weighting 1 1.104 1.100 -0.302 0.057 0.054 0.291 0.945
Inverse Probability Weighting 2 1.104 1.099 -0.448 0.058 0.057 0.325 0.955
Setting 6: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$
Oracle Estimator 1.096 1.096 -0.018 0.041 0.042 0.176 0.935
Simple Regression 1.096 2.004 82.816 0.044 0.043 82.577 0.000
Multiple Regression 1.096 1.184 8.045 0.036 0.036 0.904 0.310
IPTW Estimator 1.096 1.222 11.486 0.058 0.054 1.874 0.440
Survey-Weighted Multiple Regression 1.096 1.189 8.485 0.044 0.059 1.207 0.425
IPTW Multiple Regression 1.096 1.185 8.115 0.030 0.037 0.927 0.220
IPTW + Survey-Weighted Multiple Regression 1.096 1.182 7.881 0.040 0.048 0.972 0.440
Weighted IPTW + Survey-Weighted Multiple Regression 1.096 1.182 7.847 0.040 0.047 0.958 0.435
Naive G-Computation 1.096 0.999 -8.879 0.015 0.065 1.370 0.130
Outcome Modeling and Direct Standardization 1.096 1.079 -1.514 0.044 0.041 0.197 0.950
Inverse Probability Weighting 1 1.096 1.083 -1.209 0.066 0.072 0.540 0.940
Inverse Probability Weighting 2 1.096 1.089 -0.633 0.066 0.070 0.490 0.935
Setting 7: $\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=1$
Oracle Estimator 1.102 1.102 0.073 0.035 0.034 0.115 0.975
Simple Regression 1.102 1.058 -3.992 0.036 0.034 0.309 0.805
Multiple Regression 1.102 1.175 6.701 0.026 0.025 0.605 0.190
IPTW Estimator 1.102 1.174 6.580 0.026 0.025 0.587 0.210
Survey-Weighted Multiple Regression 1.102 1.177 6.809 0.038 0.031 0.661 0.510
IPTW Multiple Regression 1.102 1.175 6.698 0.026 0.025 0.605 0.190
IPTW + Survey-Weighted Multiple Regression 1.102 1.176 6.793 0.037 0.032 0.663 0.455
Weighted IPTW + Survey-Weighted Multiple Regression 1.102 1.177 6.808 0.034 0.031 0.658 0.385
Naive G-Computation 1.102 1.017 -7.688 0.013 0.054 1.009 0.135
Outcome Modeling and Direct Standardization 1.102 1.086 -1.405 0.038 0.033 0.135 0.970
Inverse Probability Weighting 1 1.102 1.100 -0.172 0.043 0.042 0.178 0.970
Inverse Probability Weighting 2 1.102 1.076 -2.322 0.044 0.045 0.265 0.910
Setting 8: $\tau_{X}=1$, $\beta_{A}=1$, $\beta_{X}=1$
Oracle Estimator 1.104 1.101 -0.209 0.034 0.033 0.106 0.970
Simple Regression 1.104 1.919 73.914 0.040 0.037 66.681 0.000
Multiple Regression 1.104 1.210 9.640 0.034 0.032 1.234 0.130
IPTW Estimator 1.104 1.172 6.226 0.070 0.066 0.904 0.835
Survey-Weighted Multiple Regression 1.104 1.200 8.727 0.033 0.063 1.321 0.300
IPTW Multiple Regression 1.104 1.212 9.830 0.022 0.031 1.273 0.015
IPTW + Survey-Weighted Multiple Regression 1.104 1.207 9.378 0.031 0.034 1.188 0.130
Weighted IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.214 0.030 0.041 1.204 0.160
Naive G-Computation 1.104 0.986 -10.660 0.011 0.053 1.669 0.040
Outcome Modeling and Direct Standardization 1.104 1.087 -1.531 0.038 0.032 0.131 0.970
Inverse Probability Weighting 1 1.104 1.115 1.015 0.050 0.051 0.274 0.950
Inverse Probability Weighting 2 1.104 1.083 -1.907 0.047 0.047 0.267 0.935
Table B2: Sensitivity analysis simulation results comparing the proposed methods to the oracle estimator in terms of bias, mean squared error, and coverage across different treatment heterogeneity levels (size of $\gamma_{AX}$ in the true data generating mechanism. Results are averaged over 200 independent sub-samplings from a super population of 100,000 individuals where all other simulation settings match that of Setting 8 in the main text. Results for OM are presented where the outcome model is correctly specified versus misspecified. ACD: Average Controlled Difference; SE: Standard Error; MSE: Mean Squared Error.
Method True ACD Estimated ACD Bias Analytic SE Monte Carlo SE MSE Coverage
Setting 1: $\gamma_{AX}=0$
Oracle Estimator $0.996$ $0.985$ $-0.011$ $0.034$ $0.035$ $0.001$ $0.920$
Outcome Modeling and Direct Standardization $0.996$ $0.971$ $-0.026$ $0.038$ $0.035$ $0.002$ $0.885$
Outcome Modeling and Direct Standardization, Misspecified $0.996$ $0.968$ $-0.028$ $0.034$ $0.030$ $0.002$ $0.900$
Inverse Probability Weighting 1 $0.996$ $0.983$ $-0.014$ $0.051$ $0.052$ $0.003$ $0.950$
Inverse Probability Weighting 2 $0.996$ $0.956$ $-0.041$ $0.046$ $0.045$ $0.004$ $0.835$
Setting 2: $\gamma_{AX}=0.01$
Oracle Estimator $1.000$ $0.991$ $-0.009$ $0.034$ $0.033$ $0.001$ $0.960$
Outcome Modeling and Direct Standardization $1.000$ $0.976$ $-0.024$ $0.038$ $0.032$ $0.002$ $0.940$
Outcome Modeling and Direct Standardization, Misspecified $1.000$ $0.983$ $-0.017$ $0.034$ $0.028$ $0.001$ $0.970$
Inverse Probability Weighting 1 $1.000$ $0.992$ $-0.008$ $0.050$ $0.052$ $0.003$ $0.935$
Inverse Probability Weighting 2 $1.000$ $0.955$ $-0.045$ $0.047$ $0.050$ $0.004$ $0.825$
Setting 3: $\gamma_{AX}=0.05$
Oracle Estimator $1.050$ $1.051$ $0.001$ $0.034$ $0.030$ $0.001$ $0.970$
Outcome Modeling and Direct Standardization $1.050$ $1.035$ $-0.015$ $0.038$ $0.029$ $0.001$ $0.965$
Outcome Modeling and Direct Standardization, Misspecified $1.050$ $1.059$ $0.008$ $0.035$ $0.026$ $0.001$ $0.985$
Inverse Probability Weighting 1 $1.050$ $1.041$ $-0.009$ $0.051$ $0.048$ $0.002$ $0.955$
Inverse Probability Weighting 2 $1.050$ $1.012$ $-0.037$ $0.046$ $0.043$ $0.003$ $0.870$
Setting 4: $\gamma_{AX}=0.1$
Oracle Estimator $1.106$ $1.098$ $-0.008$ $0.034$ $0.034$ $0.001$ $0.950$
Outcome Modeling and Direct Standardization $1.106$ $1.082$ $-0.022$ $0.038$ $0.033$ $0.002$ $0.925$
Outcome Modeling and Direct Standardization, Misspecified $1.106$ $1.130$ $0.021$ $0.036$ $0.030$ $0.001$ $0.930$
Inverse Probability Weighting 1 $1.106$ $1.102$ $-0.003$ $0.050$ $0.047$ $0.002$ $0.965$
Inverse Probability Weighting 2 $1.106$ $1.070$ $-0.033$ $0.047$ $0.044$ $0.003$ $0.880$
Setting 5: $\gamma_{AX}=0.5$
Oracle Estimator $1.501$ $1.489$ $-0.008$ $0.035$ $0.031$ $0.001$ $0.960$
Outcome Modeling and Direct Standardization $1.501$ $1.470$ $-0.020$ $0.037$ $0.034$ $0.002$ $0.895$
Outcome Modeling and Direct Standardization, Misspecified $1.501$ $1.720$ $0.146$ $0.041$ $0.029$ $0.049$ $0.000$
Inverse Probability Weighting 1 $1.501$ $1.497$ $-0.002$ $0.050$ $0.053$ $0.003$ $0.925$
Inverse Probability Weighting 2 $1.501$ $1.467$ $-0.023$ $0.048$ $0.047$ $0.003$ $0.900$
Appendix C Additional Data Example Results
C.1 Main Analysis of Race, SES, and Telomere Length
Table C1: Results from our (a) unweighted and (b) survey-weighted ‘propensity for race’ models. Propensity scores were constructed by fitting a logistic regression of self-reported race (Black = 1, White = 0) on twelve socioeconomic covariates. Results presented are the odds ratios (OR), corresponding 95% confidence intervals (95% CI) and $p$-values. Em-dashes (—) denote the reference group.
a. Unweighted b. Survey-Weighted
Characteristic OR1 95% CI1 p-value OR1 95% CI1 p-value
Education
   High School or GED
   Some College 0.75 0.64, 0.89 $<$0.001 0.80 0.67, 0.96 0.016
   College Graduate 0.49 0.39, 0.61 $<$0.001 0.49 0.38, 0.62 $<$0.001
   Refused/Unknown 0.96 0.12, 5.42 $>$0.9 0.29 0.02, 3.60 0.3
Marital Status
   Never Married
   Widowed/Divorced/Separated 0.93 0.75, 1.16 0.5 0.88 0.69, 1.12 0.3
   Married/Living with Partner 0.43 0.35, 0.52 $<$0.001 0.41 0.33, 0.50 $<$0.001
   Refused/Unknown 0.70 0.49, 1.01 0.057 0.82 0.54, 1.23 0.3
Household Size
   1 Person
   2 People 1.41 1.11, 1.80 0.006 1.33 1.00, 1.77 0.049
   3 People 2.55 1.95, 3.33 $<$0.001 2.27 1.66, 3.10 $<$0.001
   4 People 2.51 1.88, 3.36 $<$0.001 2.18 1.57, 3.03 $<$0.001
   5+ People 3.90 2.87, 5.29 $<$0.001 3.51 2.49, 4.94 $<$0.001
Household Income
   0 - \$20,000
   \$20,000 - \$35,000 1.13 0.88, 1.45 0.3 1.04 0.77, 1.41 0.8
   \$35,000 - \$55,000 0.81 0.60, 1.10 0.2 0.72 0.51, 1.02 0.065
   \$55,000 - \$75,000 0.70 0.49, 1.02 0.062 0.64 0.42, 0.98 0.041
   \$75,000+ 0.67 0.45, 0.98 0.040 0.65 0.40, 1.04 0.069
   Refused/Unknown 1.89 1.22, 2.94 0.005 1.83 1.10, 3.04 0.020
Poverty-Income Ratio Category
   $<$ 1.3
   1.3 $\leq$ PIR $<$ 3.5 0.96 0.76, 1.21 0.7 0.99 0.75, 1.30 $>$0.9
   $\geq$ 3.5 0.90 0.64, 1.26 0.5 0.83 0.56, 1.24 0.4
   Refused/Unknown 0.65 0.40, 1.05 0.078 0.60 0.35, 1.03 0.062
Employment Status
   Full-Time
   Part-Time 0.77 0.54, 1.08 0.13 0.80 0.55, 1.16 0.2
   Retired 0.73 0.55, 0.97 0.034 0.74 0.54, 1.01 0.058
   Not Working 0.84 0.62, 1.13 0.3 0.90 0.64, 1.28 0.6
   Refused/Unknown 0.68 0.45, 1.04 0.070 1.01 0.63, 1.62 $>$0.9
Occupation Category
   No Work
   Low Blue Collar 1.11 0.57, 2.18 0.8 1.37 0.66, 2.84 0.4
   High Blue Collar 0.73 0.35, 1.53 0.4 0.90 0.41, 2.00 0.8
   Low White Collar 0.87 0.44, 1.76 0.7 1.29 0.61, 2.74 0.5
   High White Collar 0.73 0.36, 1.49 0.4 1.04 0.48, 2.26 $>$0.9
   Refused/Unknown 1.12 0.53, 2.40 0.8 1.26 0.55, 2.86 0.6
WIC Utilization
   No WIC
   Received WIC 1.20 0.93, 1.55 0.2 1.24 0.92, 1.68 0.2
   Refused/Unknown 2.39 0.73, 8.75 0.2 1.92 0.56, 6.59 0.3
Food Security Status
   Food Secure
   Marginally Food Secure 1.80 1.36, 2.38 $<$0.001 2.18 1.59, 2.99 $<$0.001
   Food Insecure 1.40 1.10, 1.79 0.006 1.46 1.12, 1.92 0.006
   Refused/Unknown 1.58 1.07, 2.31 0.020 1.72 1.10, 2.69 0.017
Home Type
   Family Home Detached
   Family Home Attached 1.46 1.14, 1.86 0.002 1.55 1.19, 2.01 0.001
   Apartment 2.23 1.78, 2.81 $<$0.001 2.35 1.82, 3.04 $<$0.001
   Other 0.68 0.53, 0.88 0.004 0.85 0.65, 1.11 0.2
   Refused/Unknown 1.45 0.18, 13.0 0.7 2.39 0.00, 2,519 0.8
Home Ownership
   Does Not Own Home
   Owns Home 0.79 0.65, 0.97 0.023 0.77 0.63, 0.95 0.017
   Refused/Unknown 0.14 0.01, 1.23 0.090 0.04 0.00, 128 0.4
Insurance Status
   Not Insured
   Insured 1.03 0.84, 1.25 0.8 1.08 0.87, 1.34 0.5
   Refused/Unknown 2.22 0.56, 9.78 0.3 4.21 0.88, 20.2 0.072
1OR = Odds Ratio, CI = Confidence Interval
C.2 Supplemental Analysis of Lead Exposure, SES, and Telomere Length
Table C2: Population Average Treatment Effect (PATE) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of lead exposure (high versus low exposure, dichotomized at median $\mu$g/dL) on log-transformed telomere length among $n=5,267$ participants in the National Health and Nutrition Examination Survey (NHANES). IPTW: Inverse Probability of Treatment Weighting.
Method PATE Estimate 95% Confidence Interval
Multiple Regression $0.0002$ $-0.0139,0.0142$
IPTW Estimator $-0.0478$ $-0.0800,-0.0156$
IPTW Estimator w/ Race in Propensity Model $-0.0504$ $-0.0809,-0.0199$
Survey-Weighted Multiple Regression $-0.0032$ $-0.0229,0.0165$
IPTW Multiple Regression $-0.0024$ $-0.0157,0.0109$
IPTW Multiple Regression w/ Race in Propensity Model $-0.0029$ $-0.0161,0.0103$
IPTW + Survey-Weighted Multiple Regression $-0.0051$ $-0.0250,0.0149$
IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0053$ $-0.0255,0.0149$
Weighted IPTW + Survey-Weighted Multiple Regression $-0.0049$ $-0.0253,0.0155$
Weighted IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0052$ $-0.0257,0.0153$
Proposed OM $0.0001$ $-0.0121,0.0123$
Proposed OM w/ Race in Propensity Model $0.0001$ $-0.0121,0.0123$
Proposed IPW1 $-0.0358$ $-0.0469,-0.0248$
Proposed IPW1 w/ Race in Propensity Model $-0.0368$ $-0.0479,-0.0257$
Proposed IPW2 $-0.0328$ $-0.0433,-0.0223$
Proposed IPW2 w/ Race in Propensity Model $-0.0332$ $-0.0440,-0.0224$

### v2

@@
 Abstract
-A basic descriptive question in statistics often asks whether there are differences in mean outcomes between groups based on levels of a discrete covariate (e.g., racial disparities in health outcomes).
-However, when this categorical covariate of interest is correlated with other factors related to the outcome, direct comparisons may lead to biased estimates and invalid inferential conclusions without appropriate adjustment.
-Propensity score methods are broadly employed with observational data as a tool to achieve covariate balance, but how to implement them in complex surveys is less studied – in particular, when the survey weights depend on the group variable under comparison.
-In this work, we focus on a specific example when sample selection depends on race.
-We propose identification formulas to properly estimate the average controlled difference (ACD) in outcomes between Black and White individuals, with appropriate weighting for covariate imbalance across the two racial groups and generalizability.
+In this work, we are motivated by the problem of estimating racial disparities in health outcomes, specifically the average controlled difference (ACD) in telomere length between Black and White individuals, using data from the National Health and Nutrition Examination Survey (NHANES).
+To do so, we build a propensity for race to properly adjust for other social determinants while characterizing the controlled effect of race on telomere length.
+Propensity score methods are broadly employed with observational data as a tool to achieve covariate balance, but how to implement them in complex surveys is less studied – in particular, when the survey weights depend on the group variable under comparison (as the NHANES sampling scheme depends on self-reported race).
+We propose identification formulas to properly estimate the ACD in outcomes between Black and White individuals, with appropriate weighting for both covariate imbalance across the two racial groups and generalizability.
 Via extensive simulation, we show that our proposed methods outperform traditional analytic approaches in terms of bias, mean squared error, and coverage when estimating the ACD for our setting of interest.
-We are motivated by the interplay between race and social determinants of health when estimating racial differences in telomere length using data from the National Health and Nutrition Examination Survey (NHANES).
-The NHANES sampling scheme depends on self-reported race.
-We build a propensity for race to properly adjust for other social determinants while characterizing the controlled effect of race on telomere length.
-We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for confounding by socioeconomic factors and after utilizing appropriate propensity score and survey weighting techniques.
-Software to implement these methods and code to reproduce our results can be found in the R package svycdiff at https://github.com/salernos/svycdiff.
-Keywords Controlled Outcome Differences  $\cdot$ Propensity Scores  $\cdot$ Complex Surveys  $\cdot$ NHANES  $\cdot$ Racial Disparities
+In our data, we find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for confounding by socioeconomic factors and utilizing appropriate propensity score and survey weighting techniques.
+Software to implement these methods and code to reproduce our results can be found in the R package svycdiff, available through the Comprehensive R Archive Network (CRAN) at cran.r-project.org/web/packages/svycdiff/, or in a development version on GitHub at github.com/salernos/svycdiff.
+Keywords Complex Surveys  $\cdot$ Controlled Outcome Differences  $\cdot$ NHANES  $\cdot$ Propensity Scores  $\cdot$ Racial Disparities
 1 Introduction
-There is a deep and broad literature in the United States (US) documenting racial disparities along multiple socioeconomic axes (e.g., income, education; see Needham et al., 2013; Rehkopf and Needham, 2019; Zota and VanNoy, 2021).
-These dimensions are, of course, not independent, making it difficult to separate the variation in outcomes associated with race from differences attributable to other sources.
-Self-identified race, as measured in many studies, is a complex composite of biology and lived experiences, making it impossible to disentangle the two without additional data (e.g., genetic ancestry).
-Propensity scores are commonly used in causal inference, but they are also applicable under weaker assumptions to study such differences among groups by balancing complex correlates across these groups (Stuart et al., 2011; Lim et al., 2018; Zeng et al., 2021).
-While previous work has shown that propensity scores can be used to study racial disparities (see Li and Li, 2023), a gap remains in determining the ideal practice when the sampling design is complex.
-Specifically, surveys are the backbone of such observational research, but, especially as response rates decline and costs rise, they almost always involve complex designs and non-response patterns.
-Of particular interest is when the design probability of selection into the survey sample depends on the comparison variable of interest, such as race.
-In more traditional causal inference terminology, these sensitive variables may be analogous to ‘treatment’ or ‘exposure’ variables.
-However, unlike in randomized trials, which are often viewed as the gold standard to estimate controlled outcome differences (Zabor et al., 2020), certain observational scenarios preclude randomizing individuals to a particular experimental condition due to moral or biological restrictions.
+There is a deep and broad literature in the United States (US) documenting racial disparities along multiple socioeconomic axes, such as income and education (Needham et al., 2013; Rehkopf and Needham, 2019; Zota and VanNoy, 2021).
+These factors are interrelated, making it difficult to separate the effect of race from other sources.
+Further, these complex relationships between race and socioeconomic status (SES), and the porous collection of such data in the US, make it challenging to draw definitive conclusions about race differences in health outcomes (Li and Li, 2023).
+Observational studies can offer a wealth of information, but they are often limited by confounding, covariate imbalance, and a lack of representation (Imai and Ratkovic, 2013; Pattanayak et al., 2011).
+Nationally representative surveys, using techniques such as stratified and clustered sampling, allow us to study diverse populations, but generalizing results while accounting for confounding can be difficult due to their complex designs.
+Our motivating question focuses on the relationship between race, SES, and telomere length, a biomarker for health, using data from the National Health and Nutrition Examination Survey (NHANES): if we could hypothetically balance SES between Black and White individuals in a nationally representative sample, would we still see significant Black/White differences in telomere length?
+This question is statistically complex because the population characteristic of interest (i.e., race) is correlated with SES, and both factors influence the probability of being selected into the sample.
+For NHANES, the sampling mechanism is partly determined by race and socioeconomic status, with an oversampling of non-Hispanic Black individuals and those at or below 130% of the poverty level (Curtin et al., 2012).
+To address this, we propose a strategy that appropriately accounts for both confounding and selection bias when estimating the controlled difference in outcomes between groups, where an individual’s probability of selection depends on their group membership and its confounders.
+Rosenbaum and Rubin (1983) proposed using propensity scores to balance the distributions of covariates between comparison groups in the presence of confounding bias.
+In traditional causal inference, these comparison groups are analogous to randomized ‘treatments’ or ‘exposures,’ and using propensity scores to reduce confounding bias relies on several assumptions, namely exchangeability of the treatment/exposure, positivity of the treatment/exposure, and the stable unit treatment value assumption.
+However, certain observational scenarios preclude randomizing individuals due to moral or biological restrictions.
 These include practical examples such as the timing of transplant waitlisting or the inability to ‘assign’ race to an individual (Peng et al., 2017; Li and Li, 2023).
-Our paper addresses this knowledge gap by proposing a strategy for incorporating survey weights from complex survey designs with propensity score methods to achieve covariate balance among heterogeneous groups, such as race.
-Given the complex relationship between race, socioeconomic status (SES), and the porous collection of such data in the United States, the ability to make definitive statements on race differences in health outcomes is immensely challenging (Li and Li, 2023).
-Observational data can offer a wealth of information, however, many observational studies have their limitations in terms of confounding, covariate imbalance, and a lack of representation for generalizing results (Imai and Ratkovic, 2013; Pattanayak et al., 2011).
-Nationally representative surveys provide the opportunity to study individuals from a range of socioeconomic and cultural backgrounds using techniques such as stratified and clustered sampling methods, however researchers must consider the additional challenge of generalizing results outside the context of the sample.
-Such generalizations of group differences falter unless we adjust our comparison groups so that we have have similar levels of other covariates that are correlated with the comparison group and outcome of interest.
-This adjustment, which we term covariate balance can be accomplished a number of ways.
-In this paper, we use propensity scores.
-Rosenbaum and Rubin (1983) originally proposed propensity scores as a method of contrasting the conditional probability of experimental group assignment while adjusting for baseline covariates (Rosenbaum and Rubin, 1983).
-In a typical causal framework, the ability to reduce biases due to confounding by balancing covariate differences is a task that relies on several assumptions, namely (1) strong exchangeability of the treatment/exposure, (2) positivity of the treatment/exposure, and (3) stable unit treatment value.
-When transporting to a different sample or generalizing to a larger target population, we also require (4) strong selection exchangeability and (5) selection positivity.
-Rather than interpreting the difference in the observed outcomes as a causal effect, we seek to estimate a controlled descriptive comparison to adjust for differences in covariates between groups.
-Thus, we leverage propensity scores to compute what Li et al.
+Rather than interpreting differences in outcomes across these groups as causal effects, we seek a controlled descriptive comparison.
+Propensity scores can still be used to compute what Li et al.
 (2013) call an average controlled difference (ACD).
-This quantity accounts for associations between variables of interest under weaker assumptions – positivity (1) and (5) and weak selection exchangeability (4) – but does not render a causal interpretation to the estimated group differences.
-While we later show that our proposed methods also target a causal quantity such as the population average treatment effect (PATE) under stronger assumptions, we do not require this for validity of our proposed estimation procedure, as one or more of these assumptions may be violated in observational data settings – particularly when measuring racial differences in health outcomes.
-We will later present an auxiliary example where we illustrate PATE estimation for an environmental exposure.
-To further complicate things in the study of health disparities, it is often the case that the population characteristics of interest, such as race, are factors that influence the probability of being selected into the sample.
-For instance, the sampling mechanism for the National Health and Nutrition Examination Survey (NHANES) is partially determined by race and indicators of socioeconomic status.
-For example, NHANES oversamples non-Hispanic Black individuals and those individuals at or below 130% of the poverty level (Curtin et al., 2012).
-Previous studies have discussed using propensity score methods for analyzing complex survey data (Austin and Stuart, 2015; Lee et al., 2011; DuGoff et al., 2014; Fuentes et al., 2022), however, not when sample selection depends on race (or other variables commonly used for comparing outcome differences).
+This quantity achieves covariate balance under the weaker assumption of positivity, but it does not render a causal interpretation to the estimated group differences.
+We will refer to race as the ‘group’ (rather than ‘treatment’) of interest hereafter.
+While Li and Li (2023) showed that propensity scores can be used in this way to study racial disparities, a gap remains in determining the ideal practice when the sampling design is complex.
+Many studies have discussed using propensity score methods to generalize inference from randomized trials (Stuart et al., 2011; Lim et al., 2018; Zeng et al., 2021), and complex surveys (Austin and Stuart, 2015; Lee et al., 2011; DuGoff et al., 2014; Fuentes et al., 2022), however, not when sample selection depends on the group of interest, such as race.
 For example, Rossen and Talih (2014) used propensity scores to assess disparities in weight among US children and adolescents from 2001 to 2010 with data from NHANES.
-However, there was not consensus on how to appropriately incorporate survey weights at the propensity score estimation stage.
-There are several challenges in the problem we consider that are distinct from a traditional PATE estimation problem.
-A key assumption and function of the propensity score, traditionally, is conditional independence of the outcome and comparison group of interest conditional on the propensity score or any balancing score.
-Due to the complex nature of these relationships, it is not immediately clear which variables should be conditioned on in the propensity score versus the downstream outcome model.
-Second, it is often useful to express the research question and corresponding dependencies in terms of conceptual diagrams or directed acyclic graphs (such as Figure 1).
-In estimating the ACD, we include diagrams to demonstrate associations between the variables of interest.
-However, given the nature of the variables considered, we caution against interpreting these relationships as causal.
-If considering these diagrams broadly and even causally, in some instances it would be possible to open backdoor pathways when conditioning on selection into the survey, allowing for the possibility of violating assumptions that causal conclusions rely on.
-Addressing this in practice often relies on multi-stage modeling techniques such as inverse probability weighting (IPW) or direct standardization via ‘g-methods’ (see Richardson and Rotnitzky, 2014, for a more detailed history).
-In this paper, we introduce a strategy for incorporating sampling mechanisms, such as survey weights, and propensity scores to make statements on the controlled difference in outcomes when an individual’s probability of selection depends on the group membership variable of interest (herein, race; see Figure 1).
-We explore how the proposed methods compare to current analytic strategies.
-Our motivation comes from a research question addressing the relationship between race, SES, and telomere length, a biomarker for health measured in a particular nationally representative survey (NHANES).
-The remainder of this manuscript is structured as follows.
+However, there was no consensus on whether to incorporate survey weights when estimating the propensity score and/or outcome model.
+Our paper addresses this knowledge gap by proposing a strategy for incorporating survey weights from complex survey designs with propensity score methods to achieve covariate balance and generalizability, specifically when survey selection depends on the group of interest.
+We offer four approaches for estimating the ACD, which rely on multi-stage modeling techniques such as inverse probability weighting (IPW) or direct standardization via ‘g-methods’ (see Richardson and Rotnitzky, 2014) and only require the assumptions of positivity, selection positivity, and weak selection exchangeability.
+While we later show that our proposed methods also target a causal quantity such as the population average treatment effect (PATE) under stronger assumptions, we do not require this for validity of our proposed estimation procedure, as one or more of these assumptions may be violated in observational data settings – particularly when studying racial differences in health outcomes.
+In this work, we introduce a novel framework that enhances existing approaches for estimating racial disparities in health outcomes in complex survey settings.
+Our contributions are threefold.
+First, we derive new identification formulas that provide theoretical insight into the estimation of controlled group differences under conditions of confounding and selection bias, addressing a critical gap in the literature where selection depends on the group of interest, such as race.
+Second, we present a comprehensive set of simulation settings that consider different data-generating mechanisms and allow for a thorough examination of the proposed methodology’s performance under various real-world scenarios.
+Lastly, we implement these methods within a comprehensive R package, enabling researchers to rigorously and efficiently apply these techniques to their own data.
+Together, these contributions offer a robust framework for analyzing health disparities using complex survey data.
+The rest of this manuscript is structured as follows.
 Section 2 reviews related work in this area.
 Section 3 introduces and justifies our proposed methods for estimating the quantity of interest, which we will refer to as the ACD in the spirit of Li and Li (2023), and discusses differences with other common approaches.
-Section 4 provides simulation studies to support our recommendations.
+Section 4 explores how the proposed methods compare to current analytic strategies and provides simulation studies to support our recommendations.
 Section 5 introduces the data motivating this work and provides results for this example.
+We then present an auxiliary example where we estimate the PATE for an environmental exposure.
 We close with a summary discussion of the concepts outlined here and some final recommendations to researchers who may be working under similar considerations.
-Figure 1: Conceptual diagrams.
-(A) Assumed variable relationships: $S$ = sample selection indicator; $A$ = group membership variable of interest; $X$ = covariates; $Y$ = observed outcome.
-(B) Conceptual diagram for our motivating data example: SES = Socioeconomic Status.
-(C) Directed acyclic graph for our auxiliary data example.
 2 Related Work
 To our knowledge, Zanutto (2006) and DuGoff et al.
-(2014) were among the first to discuss propensity score methods as an important tool for addressing confounding in observational studies that may include survey weights.
-Both authors suggest that, intuitively, since the purpose of incorporating propensity scores is to achieve in-sample balance rather than in the population, weighting the propensity score by the survey weight is unnecessary (Zanutto, 2006; DuGoff et al., 2014).
-The authors also recommend including survey weights as covariates in the propensity model, drawing a parallel between this intuition to the recommendation by Lumley (2011) of including these covariates to account for non-response in surveys (Lumley, 2011).
+(2014) were among the first to discuss using propensity score methods to address confounding in observational studies with survey weights.
+Both authors suggest that, intuitively, since the purpose of propensity scores weighting is to achieve in-sample balance, survey weighting the propensity score is unnecessary.
+The authors also recommend including survey weights as covariates in the propensity model, drawing a parallel between this intuition to the recommendation by Lumley (2011) of including these covariates to account for non-response in surveys.
 However, this simple inclusion is often problematic in practice, as it may lead to separation in the data or large biases (Ridgeway et al., 2015).
 DuGoff et al.
-(2014) also addressed limitations regarding how to handle extreme weights, model misspecification, nonlinear propensity score estimation, and additional covariates in the models.
-Unfortunately, these recommendations are largely based on intuition, necessitating more rigorous arguments of consistency and conditional independence.
+(2014) also addressed limitations regarding how to handle extreme weights, model misspecification, nonlinear propensity score estimation, and additional model covariates.
+Such recommendations are largely based on intuition for our particular setup and require more rigorous study.
 Subsequently, Ridgeway et al.
@@
 A natural question is in what stages of the analysis (i.e., the propensity score and/or outcome model stage) should the survey weights be included?
-The authors recommended that researchers use survey weights in both stages of the based on their results.
+The authors recommended that researchers use survey weights in both stages based on their results.
 However, the simulation results pertaining to our scenario are inconclusive as to which approach achieves the lowest bias, and the derivations of consistency are not specific to this scenario.
-Further, they consider situations with only the population average treatment effect on the treated (PATT) as their target estimand (Ridgeway et al., 2015).
+Further, they consider only the population average treatment effect on the treated.
 Similarly, others have addressed the task of combining causal inference and survey methodology.
@@
 (2018) further studied the causal inference framework for estimating population versus sample effects and explored the impact of non-response mechanisms with matching techniques.
-Their simulation setting followed that of Austin et al.
-(2018) and conclude that the degree of misspecification impacted their analysis results and population balance.
-In our setting, the sample selection weights are supplied by NHANES, however the true sample selection mechanism is unknown.
-Further, there are many SES variables collected in the survey, which are potential covariates in selection and propensity score models.
-Here, considerations as to the limited degrees of freedom in the modeling process with the addition of survey weights, strata, and PSU’s may be especially important.
-Lastly, race is part of the selection mechanism in the NHANES design, which complicates the relationships underlying our motivating research question.
+Their simulation setting followed Austin et al.
+(2018) and concluded that the degree of misspecification impacted their analysis results and population balance.
+More recently, Yang et al.
+(2024) conducted simulations, also following the design of Austin et al.
+(2018), to compare various propensity score weighting approaches for estimating population-level treatment effects on a binary outcome with survey weighted data.
+The authors found that two-stage methods, with predicted outcomes weighted by survey weights, outperformed the other methods for estimating the population average treatment effect on the treated, but for population average treatment effects, the best performing method depended on the degree of model misspecification and propensity score overlap.
+McCaffrey et al.
+(2024) further extended the use of survey and attrition weights from binary treatment propensity score analyses to the continuous exposure setting.
+They first derive the conditions under which sampling or nonresponse weights must be incorporated into both the generalized propensity score (GPS) estimation and the subsequent outcome model to obtain unbiased estimates of a continuous dose-response curve.
+Through an extensive simulation study, they show that while including weights in both stages generally ensures against bias, in some settings (e.g., when the same covariates drive both treatment assignment and sample selection), unbiased estimation is possible without weighting the GPS itself.
+Conversely, naively omitting weights can lead to bias or inefficiency when selection and/or attrition depend on treatment or covariates.
+They further demonstrate that, although weighting both stages often reduces bias, it can inflate variance when weights are highly variable.
+In our setting, the sample selection weights are supplied by NHANES, and race is part of the selection mechanism in the NHANES design, which complicates the relationships underlying our motivating research question.
 3 Methods
 3.1 Notation
-Let $Y$ denote our observed outcome, here telomere length.
-Our goal is to estimate the de-confounded, or controlled difference in mean telomere length between non-Hispanic White and non-Hispanic Black-identifying individuals in the United States.
+Let $Y$ denote our outcome, here telomere length.
+Our goal is to estimate the deconfounded, or controlled difference in mean telomere length between non-Hispanic White and non-Hispanic Black-identifying individuals in the US.
 To that end, let $\mathcal{A}$ be the discrete set of potential comparison groups of interest, such that $a\in\mathcal{A}$ denotes a specific group.
 We are interested in comparing the values of $Y$ across levels of $A$.
-Our motivating analysis departs from the traditional causal literature in that $\mathcal{A}$ represents race, with $a=1$ denoting Black race and $a=0$ denoting White race, rather than a potential ‘treatment,’ per se (see Li and Li, 2023, for further discussion on using propensity scores for racial disparities analysis).
-Given the nature of our variables, and race in particular, this contrast of interest will primarily be viewed as a controlled comparison of the observed outcomes after balancing the two groups with respect to potential confounding variables (Li et al., 2013; Li and Li, 2023).
+Our motivating analysis departs from the traditional causal literature in that $\mathcal{A}$ represents race, with $a=1,0$ denoting Black versus White race, respectively, rather than a potential ‘treatment,’ per se.
+Given the nature of these variables, our contrast of interest will be viewed as a controlled comparison of the observed outcomes after balancing the two groups with respect to potential confounding variables (Li et al., 2013; Li and Li, 2023).
 As the lived experience of Black and White individuals differs significantly in terms of social determinants of health, we believe a controlled descriptive comparison can help inform the dynamics of these complex relationships (Li and Li, 2023).
-While race may not be interpreted as a modifiable treatment, we conceptualize race as an important construct worthy of study.
-Moreover, the methods described herein are equally applicable in settings where a participant’s sample selection is dependent on a treatment condition or exposure status in a more traditional causal inference framework (see Section 3.2 for more details).
-Our observed data, $\mathcal{D}=\{(Y_{i},A_{i},X_{i},S_{i}=1);i=1,\ldots,n\}$, consist of $n$ independently and identically distributed copies of $(Y,A,X,S)$ sampled from a population distribution, $\mathcal{P}$, of $N$ individuals where $Y$ is the observed outcome, $A$ is a participant’s group membership, $X$ is a $p$-dimensional vector of covariates, and $S$ is an indicator for being selected into the sample.
-In our motivating example, $\Pr(S_{i}=1)$ is known for each individual $i$ in the super population through the explicit NHANES sampling design.
+While race may not be interpreted as a modifiable treatment, the methods described herein are equally applicable in settings where sample selection depends on a treatment in a more traditional sense (see Section 3.2).
+Our observed data are $\mathcal{O}=\{(Y_{i},A_{i},\boldsymbol{X}_{i});i=1,\ldots,n\}\overset{\text{i.i.d.}}{\sim}\mathcal{Q}:=\Pr(\ \cdot\mid S=1)$, where $\mathcal{Q}$ is the sample law for units selected into the analytic sample ($S_{i}=1$), $S\in\{0,1\}$ indicates selection, and $\boldsymbol{X}\in\mathbb{R}^{p}$ denotes our (potentially imbalanced) covariates.
+The underlying population law is $\mathcal{P}:=\Pr(\cdot)$ on $N$ individuals.
+In our motivating data, survey weights, and hence the design probabilities, $\Pr(S=1\mid\cdot)$, are known from the sampling plan.
 There is also the possibility of having missing outcome values among sampled individuals, but this situation will not be explored here and is instead left to future work.
-Further, our assumed variable relationships are given in Figure 1.
-Here, we can see that our outcome ($Y$), the variable of interest ($A$), and the sample selection indicator ($S$) all depend on a particular set of covariates ($X$), and sample selection also depends on race.
+Our assumed variable relationships are given in Figure 1.
+Here our outcome ($Y$), the variable of interest ($A$), and the sample selection indicator ($S$) all depend on covariates ($\boldsymbol{X}$), and sample selection also depends on $A$.
+Unlike in a causal setting, where $\boldsymbol{X}$ consists of variables which affect both the treatment assignment and the outcome (for example, consider the DAG in the third panel of Figure 1), in the ACD setting, $\boldsymbol{X}$ consists of covariates which are unbalanced across $A$, and will bias the ACD estimate if not properly accounted for.
+Further, $\boldsymbol{X}$ depends on the research question.
+That is, $\boldsymbol{X}$ is the set of covariates for which we wish to balance to estimate the controlled difference between outcomes.
+To some extent, this answers the question of, if we balance the distribution of $\boldsymbol{X}$ between groups, what is the “residual difference” in outcomes?
+Here, if we do not account for the contribution of $\boldsymbol{X}$ to the group differences, then we cannot estimate a balanced, marginal effect of the comparison group membership.
+Figure 1: Conceptual diagrams.
+(A) Assumed variable relationships: $S$ = sample selection indicator; AA = group membership variable of interest; $\boldsymbol{X}$ = covariates; $Y$ = observed outcome.
+(B) Conceptual diagram for our motivating data example: SES = Socioeconomic Status.
+(C) Directed acyclic graph for our auxiliary data example.
 3.2 Proposed Estimators
-We are interested in drawing inference about the super population of individuals from which we have selected our study participants.
-Here, the targets of inference are the conditional outcome means, i.e., $\mathbb{E}\left[Y\mid A=a,X\right]$ for $a\in\mathcal{A}$ and their expected difference over $X$, i.e.,
-${\rm ACD}=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=1,X\right]-\mathbb{E}\left[Y\mid A=0,X\right]\right].$ (1)
-Our goal is to estimate the individual group specific conditional means, of which ACD is the difference. We make the following assumptions:
+We wish to draw inference on the population from which our analytic sample was selected. For $a\in\{0,1\}$, we define
+$\mu(a)=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right],$
+where $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})$ and $w(\boldsymbol{X})$ is a known design selection factor.
+The target of inference, the ACD, is given by ${\rm ACD}=\mu(1)-\mu(0)$.
+We make the following assumptions:
   1. (i)
-    Positivity: $\Pr(A=a\mid X=x)>0\ \forall a\in\mathcal{A}$ and every $x$ such that $f_{X}(x)>0$
+    Positivity: $e_{a}^{\mathcal{Q}}(\boldsymbol{X}):=\mathcal{Q}(A=a\mid\boldsymbol{X})>0$ $\mathcal{Q}$-a.s. $\forall a\in\mathcal{A}$
   2. (ii)
-    Positivity of Sample Selection: $\Pr(S=1\mid A=a,X=x)>0$ for every $a,x$ such that $f_{A,X}(a,x)>0$
+    Selection Positivity: $\pi_{a}(\boldsymbol{X}):=\Pr(S=1\mid A=a,\boldsymbol{X})>0$ on the support of $A,\boldsymbol{X}$
   3. (iii)
-    Weak Selection Exchangeability: $\mathbb{E}[Y\mid A=a,X]=\mathbb{E}[Y\mid A=a,S=1,X]$
-We require (i) and (ii) to ensure that our observed data functional, denoted in the next section by $\mu(a)$, is well defined.
+    Weak Selection Exchangeability: $\mathbb{E}_{\mathcal{P}}(Y\mid A=a,\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})=g_{a}(\boldsymbol{X})$.
+We require (i) and (ii) to ensure that our observed data functional, $\mu(a)$, is well-defined.
 Condition (iii) allows us to generalize the conditional mean of the outcome from the study sample to the target population.
-This assumption is less stringent than exchangeability in distribution (i.e, $Y\perp S\mid A,X$), as it does not require the entire conditional distribution of the outcomes to be independent of selection, only their conditional expectations.
-Depending on the factorization of the joint probability of being sampled $(S=1)$ and belonging to a particular group $(A=a)$, we propose three formulas, which can be estimated using either outcome model–based estimators (i.e., g-computation) or inverse probability weighted estimators:
-$\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right],$ (2)
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right],$ (3)
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right].$ (4)
-We will refer to the methods for estimating Equations (2), (3), and (4) as Outcome Modeling and Direct Standardization (OM), Inverse Probability Weighting 1 (IPW1), and Inverse Probability Weighting 2 (IPW2), respectively.
-For additional details on this derivation, see Appendix A.1.
+Specifically, the weak selection exchangeability assumption posits that the expected outcome, conditional on confounders, $\boldsymbol{X}$, and the exposure/group membership variable of interest, $A$, is the same between the sample and the target population.
+This assumption is less stringent than exchangeability in distribution (i.e, $Y\perp S\mid A,\boldsymbol{X}$), as it does not require the entire conditional distribution of the outcomes to be independent of selection, only their conditional expectations.
+Under survey sampling, this assumption automatically holds when the survey weights fully encode the (known) design and response adjustments.
+Further, when the parameter of interest is a conditional quantity, incorporation of the survey weights will not affect the bias or standard error if the model is correctly specified.
+However, we are interested in ACD, which is a marginal quantity, thus the weighting is needed to standardize to the target population.
+In other words, even though the survey weights adjust for the sampling design, incorporating them into the ACD is necessary to correctly weight the final estimator to generalize and deconfound estimation.
+How the survey weights are incorporated depends on the factorization of the joint probability of exposure/group membership and selection.
+Depending on how we factorize the joint probability of being sampled $(S=1)$ and belonging to a particular group $(A=a)$, $\mu(a)$ admits the following equivalent identification formulas under the sample law, $\mathcal{Q}$, given by
+$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right],\quad\text{or}$ (1)
+$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y\Pr(S=1)}{\Pr(S=1,A=a\mid\boldsymbol{X})}\right],$ (2)
+where $w(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}=\tfrac{\Pr(S=1)}{\pi(\boldsymbol{X})}$, $\mathbb{I}(\cdot)$ is the indicator function, and Equation (2) is equivalent to the following two factored forms:
+$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y}{\mathcal{P}(A=a\mid\boldsymbol{X})}\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid A=a,\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathbb{I}(A=a)Y}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}\right],\quad\text{or}$ (3)
+$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y}{\mathcal{Q}(A=a\mid\boldsymbol{X})}\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)Y}{e_{a}^{\mathcal{Q}}(\boldsymbol{X})}\right],$ (4)
+where $e_{a}^{\mathcal{Q}}(\boldsymbol{X}):=\mathcal{Q}(A=a\mid\boldsymbol{X})$ is the within-sample propensity, $e_{a}^{\mathcal{P}}(\boldsymbol{X}):=\mathcal{P}(A=a\mid\boldsymbol{X})$ is the population propensity, and $w_{a}(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\Pr(S=1\mid A=a,\boldsymbol{X})}$.
+We refer to (1), (3), and (4) as Outcome Modeling (OM), Inverse Probability Weighting 1 (IPW1), and IPW2, respectively.
+For the full derivation of (1) - (4), see Appendix A.1.
 The assumptions required to estimate the ACD are relatively weaker than those required to identify a causal quantity.
-Figure 1, Panel (C) gives an example directed acyclic graph (DAG) where a causal target may be warranted, such as the case where $A$ is an environmental exposure, like lead, which can also alter telomere length.
+Figure 1, Panel (C) gives an example of a directed acyclic graph (DAG) where a causal target may be warranted, such as the case where $A$ is an environmental exposure, like lead, which can also alter telomere length.
 We provide a supplementary analysis in which we estimate the causal effect of lead exposure on telomere length for illustration, as a practitioner may also be interested in such causally interpretable estimated differences.
-If the group variable under study is a treatment or exposure, we show under stronger assumptions that Formulas (2) - (4) can equivalently target the population potential outcome means, denoted by $\mathbb{E}[Y^{a}]\ \forall a\in\mathcal{A}$.
-Along with the previous assumptions, the following conditions,
-  1. (iii*)
-    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}[Y^{a}\mid A=a,X]=\mathbb{E}[Y^{a}\mid A=a,S=1,X]$
+If the group variable under study is a treatment or exposure, we show under stronger assumptions that Formulas (1) - (4) can equivalently target the population potential outcome means, denoted by $\mathbb{E}_{\mathcal{P}}[Y^{a}]\ \forall a\in\mathcal{A}$.
+Along with the previous assumptions (i) and (ii), the following conditions,
+  1. (iii’)
+    Weak Selection Exchangeability: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]=\mathbb{E}_{\mathcal{Q}}[Y^{a}\mid A=a,\boldsymbol{X}]$
   2. (iv)
-    Stable Unit Treatment Value Assumption (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
+    Stable Unit Treatment Value Assumption (SUTVA): $Y_{i}=Y_{i}^{a}\ \forall i,\ A_{i}=a\in\mathcal{A}$
   3. (v)
-    Weak Exchangeability: $\mathbb{E}[Y^{a}\mid X]=\mathbb{E}[Y^{a}\mid A=a,X]$,
-are sufficient to identify $\mathbb{E}[Y^{a}]$.
-The additional SUTVA assumption underpins the existence of the counterfactual outcomes, $Y^{a}$.
+    Weak Exchangeability: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid\boldsymbol{X}]=\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]$,
+are sufficient to identify $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+SUTVA underpins the existence of the counterfactual outcomes, $Y^{a}$.
 The compact representation in (iv) captures two aspects of SUTVA – no interference, which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and consistency, which states that the potential outcome is well-defined for each treatment level, $a$ (Rubin, 1980).
 The weak exchangeability assumption allows us to represent the conditional distribution of the unobserved potential outcome using that of the observed potential outcome.
-As before, we only require a weaker version of this assumption, though strong exchangeability, i.e., $Y^{a}\perp A\mid X$, is often assumed.
+As before, we only require a weaker version of this assumption, though strong exchangeability, i.e., $Y^{a}\perp A\mid\boldsymbol{X}$, is often assumed.
 For additional details and the full identification proof, see Appendix A.2.
@@
 3.3 Comparisons to Existing Approaches
-The three formulas derived above differ from standard, existing approaches that one might take in practice, which warrants some brief discussion here.
+Formulas (1) - (4) differ from standard approaches one might take in practice, which warrants a brief discussion here.
 Table 1 compares our proposed approaches to such existing methods.
-As shown, there are two sources of bias if our assumed conceptual diagram holds – bias from covariate imbalance between comparison groups (referred to herein as confounding bias) and unequal selection induced by the design of the complex survey (referred to herein as selection bias), which impacts the generalizability of our within-sample results to a larger, target population.
-Existing approaches may incorporate either or both balancing and/or generalizability weights to account for these two sources of bias, however if either are foregone, then we expect to obtain biased estimates of our target quantity of interest.
-For instance, the estimated (unadjusted) effect of race can be either exaggerated or attenuated depending on the direction of the true association between the confounders (e.g., SES), race, and the outcome in simple regression (Approach 1 in Table 1; Williams et al., 2019).
-Multiple regression (Approach 2) further adjusts for covariates, but unless the outcome model includes the appropriate interaction terms, this approach will suffer from confounding bias as well (Rao, 1971).
-With inverse probability weighting (Approach 3), the race-covariate relationship is modeled, rather than the covariate–outcome relationship, foregoing the need to specify any race $\times$ covariate interactions (Mansournia and Altman, 2016).
+If our assumed conceptual diagram holds, there are two sources of bias: bias from covariate imbalance between comparison groups (i.e., confounding bias) and unequal selection induced by the design of the complex survey (i.e., selection bias), which impacts the generalizability of our within-sample results to a larger, target population.
+Existing approaches may use balancing and/or generalizability weights to address these biases, but omitting either can result in biased estimates of our target quantity.
+For instance, the effect of race can be either exaggerated or attenuated depending on the direction of the true association between confounders, race, and the outcome in simple regression (Approach 1 in Table 1; Williams et al., 2019).
+Multiple regression (Approach 2) adjusts for covariates but will suffer from bias due to model misspecification unless the outcome model includes, for example, appropriate interaction terms (Rao, 1971).
+With inverse probability weighting (Approach 3), the race-covariate relationship is modeled, rather than the covariate-outcome relationship, foregoing the need to specify any race $\times$ covariate interactions (Mansournia and Altman, 2016).
 However, with all three of these approaches, the target of inference is the within-sample difference in outcomes, which cannot be generalized to a larger population if a probability sample was taken (Shi et al., 2023; Degtiar and Rose, 2023).
-Survey-weighted regression (Approach 4) accounts for selection bias by accounting for how many individuals in the population that one individual in the sample represents (Lumley and Scott, 2017), allowing for generalizability, but these results will be biased if there is covariate imbalance in the sample (DuGoff et al., 2014; Ridgeway et al., 2015).
-Table 1: Comparison of existing approaches for analysis to our proposed approach. Dashes denote that the model/weight is not used.
-IPTW = Inverse Probability of Treatment Weighting
-$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X)$
+Survey-weighted regression (Approach 4) accounts for selection bias by accounting for how many individuals in the population that one individual in the sample represents (Lumley and Scott, 2017), allowing for generalizability, but will be biased if there is covariate imbalance (DuGoff et al., 2014).
+Table 1: Comparison of existing approaches for analysis to our proposed approach. Dashes (–) denote that the model is not used.
+Model Bias Being Addressed
+Approach Outcome Propensity Selection Confounding Selection
+Existing Approaches
+1. Simple Regression $\mathbb{E}[Y\mid A,S=1]$ No No
+2. Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ No No
+3. IPTW Estimatora $\Pr(A=1\mid\boldsymbol{X},S=1)$ Yes No
+4. Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(S=1\mid A,\boldsymbol{X})$ No Yes
+5. IPTW Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ Yes No
+6. IPTW + Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
+7. Weighted IPTW + Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(A=1\mid\boldsymbol{X})$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
+Proposed Approaches
+8. Outcome Modeling and Direct Standardization $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(S=1\mid\boldsymbol{X})^{{\bf b}}$ Yes Yes
+9. Inverse Probability Weighting 1 $\Pr(A=1\mid\boldsymbol{X})$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
+10. Inverse Probability Weighting 2 $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid\boldsymbol{X})$ Yes Yes
+11. Augmented Inverse Probability Weighting $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid\boldsymbol{X})$ Yes Yes
+a IPTW: Inverse Probability of Treatment Weighting.
+b $\Pr(S=1\mid\boldsymbol{X})=\Pr(S=1\mid A=1,\boldsymbol{X})\Pr(A=1\mid\boldsymbol{X})+\Pr(S=1\mid A=0,\boldsymbol{X})\Pr(A=0\mid\boldsymbol{X})$.
+In our research question, we have an intersecting consideration of both selection and confounding bias; thus, neither source of bias should be disregarded.
 A more principled approach would be one that accounts for both confounding and selection bias, as given by Approaches 6 and 7, which utilize a combination of propensity score and survey weighting.
-These approaches are advocated for by Zanutto (2006); DuGoff et al.
-(2014) and Ridgeway et al.
-(2015), respectively, and may be viewed as current analytic standard since they suggest simple survey-weighted generalizations of the usual use of IPTW, either at the first (propensity model) and/or second (outcome model) stage.
-Though it may seem that either simple extension addresses both sources of bias, there has been little theoretical justification of how weighting one or multiple models within IPTW accomplishes both tasks.
-This is also true of the ‘naive g-computation’ approach (Approach 8) we study, which weights the predicted conditional expectation for each individual’s outcome and comparison group/race by their survey contribution.
+These approaches, advocated by Zanutto (2006), DuGoff et al.
+(2014), and Ridgeway et al.
+(2015) are considered the current analytic standard.
+They suggest survey-weighted generalizations of IPTW, either at the propensity model stage and/or the outcome model stage.
+Although it may seem that either extension addresses both sources of bias, there is little theoretical justification for how weighting one or multiple models within IPTW achieves both tasks in our setting.
 Additionally, propensity scores are often seen as in-sample balancing scores (Zanutto, 2006; DuGoff et al., 2014).
 From this perspective, their generalizability (and corresponding selection bias correction based on survey weights) is often ignored.
-Survey-weighting both the propensity score and outcome model is reasonable (Ridgeway et al., 2015), however directly combining survey-weighted IPWs and survey weights in a second stage outcome regression may also lead to biased results if additional correction is not made.
-In our research question, we have an intersecting consideration of both selection and confounding bias; thus, neither source of bias should be disregarded.
-From a practical standpoint, one question at hand is: “should we implement survey-weighted propensity score estimation in IPTW or not”.
-Our analytic results show that for estimation of the ACD, the propensity score does not have to be survey-weighted.
+Survey-weighting both the propensity score and outcome model is reasonable (Ridgeway et al., 2015), however directly combining survey-weighted IPWs and survey weights in a second stage outcome regression may also lead to biased results.
+From a practical standpoint, one question is: ‘should we implement survey-weighted propensity score estimation in IPTW or not?’ Our analytic results show that for estimation of the ACD, the propensity score does not have to be survey-weighted.
 We see this in Equations (3) and (4), which differ with respect to the factorization of the joint distribution into propensity and survey weight models.
-If the propensity score is not survey-weighted, based on Equation (4), we need $\Pr(S=1\mid X)$, i.e., one may need to marginalize $\Pr(S=1\mid A,X)$ over the group membership variable, $A$.
-If survey weights are included in the propensity score estimation, then we must know $\Pr(S=1\mid A,X)$, i.e., the probability of selection conditional on each level of the group membership variable and the covariates.
-Moreover, our proposed approaches carry additional differences amongst themselves, and prove to be useful in different settings.
-Formula (2), outcome modeling and direct standardization (OM), requires individual-level data on $X$, or knowledge of $f_{X}(x)$, as an outcome regression, $g_{a}(X)$, is fit in practice.
-Formulas (3) and (4), IPW1 and IPW2, do not require estimation of an outcome model and may be the best options depending on how much information is available in practice.
+If the propensity score is not survey-weighted, based on Equation (4), we need $\Pr(S=1\mid\boldsymbol{X})$, i.e., one may need to marginalize $\Pr(S=1\mid A,\boldsymbol{X})$ over the group membership variable, $A$.
+If survey weights are included in the propensity score estimation, then we must know $\Pr(S=1\mid A,\boldsymbol{X})$, i.e., the probability of selection conditional on each level of the group membership variable and the covariates.
+Moreover, our proposed approaches differ from each other and prove useful in different settings.
+Formula (1), outcome modeling and direct standardization (OM), requires knowledge of $f_{X}(x)$, as an outcome regression, $g_{a}(\boldsymbol{X})$, is fit in practice.
+Formulas (3) and (4), IPW1 and IPW2, do not require estimation of an outcome model and may be the best options depending on the level of information available in practice.
 We view IPW1 and IPW2 as simple corrections to what one might do in practice, namely the traditional inverse probability of treatment weighted (IPTW) estimator, and thus may be a better starting point for an analysis of this type as they additionally account for selection bias.
-Further, IPW1 and IPW2 do not assume additivity, and thus may be more robust if the outcome model for (2) is misspecified.
-However, while OM, IPW1, and IPW2 are all consistent, OM is most efficient if correctly specified.
+While IPW1 does align better with the NHANES survey weighting structure, IPW1 versus IPW2 show that the incorporation of the survey weights into each stage of modeling depends on two alternate factorizations for the joint distribution of being sampled and belonging to the exposure/group of interest.
+The choice of IPW1 versus IPW2 is also driven by design, as IPW1 is better suited for retrospective designs, where IPW2 is better suited for prospective designs.
+Further, IPW1 and IPW2 may be more robust if the outcome model for (1) is likely misspecified, however, the additional modeling of the two propensity score models for IPW2 may result in efficiency loss.
+While OM, IPW1, and IPW2 are all consistent, OM is most efficient if correctly specified.
 Thus, depending on the availability of the data or the ability to estimate certain key quantities, a researcher may prefer one approach over the others.
-Regarding the incorporation of propensity and sample selection weights into the analysis, our weights are not the same as those in the traditional g-computation or IPTW frameworks.
-We derived a new estimator with weights both for within-sample inference and for generalizability to our target population.
-Specifically, note the term $\Pr(S=1\mid X)$ in the denominators of Formulas (2) and (4).
-If selection does not depend on $A$, then identification reduces to the standard proof for the IPTW identification formula.
-However, in our setting, selection depends on the participant’s self-reported race.
-Thus, we need the following decomposition to calculate $\Pr(S=1\mid X)$:
-$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X).$
-As with OM, IPW2 requires that the probability of selection be marginalized over $A$, i.e., $\Pr(S=1\mid X)$, and that the propensity model be fit without the incorporation of selection weights, i.e., $\Pr(A=a\mid S=1,X)$.
-IPW1 requires the selection mechanism be known (e.g., in the case of survey weights), conditional on $A$, i.e., $\Pr(S=1\mid A=a,X)$, and that a survey-weighted propensity model be fit, i.e., $\Pr(A=a\mid X)$.
+A doubly robust estimator, by combining both an outcome model and an inverse-probability weighting component, requires only one of these two models to be correctly specified in order to yield a consistent ACD estimate (see Section 3.4.1).
+In practice, this flexibility eliminates the need for a choice strictly between IPW1 and IPW2.
+If either the propensity model or the outcome model are well-specified, the doubly robust approach remains consistent and often achieves smaller variance than either IPTW estimator alone.
+For these reasons, we generally recommend the doubly robust estimator as a practical default.
+Regarding the incorporation of propensity and sample selection weights, our weights differ from those in traditional g-computation or IPTW frameworks.
+We derived a new estimator with weights for both within-sample inference and generalizability to our target population.
+Specifically, note the term $\Pr(S=1\mid\boldsymbol{X})$ in the denominators of Formulas (1) and (4).
+If selection does not depend on $A$, identification simplifies to the standard IPTW identification formula.
+However, in our setting, selection depends on the participant’s self-reported race, requiring the probability of selection to be marginalized over $A$, where
+$\Pr(S=1\mid\boldsymbol{X})=\Pr(S=1\mid A=1,\boldsymbol{X})\Pr(A=1\mid\boldsymbol{X})+\Pr(S=1\mid A=0,\boldsymbol{X})\Pr(A=0\mid\boldsymbol{X}).$
+As with OM, IPW2 uses the marginalized $\Pr(S=1\mid\boldsymbol{X})$ above, rather than directly modeling $\Pr(S=1\mid\boldsymbol{X})$ by omitting $A$ (Wilms et al., 2021).
+Thus, IPW2 requires that both the propensity model be fit without incorporating selection weights, i.e., $e_{a}^{\mathcal{Q}}(\boldsymbol{X})$, for inverse probability weighting, and the survey-weighted propensity model be fit to calculate $\Pr(S=1\mid\boldsymbol{X})$.
+In contrast, IPW1 requires the selection mechanism be known (e.g., in the case of survey weights), conditional on $A$, i.e., $\Pr(S=1\mid A=a,\boldsymbol{X})$, and that a survey-weighted propensity model be fit, i.e., $e_{a}^{\mathcal{P}}(\boldsymbol{X})$.
 The key difference here comes with respect to how the joint propensity and selection mechanisms are factorized.
 3.4 Estimation
-To estimate the ACD based on the three formulas presented above (OM, IPW1, and IPW2), we propose the following. For OM, define the functional
-$\mu(a)=\mathbb{E}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right].$
-We denote the estimator for the expectation of the outcome, $\mathbb{E}[Y\mid A=a,S=1,X]$, by $\hat{g}_{a}(X)$.
-In practice, $\hat{g}_{a}(X)$ can be estimated, for example, using a standard parametric outcome model with a finite-dimensional parameter vector.
-Our outcome model-based estimator for $\mu(a)$ is given by
-$\hat{\mu}_{\rm OM}(a)=\frac{1}{n}\sum_{i=1}^{n}\hat{g}_{a}(X_{i})\frac{\Pr(S_{i}=1)}{\Pr(S_{i}=1\mid X_{i})}.$ (5)
-[…]
-$\hat{\mu}_{\rm IPW1}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(A_{i}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}^{w}_{a}(X_{i})Pr(S_{i}=1\mid A_{i}=a,X_{i})}\quad{\rm or}$ (6)
-$\hat{\mu}_{\rm IPW2}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(\mathcal{A}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}_{a}(X_{i})Pr(S_{i}=1\mid X_{i})}.$ (7)
-[…]
+To estimate the ACD based on the three formulas presented above (OM, IPW1, and IPW2), we propose the following.
+Let $\hat{g}_{a}(\boldsymbol{X})$ be an estimator of $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})$.
+In practice, $\hat{g}_{a}(\boldsymbol{X})$ can be estimated, for example, using a parametric outcome model with a finite-dimensional parameter vector.
+Our outcome model-based estimator for $\mu(a)$ represented in Equation (1) is
+$\hat{\mu}_{\rm OM}(a)=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\hat{g}_{a}(\boldsymbol{X}_{i}).$ (5)
+When the outcome model is correctly specified, $\hat{g}_{a}(\boldsymbol{X})\overset{p}{\rightarrow}g_{a}(\boldsymbol{X})$, leading to the result that $\hat{\mu}_{\rm OM}(a)\overset{p}{\rightarrow}\mu(a)$.
+Alternatively, IPW1 and IPW2 rely on inverse probability (IP) weighting to control for confounding (Horvitz and Thompson, 1952).
+Let $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})$ denote the within-sample propensity model, which provides an estimator for $e_{a}^{\mathcal{Q}}(\boldsymbol{X})$.
+Similarly, let $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})$ denote the survey-weighted estimator of $e_{a}^{\mathcal{P}}(\boldsymbol{X})$.
+These are both estimable, as we assume a probability sample with known survey weights.
+A common approach may be to model the propensity via logistic regression, though many non-parametric approaches have been proposed (Lee et al., 2010).
+Depending on how we factorize the joint probability of sample selection and group membership, we can write the IPW estimators as
+$\hat{\mu}_{\rm IPW1}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}w_{a}(\boldsymbol{X}_{i})\frac{\mathbb{I}(A_{i}=a)Y_{i}}{\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X}_{i})},\quad{\rm or}$ (6)
+$\hat{\mu}_{\rm IPW2}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\frac{\mathbb{I}(A_{i}=a)Y_{i}}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}.$ (7)
+If the propensity model is correct, that is $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})\pi_{a}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})\pi(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$, then $\hat{\mu}_{\rm IPW1}(a)$ and $\hat{\mu}_{\rm IPW2}(a)$ are consistent.
+3.4.1 Augmented Inverse Probability Weighted Estimator
+Following Dahabreh et al.
+(2020), we can also derive the efficient influence function (EIF) for our setting.
+Specifically, the EIF is the canonical gradient for the parameter $\mu(a)$.
+Here, the EIF for $\mu(a)$ is given by
+$\phi_{a}(Y,A,\boldsymbol{X})=\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\left\{\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left[Y-g_{a}(\boldsymbol{X})\right]+g_{a}(\boldsymbol{X})\right\}-\mu(a)=w(\boldsymbol{X})\left\{\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left[Y-g_{a}(\boldsymbol{X})\right]+g_{a}(\boldsymbol{X})\right\}-\mu(a).$
+This also provides us the form of the doubly robust estimator, given by
+$\hat{\mu}_{DR}(a)=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\left\{\hat{g}_{a}(\boldsymbol{X}_{i})+\frac{\mathbb{I}(A_{i}=a)}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}\left[Y_{i}-\hat{g}_{a}(\boldsymbol{X}_{i})\right]\right\},$ (8)
+where $\hat{g}_{a}(\boldsymbol{X}_{i})$ is an estimator for $g_{a}(\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})$ is an estimator for the within-sample propensity score.
+For the ACD comparing groups 1 and 0, the final estimator is given by
+$\widehat{ACD}_{DR}=\hat{\mu}_{DR}(1)-\hat{\mu}_{DR}(0).$
+We provide the full derivation and proof in Appendix A.3.
 3.4.2 Estimation of Uncertain Sample Selection or Non-Probability Samples
-[…]
+Throughout this work, we are motivated by settings where the sample arises from a known probability survey design.
+That is, we know $\pi_{a}(\boldsymbol{X})=\Pr(S=1\mid A=a,\boldsymbol{X})$.
+In practice, however, the sample selection mechanism may be unknown.
+For example, in non-probability samples, such as those arising from the study of electronic health records, we must estimate $\pi_{a}(\boldsymbol{X})$ with an assumed model ((e.g., through beta regression, as recommended by Elliot, 2009).
+In these settings, we replace $\pi_{a}(\boldsymbol{X})$ by a consistent estimator, $\hat{\pi}_{a}(\boldsymbol{X})$.
+Further, we can replace $\pi(\boldsymbol{X})=\Pr(S=1\mid\boldsymbol{X})$, the sample selection probability marginalized over $A$, by $\hat{\pi}(\boldsymbol{X})$, and $\Pr(S=1)$ by its estimate, $\hat{\pi}=n/N$.
+Here, our asymptotic results hold if the propensity and selection models are both correctly specified, then we have that $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})\hat{\pi}_{a}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})\hat{\pi}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$, leading to $\hat{\mu}(a)$ being consistent estimators of $\mu(a)$.
+As consistency depends on correctly specifying the selection model, alternative approaches, such as simplex regression or nonparametric methods, may provide more robust estimates (Kundu et al., 2023; Salvatore et al., 2024).
+If selection depends on additional, unmeasured covariates, $U$, that do not affect the group membership or the outcome, then our procedure remains unbiased, but we may underestimate the analytic standard errors.
+For inference, we must account for this additional step to obtain correct standard errors (see below).
 3.5 Inference
 […]
-$\phi(Y,X,\theta,\mathcal{P})=M^{-1}\psi[Y,X,\theta(\mathcal{P})],$
-where
+$\sum_{i=1}^{n}\psi(O_{i};\boldsymbol{\theta})=0,$
+where $\psi(O;\boldsymbol{\theta})$ stacks the estimating functions for the nuisance models (e.g., outcome and/or propensity, and selection where applicable; see Section 3.4.2) and the target components, and $\boldsymbol{\theta}=\boldsymbol{\theta}(\mathcal{Q})$ denotes the probability limit under $\mathcal{Q}$.
+Under standard regularity conditions, $\hat{\boldsymbol{\theta}}$ is asymptotically linear,
+$\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}(\mathcal{Q})=\frac{1}{n}\sum_{i=1}^{n}\phi(O_{i})+o_{p}(n^{-1/2}),$
+with influence function
+$\phi^{\mathcal{Q}}(O)=-M^{-1}\psi[O;\boldsymbol{\theta}(\mathcal{Q})];\quad M=-\mathbb{E}_{\mathcal{Q}}[\nabla_{\theta}\psi(O;\boldsymbol{\theta})]_{\boldsymbol{\theta}=\boldsymbol{\theta}(\mathcal{Q})}.$
+Here, $M$ is the Jacobian matrix of the estimating equations (Huber, 1996). The population asymptotic covariance matrix is then given by
+$V(\theta)=\frac{1}{n}\Sigma(\theta);\quad\Sigma(\theta)=\mathbb{E}_{\mathcal{Q}}\left[\left(\phi^{\mathcal{Q}}\right)\left(\phi^{\mathcal{Q}}\right)^{\top}\right]=\left[M(\theta)^{-1}\right]B(\theta)\left[M(\theta)^{-1}\right]^{\top};\quad B(\theta)=\mathbb{E}_{\mathcal{Q}}\left[\psi\psi^{\top}\right].$
 […]
-is the Jacobian matrix of the estimating equations (Huber, 1996).
-Using numerical optimization, the covariance matrix of the estimated parameters, $\hat{V}(\hat{\theta})$, including the ACD estimator, is given by
-$\hat{V}(\hat{\theta})=\sum_{i=1}^{n}\phi_{i}(Y,X,\hat{\theta},\mathcal{P})\phi_{i}(Y,X,\hat{\theta},\mathcal{P})^{\prime},$ (8)
-where $\phi_{i}(Y,X,\hat{\theta},\mathcal{P})$ is the empirical evaluation of the $m\times 1$ vector of the $i$th observation’s values of the influence function for $\hat{\theta}$ by replacing the unknown parameters with sample estimates.
+$\hat{\Sigma}_{n}=\frac{1}{n}\sum_{i=1}^{n}\hat{\phi}_{i}\hat{\phi}_{i}^{\top}=\left[\hat{M}^{-1}\right]\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{i}\hat{\psi}_{i}^{\top}\right\}\left[\hat{M}^{-1}\right]^{\top}.$
+Consequently, the estimated covariance matrix of $\hat{\theta}$ is
+$\hat{V}(\hat{\theta})=\frac{1}{n}\hat{\Sigma}_{n}=\frac{1}{n^{2}}\sum_{i=1}^{n}\hat{\phi}_{i}\hat{\phi}_{i}^{\top}=\frac{1}{n}\left[\hat{M}^{-1}\right]\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{i}\hat{\psi}_{i}^{\top}\right\}\left[\hat{M}^{-1}\right]^{\top}.$
+[…]
 3.5.1 Inference in Stratified, Clustered Multi-Stage Survey Samples
-NHANES employs a complex stratified, clustered sampling design.
-While bias incurred from the differential sample selection probabilities (encoded in the survey weights) is accounted for in the estimation procedure, inference is affected by the fact that individuals within strata and primary sampling units (PSUs) are correlated.
-To account for this, we modify this linearized variance estimator, known as the robust or sandwich estimator outside of survey-based analysis, by summing the contributions of the $i$th individual $(i=1,\ldots,n_{j})$, nested in the $j$th primary sampling unit (PSU; $j=1,\ldots,J_{k}$), nested in the $k$th sampling stratum $(k=1,\ldots,K)$.
-Thus, the estimated covariance matrix for our parameters is expressed in terms of the variability of the between PSU-level sums of the first order Taylor approximations within the sampling strata.
-This becomes
-$\sum_{k=1}^{K}\frac{J_{k}}{J_{k}-1}\sum_{j=1}^{J_{k}}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}^{\prime}$ (9)
-where $K$ and $J_{k}$ are the number of strata and number of PSUs within the $k$th stratum, respectively, $\phi_{\cdot jk}(y,\hat{\theta},\mathcal{P})=\sum_{i=1}^{n_{j}}\phi_{ijk}(y,\hat{\theta},\mathcal{P})$ is the sum of the influence function contributions among individuals in the $j$th cluster, and $\bar{\phi}_{\cdot\cdot k}(y,\hat{\theta},\mathcal{P})=\frac{1}{J_{k}}\sum_{j=1}^{J_{k}}\phi_{\cdot jk}$ is the “average” influence function value within the $k$th stratum.
-Software for these methods can be found in the R package svycdiff at https://github.com/salernos/svycdiff, in addition to the code and data necessary to reproduce the simulation studies and data analysis described below.
+[…]
+$\hat{V}(\hat{\theta})=\frac{1}{n^{2}}\sum_{k=1}^{K}\frac{J_{k}}{J_{k}-1}\sum_{j=1}^{J_{k}}\left(\hat{U}_{jk}-\bar{U}_{k}\right)\left(\hat{U}_{jk}-\bar{U}_{k}\right)^{\top}.$ (9)
+3.6 Software
+Software for these methods can be found in the R package svycdiff, available through the Comprehensive R Archive Network (CRAN) at cran.r-project.org/web/packages/svycdiff, or in a development version on GitHub at github.com/salernos/svycdiff.
+As our proposed approach uses an estimating equations framework, the svycdiff accommodates various outcome types, including continuous, binary, and count outcomes.
+This includes the main functions for estimation and inference, as well as a function for simulating population-level data under a wide range of scenarios.
+Moreover, these accommodate the setting described in Section 3.4.2.
+Lastly, the codes necessary to reproduce our later simulations and data analysis are also available at the GitHub link.
 4 Simulation Studies
-We conduct a series of simulation studies to better understand the finite sample performance of our proposed estimators, as compared to approaches typically used in practice.
-Based on the simulation studies conducted in DuGoff et al.
-(2014) and Ridgeway et al.
-(2015), we extend this work in several important ways.
-First, we have our proposed estimators, which represent novel alternatives to various combinations of weighting schemes outlined in Table 1.
-Second, we have varied the set of associations in our data to systematically include or exclude associations between $A$, $X$, and $S$ in our data generating mechanism to understand how certain relationships impact our estimation and inference.
-The purpose of the overall simulation study is to quantify the bias and variance associated with the use of our proposed approach, as compared to the various weighting schemes which may be currently employed.
+We conduct a series of simulations based on the setup of Austin et al.
+(2018), and subsequently Zeng et al.
+(2025), to understand the finite sample performance of our proposed estimators compared to those typically used in practice.
+We systematically vary the associations between $A$, $\boldsymbol{X}$, and $S$ in our data-generating mechanism to understand how these relationships impact estimation and inference.
+We quantify the bias and variance of our proposed approach compared to existing approaches.
 4.1 Data Generation
-As in DuGoff et al.
-(2014) and Ridgeway et al.
-(2015), let $X$ denote a normally-distributed covariate which relates to our outcome, $Y$, race, $A$, and sample selection indicator, $S$.
-We generate population-level data with $X\sim\mathcal{N}(1,1)$ for a super population of $N=100,000$ individuals from which we repeatedly take samples.
-We then generate the remaining data from three models:
-  1. 1.
-    Propensity Model: $\Pr(A\mid X)={\rm logit}^{-1}(\tau_{0}+\tau_{X}X)$,
-  2. 2.
-    Sample Selection Model: $\Pr(S\mid A,X)={\rm logit}^{-1}(\beta_{0}+\beta_{A}A+\beta_{X}X+\varepsilon_{S});\ \varepsilon_{S}\sim\mathcal{N}(0,0.1)$,
-  3. 3.
-    Outcome Model: $f(Y\mid A,X)$,
-where ${\rm logit}^{-1}(u)=[1+\exp\{-u\}]^{-1}$ is the inverse logit function and we add a small amount of noise to the generation of the sample selection indicator to model this variable in our later procedure.
-From these models, we simulate the binary variable of interest, $A\mid X\sim\text{Bin}[N,\Pr(A\mid X)]$, the sample selection indicator, $S\mid A,X\sim\text{Bin}[N,\Pr(S\mid A,X)]$, and sample selection weights, $\omega_{S}=\Pr^{-1}(S\mid A,X)$.
-In simulating the outcome $Y$, we consider a heterogeneous treatment effect which varies with $X$, such that
-$Y=\gamma_{0}+\gamma_{X}X+\gamma_{A}A+\gamma_{AX}AX+\varepsilon_{O}\$
-where $\varepsilon_{O}\sim\mathcal{N}(0,1)$.
-We consider eight simulation settings, over which we vary the coefficients $\tau_{X},\beta_{A}$, and $\beta_{X}$, corresponding to the parameters in the true propensity and selection models (see Figure 2a-b), where each setting corresponds to a progressively more complex relationship among the variables.
-Across all simulation settings, we fixed $\tau_{0}=-1$, $\beta_{0}=-4.5$, $\gamma_{0}=1$, $\gamma_{A}=1$, $\gamma_{X}=1$, and $\gamma_{AX}=0.1$.
-Within each setting, we drew $200$ sub-samples based on our simulated sample selection indicator, $S$.
-We fit the various combinations of approaches listed in Figure 2.
-Our target quantity of interest was the ACD.
-We compared the percent bias, mean squared error (MSE), and coverage probability at a significance level of 0.05 for each method.
+We first generate population-level data for $N=1,000,000$ individuals.
+We organize the individuals into 10 strata $(j=1,\ldots,10)$ that are further subdivided into 20 clusters, so that there are 200 unique clusters $(k=1,\ldots,200)$ and 5,000 individuals per cluster $(i=1,\ldots,5,000)$.
+For each individual, we generate two baseline covariates $(l=1,2)$ from normal distributions, allowing the means of these covariates to vary at both the stratum and cluster levels.
+That is, we have
+$X_{l,ijk}\sim\mathcal{N}\left(\nu_{l,j}^{s}+\nu_{l,j}^{c},1\right);\quad\nu_{l,j}^{s}\sim\mathcal{N}\left(0,\sigma_{l}^{s}\right),\quad\nu_{l,j}^{c}\sim\mathcal{N}\left(0,\sigma_{l}^{c}\right),$
+where $\nu_{l,j}^{s}$ and $\nu_{l,j}^{c}$ are stratum- and cluster-specific random effects for the $l$th covariate.
+We set $\sigma_{l}^{s}=0.35$ and $\sigma_{l}^{c}=0.15$ across all simulation settings.
+Note that $\boldsymbol{X}_{i}=\{X_{1,i},X_{2,i}\}$ relate to our outcome, $Y_{i}$, group variable of interest, $A_{i}$, and sample selection indicator, $S_{i}$.
+We then generate $A_{i}\in\{0,1\}$ from a logistic model, so the propensity model for $A_{i}=1$ is given by
+$\Pr(A_{i}=1\mid\boldsymbol{X}_{i})={\rm logit}^{-1}\left[\tau_{0}+\tau_{A}\left(\tau_{X_{1}}X_{1,i}+\tau_{X_{2}}X_{2,i}+\tau_{X_{1}X_{2}}X_{1,i}X_{2,i}\right)\right],$
+where ${\rm logit}^{-1}(u)=[1+\exp\{-u\}]^{-1}$ is the inverse logit function.
+Here, $\tau_{0}$ controls the overall group prevalence, $\tau_{A}$ reflects the strength of confounding and, correspondingly, the degree of overlap in the propensity score, $\tau_{X_{1}}$ and $\tau_{X_{2}}$ are coefficients for $X_{1}$ and $X_{2}$, respectively, and $\tau_{X_{1}X_{2}}$ an interaction effect between $X_{1}$ and $X_{2}$.
+For this main simulation, we fix $\tau_{0}=0$, $\tau_{A}=0.5$, and $\tau_{X_{1}X_{2}}=0$, and we vary $\tau_{X_{1}}$ and $\tau_{X_{2}}\in\{0,1\}$.
+We then simulate the $A_{i}$ from a Bernoulli distribution with $A_{i}\sim{\rm Bernoulli}[\Pr(A_{i}=1\mid\boldsymbol{X}_{i})]$.
+We next generate a survey weight for each individual, $\omega_{i}$, by specifying a logistic selection model, with
+$\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})={\rm logit}^{-1}(\beta_{0}+\beta_{A}A_{i}+\beta_{X_{1}}X_{1,i}+\beta_{X_{2}}X_{2,i}+\varepsilon_{S});\quad\varepsilon_{S}\sim\mathcal{N}(0,0.1),$
+where $\beta_{0}$ controls the marginal sampling fraction, $\beta_{A}$, $\beta_{X_{1}}$, and $\beta_{X_{2}}$ are coefficients for $A$, $X_{1}$ and $X_{2}$, respectively, and $\varepsilon_{S}$ is small amount of random noise added to the generation of the survey weights, to support later sensitivity analyses.
+We fix $\beta_{0}=-8$ and vary $\beta_{A}$, $\beta_{X_{1}}$, and $\beta_{X_{2}}\in\{0,1\}$, and we set $\omega_{i}=\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})^{-1}$.
+Lastly, we simulate continuous outcomes, $Y_{i}$, for each individual in the population as a linear combination of $A_{i}$, $\boldsymbol{X}_{i}$, and interaction terms.
+Specifically, we let
+$Y_{i}=\alpha_{0}+\alpha_{X_{1}}X_{1,i}+\alpha_{X_{2}}X_{2,i}+A_{i}\left[\alpha_{A}+\alpha_{AX}\left(\alpha_{X_{1}}X_{1,i}+\alpha_{X_{2}}X_{2,i}\right)\right]+\varepsilon_{O};\quad\varepsilon_{O}\sim\mathcal{N}(0,1),$
+where we set $\alpha_{0}=1$ and $\alpha_{X_{1}}=\alpha_{X_{2}}=1$ for the covariate coefficients.
+We further set $\alpha_{A}=1$ to reflect the true effect of $A$ on $Y$ and $\alpha_{AX}=0.1$ to induce a heterogeneous group/exposure effect which varies with $\boldsymbol{X}$.
+For this main simulation study, we consider a fully factorial design where we vary $\tau_{X_{1}}=\tau_{X_{2}}\in\{0,1\}$, $\beta_{A}\in\{0,1\}$, and $\beta_{X_{1}}=\beta_{X_{2}}\in\{0,1\}$, yielding eight parameter settings, where each setting corresponds to a progressively more complex relationship among the variables: two with no confounding or selection bias, two with confounding bias but no selection bias, two with selection bias but no confounding bias, and two with both selection and confounding bias (see Table LABEL:tab:simulation_results).
+Within each simulation setting, we draw 500 random samples from our generated population by repeatedly simulating sample selection indicators, $S_{i}$.
+As the focus of the current work is in setting of known survey weights, which can be stratified or multistage, we extract the probability of selection for the $i$th individual and generate $S_{i}$ from a Bernoulli distribution with $S_{i}\sim{\rm Bernoulli}[\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})]$, retaining those observations with $S_{i}=1$ in each replicate.
+Note that this deviates slightly from a multistage stratified, clustered design.
+However, we opted for this simplification to focus on the more important aspect of the problem, namely that the final weights need to appropriately factorize the joint distribution of being sampled and belonging to a particular group.
+This holds over a range of sampling designs (see our sensitivity analyses below).
+We compare our proposed (5) – (8) to the standard analytic approaches in Table 1.
+Across all methods where survey weights are considered, we apply the corresponding appropriate standard error calculations.
+For example, for our proposed approaches, we use the variance formula we have derived for stratified, clustered designs (9).
+For comparison methods such as survey-weighted regression, we use design-based standard errors (Lumley, 2011).
+The oracle estimator, which knows the full data generating process, serves as our benchmark throughout.
+Our target quantity of interest is the ACD.
+We compare the relative bias, average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE), and coverage probability at a significance level of 0.05 for each method.
 4.2 Results
-Results from our simulation studies are given in Figure 2c-d and Supplemental Table LABEL:tab:simulation_results.
+Results are given in Table LABEL:tab:simulation_results.
 Here, we discuss these results in the context of each simulation setting, where the data are generated to vary what sources of bias we expect to encounter when estimating the ACD.
-4.2.1 Settings 1 and 3: No Bias Expected
-In Settings 1 and 3, there is neither confounding nor selection bias in the data generating mechanism.
-Setting 1 provides a fair comparison across all the methods considered, as both the group propensity, $\Pr(A\mid X)$, and sample selection probabilities, $\Pr(S\mid A,X)$, were generated independent of any covariates.
-As shown, all of the methods except for the naive g-computation approach perform comparably in this setting with respect to the percent bias of the ACD estimator.
-Specifically, the percent bias among the three proposed estimators ranged from 0.499-0.647%, as compared to 0.166-0.538% among the existing methods (excluding the naive g-computation), suggesting that all of the methods considered accurately estimate the average controlled difference.
-This is expected, as the methods which do not account for confounding or selection bias are correctly specified, and the methods that do are unencumbered by the additional weights.
-The proposed approaches also had among the lowest MSEs, ranging from 0.475-0.480, as compared to 0.482-1.103 among the existing approaches.
-Lastly, across 200 independent replications, the proposed methods had coverage rates of 92.0-92.5%, as compared to 87.5-93.0% for the existing methods.
-Similarly, in Setting 3, where selection depends on $A$ but neither selection nor $A$ depend on $X$, all methods tend to perform well.
-Here, we see that the percent bias for the proposed methods ranged from 0.217-1.497%, on par with the other methods (0.138-0.417%).
-Mean squared errors ranged from 0.219-0.243 for the proposed methods, similar to MSEs of 0.222-0.547 among the existing methods.
-Further, all of the methods had nominal coverage rates of 94.5-99.0%.
-4.2.2 Setting 2 and 4: Confounding Bias Expected
-In Settings 2 and 4, we generate $A$ conditional on $X$, with selection further being dependent on $A$ in Setting 4.
-In these settings, we now expect there to be confounding bias, as $\Pr(A\mid X)$ is no longer uniform over the range of $X$ values.
-As such, we see that all methods which account for both $A$ and $X$ tend to perform comparatively well, whereas the simple linear regression (SLR) of $Y$ on $A$ performs poorly due to confounding bias in $X$.
-Specifically, we see that the percent bias in the SLR estimator has increased from 0.138-0.166% in Settings 1 and 3 to 78.249-79.480% in Settings 2 and 4, whereas the other estimators (including the proposed methods) still see biases ranging in magnitude from 0.977-2.778% in Setting 2 and 0.210-1.526% in Setting 4.
-We further see an increase in MSE for the SLR estimator, with values ranging from 75.408-75.638, while the MSE for the other existing methods (0.404-1.176) and the proposed estimators (0.362-1.220) remained low.
-Once again, the coverage probabilities for the proposed (92.5-94.5%) and other existing (88.5-94.0%) remained at the nominal level, whereas the SLR did not cover the true ACD (0%).
-4.2.3 Settings 5 and 7: Selection Bias Expected
-In Settings 5 and 7, we now generate $S$ conditional on $X$, with selection further being dependent on $A$ in Setting 7.
-Here, we now expect there to be selection bias induced by the $S\mid X$ relationship.
-In Setting 5, the percent bias among the existing approaches ranges from 9.182-9.544%, while the bias among the proposed methods ranges from 0.302-1.252%, demonstrating empirically that the proposed estimators are unbiased for the target ACD.
-Moreover, these results persist in Setting 7, with the percent bias among the existing approaches (3.992-6.809%) being higher than the proposed approaches (0.172-2.322%).
-Similar patterns can be seen when comparing the MSEs of the existing approaches (1.135-1.311 in Setting 5 and 0.587-0.663 in Setting 7) to those proposed in this work (0.236-0.325 in Setting 5 and 0.135-0.265 in Setting 7).
-Further, only the proposed methods achieved the nominal coverage rate (93.5-95.5% in Setting 5 and 91.0-97.0% in Setting 7), while the other methods show coverage ranging from 9.0-36.5% and 19.0-80.5% in Settings 5 and 7, respectively.
-4.2.4 Settings 6 and 8: Both Confounding and Selection Bias Expected
-Lastly, Settings 6 and 8 represent the most complex scenarios of confounding and selection bias, with Setting 8 mirroring the conceptual diagram for our motivating data.
-In this setting, our proposed estimators outperform all current approaches with respect to the percent bias, MSE, and coverage for estimating the ACD.
-In particular, percent bias among existing approaches ranges from 7.847-82.816% in Setting 6 and 6.226-73.914% in Setting 8, whereas the bias among the proposed estimators ranges from 0.633-1.514% and 1.083-1.115%, respectively.
-This is expected as, the proposed methods have the correct balancing and generalizability weights under these settings, whereas the existing approaches fail to incorporate one or both weights, or fail to incorporate them correctly into the form of the estimator.
-The proposed approaches again had the lowest MSEs, ranging from 0.131-0.540 across both settings, as compared to 0.927-82.577 for the existing approaches.
-Lastly, the proposed approaches achieve nominal coverage (93.5-97.0% across both settings), whereas the existing methods do not (0-83.5%).
-For full results, see Supplemental Table LABEL:tab:simulation_results.
-4.3 Sensitivity Analysis
-In a sensitivity analysis, we study our proposed approaches under increasing model misspecification by introducing effect heterogeneity.
-Specifically, we systematically included larger interactions of race and covariates, where we varied $\gamma_{AX}\in\{0,0.01,0.05,0.1,0.5\}$.
-We estimate IPW1 and IPW2 as is, and we fit two versions of OM where we (1) correctly specify versus (2) misspecify the outcome model by not including the interaction term present in the data generating mechanism.
-Here we seek to assess the robustness of these approaches against the oracle estimator.
-The results from this analysis show that overall, all methods tend to perform well in terms of the bias, MSE, and coverage of the estimator, with OM performing acceptably under moderate misspecification.
-In particular, for $\gamma_{AX}\in\{0,0.01,0.05,0.1\}$, the method with the misspecified outcome model attained nominal coverage (ranging from 93.5-96.5%), with biases below 0.05% (0.008-0.028) and comparable MSEs to the other methods (0.001-0.003).
-However this approach incurs bias as treatment heterogeneity becomes more evident.
-Specifically, for the setting in which $\gamma_{AX}=0.5$, we see that the bias increases to 0.14% and the method does not cover the true ACD.
-These results suggest that the IPTW-type estimators (IPW1 and IPW2), which do not assume additivity, may be more robust.
-Full simulation results are presented in Supplemental Table LABEL:tab:sim_sensitivity.
+Settings 1 and 3: No Bias Expected.
+When there is neither confounding nor selection bias in the data generating mechanism, most estimators recover the true ACD with negligible bias and nominal 95% coverage.
+Specifically, the relative bias among our four proposed estimators ranges from -0.002 to -0.001, as compared to -0.003 to 0.005% among the existing methods, as expected.
+The proposed approaches also have the lowest MSEs (0.012), as compared to MSEs ranging from 0.012-0.069 for the existing approaches, as well as standard errors that are comparable to the more efficient approaches (ASE: 0.109-0.113).
+Simple linear regression (SLR) has larger standard errors, though the ASE is within 5% of the MCSE, whereas the ASE of the traditional IPTW estimator is less than 50% that of its MCSE, suggesting the variance is underestimated.
+Lastly, the proposed methods have coverage rates of 95.6-96.2%, as compared to 95.6-96.4% for the existing methods except the IPTW estimator (58.6%).
+In contrast, the traditional IPTW estimator, though, unbiased on average, is particularly susceptible to underestimating the true variance when ignoring the stratified, clustered sampling design.
+in Setting 3, where selection depends on $A$, all methods except IPTW perform well.
+The relative biases for the proposed methods are all -0.005%, equal to the other methods, except IPTW (1.228).
+Mean squared errors are low (0.008) for the proposed methods, similar to MSEs of 0.008 to 0.024 among the existing methods, except IPTW (1.659).
+Further, all the methods had nominal coverage (95.2-96.4%), except IPTW (0%).
+Settings 2 and 4: Confounding Bias Expected.
+When we generate $A$ conditional on $X_{1}$ and $X_{2}$, with selection also depending on $A$ in Setting 4, we now expect there to be confounding bias.
+We see that methods which account for $A$ and $\boldsymbol{X}$ perform comparatively well, whereas SLR performs poorly, as $\Pr(A\mid X_{1},X_{2})$ is no longer uniform over the range of $\boldsymbol{X}$ values.
+Specifically, the relative bias in the SLR estimator increases from 0.002-0.005 in Settings 1 and 3 to 2.021-2.123 in Settings 2 and 4, whereas the other estimators (including the proposed methods) have biases ranging in magnitude from -0.004 to 0 in Setting 2 and 0.001 to 0.027 in Setting 4.
+We further see a higher in MSE for the SLR (1.089-1.357) and IPTW (0.078-1.421) estimators, while the MSE for the other existing methods (0.010-0.015) and the proposed estimators (0.010-0.022) remain low.
+Coverage for the proposed (93.4-94.4%) and other existing (91.4-93.6%) are also at the nominal level, whereas SLR (0%) and IPTW (0-63.4%) do not cover the true ACD.
+Among our proposed estimators, outcome modeling/direct standardization and AIPW deliver unbiased estimates with slightly lower ASE/MCSE than the purely IPW-based estimators, indicating better efficiency in this confounding scenario.
+Settings 5 and 7: Selection Bias Expected.
+When the $\boldsymbol{X}$ variables influence the probability of inclusion but do not confound the relationship between $Y$ and $A$, we expect there to be selection bias.
+In Setting 5, the relative bias among the existing approaches ranges from 0.001 to 0.231, with unweighted methods having higher bias (0.217-0.231) than the survey-weighted approaches (0.001-0.008).
+The bias among the proposed methods ranges from 0.000 to 0.027, demonstrating empirically that accounting for the survey weights corrects this selection bias when estimating the target ACD.
+These results persist in Setting 7, where selection also depends on $A$, with the relative bias among the existing unweighted approaches (0.173-3.560) being higher than existing survey-weighted approaches (0.001-0.005) and the proposed approaches (0.003-0.029).
+Similar patterns are seen in the MSEs of the existing approaches (0.019-0.162 in Setting 5 and 0.008-0.111 in Setting 7) versus the proposed methods (0.043-0.128 in Setting 5 and 0.018-12.637 in Setting 7).
+However, only the proposed methods have nominal coverage (94.2-97.4% in Setting 5 and 93.4-97.8% in Setting 7), while all other methods have coverage ranging from 19.4-85.8% and 0.0-90.2% in Settings 5 and 7, respectively.
+Among our proposed estimators, Outcome modeling again shows highest efficiency (ASE = 0.154 vs.
+MCSE = 0.137), as it is correctly specified, but AIPW approach strikes a balance between bias control and efficiency.
+Settings 6 and 8: Both Confounding and Selection Bias Expected.
+Lastly, Settings 6 and 8 are the most complex, with Setting 8 mirroring the conceptual diagram for our motivating data.
+In these settings, our proposed estimators outperform all current approaches with respect to the relative bias, MSE, and coverage for estimating the ACD.
+In particular, the relative biases for existing approaches that not use survey weights range from 0.203-1.198 in Setting 6 and 0.230-2.872 in Setting 8, and their coverages range from 0-44.6% and 0-14.4%, respectively.
+Further, while survey-weighting reduces bias (-0.004 to 0.021), the existing methods still under-cover (87.4-92.0%), indicating that design adjustments alone are insufficient when confounding and selection bias co-occur.
+In contrast, the bias among the proposed estimators ranges from -0.005 to 0.025, with coverages ranging from 94.6-98.2%.
+This is expected, as the proposed methods have the correct balancing and generalizability weights under these settings, whereas the existing approaches fail to incorporate one or both weights, or fail to incorporate them correctly into the form of the estimator.
+Among our proposed approaches outcome modeling/direct standardization is most efficient, while AIPW again achieves low bias, correct SEs, and nominal coverage, while offering better efficiency than the IPW-based approaches.
+Table 2: Simulation results comparing different analytic methods for calculating the population average controlled difference (ACD) in $Y$ across AA for different data generating mechanism settings.
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
+Setting 1: No Bias Expected ($\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=0$)
+Oracle Estimator 1.003 1.000 -0.003 0.109 0.107 0.012 0.964
+Simple Reg. 1.003 1.004 0.002 0.204 0.206 0.042 0.948
+Multiple Reg. 1.003 1.001 -0.002 0.109 0.107 0.012 0.958
+IPTW Estimator 1.003 1.008 0.005 0.110 0.263 0.069 0.586
+Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
+IPTW Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.958
+IPTW + Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
+Outcome Modeling and Direct Standardization 1.003 1.001 -0.002 0.109 0.107 0.012 0.962
+Inverse Probability Weighting 1 1.003 1.001 -0.002 0.110 0.108 0.012 0.956
+Inverse Probability Weighting 2 1.003 1.002 -0.001 0.113 0.110 0.012 0.960
+Augmented Inverse Probability Weighting 1.003 1.001 -0.002 0.110 0.108 0.012 0.958
+Setting 2: Confounding Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=0$)
+Oracle Estimator 0.996 0.995 0.000 0.115 0.120 0.014 0.930
+Simple Reg. 0.996 2.021 1.030 0.195 0.194 1.089 0.000
+Multiple Reg. 0.996 0.995 -0.001 0.115 0.121 0.015 0.932
+IPTW Estimator 0.996 0.982 -0.014 0.135 0.280 0.078 0.634
+Survey-Weighted Multiple Reg. 0.996 0.995 -0.001 0.116 0.122 0.015 0.932
+IPTW Multiple Reg. 0.996 0.996 0.000 0.116 0.123 0.015 0.936
+IPTW + Survey-Weighted Multiple Reg. 0.996 0.995 -0.001 0.117 0.124 0.015 0.928
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.996 0.995 0.000 0.117 0.124 0.015 0.928
+Outcome Modeling and Direct Standardization 0.996 0.995 -0.001 0.116 0.121 0.015 0.934
+Inverse Probability Weighting 1 0.996 0.993 -0.003 0.136 0.146 0.021 0.942
+Inverse Probability Weighting 2 0.996 0.992 -0.004 0.138 0.147 0.022 0.940
+Augmented Inverse Probability Weighting 0.996 0.996 0.000 0.118 0.124 0.015 0.934
+Setting 3: No Bias Expected ($\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=0$)
+Oracle Estimator 0.991 0.986 -0.005 0.090 0.090 0.008 0.962
+Simple Reg. 0.991 0.996 0.005 0.169 0.167 0.028 0.962
+Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.964
+IPTW Estimator 0.991 2.267 1.288 0.105 0.177 1.659 0.000
+Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.954
+IPTW Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.966
+IPTW + Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.952
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.952
+Outcome Modeling and Direct Standardization 0.991 0.986 -0.005 0.093 0.089 0.008 0.962
+Inverse Probability Weighting 1 0.991 0.986 -0.005 0.093 0.089 0.008 0.962
+Inverse Probability Weighting 2 0.991 0.986 -0.005 0.095 0.090 0.008 0.954
+Augmented Inverse Probability Weighting 0.991 0.986 -0.005 0.093 0.089 0.008 0.960
+Setting 4: Confounding Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$)
+Oracle Estimator 0.969 0.969 0.001 0.094 0.100 0.010 0.938
+Simple Reg. 0.969 2.123 1.191 0.167 0.161 1.357 0.000
+Multiple Reg. 0.969 0.995 0.027 0.097 0.103 0.011 0.930
+IPTW Estimator 0.969 2.146 1.215 0.148 0.188 1.421 0.000
+Survey-Weighted Multiple Reg. 0.969 0.970 0.001 0.094 0.100 0.010 0.940
+IPTW Multiple Reg. 0.969 0.995 0.027 0.098 0.106 0.012 0.914
+IPTW + Survey-Weighted Multiple Reg. 0.969 0.995 0.027 0.098 0.107 0.012 0.920
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.969 0.969 0.001 0.095 0.102 0.010 0.934
+Outcome Modeling and Direct Standardization 0.969 0.970 0.001 0.096 0.100 0.010 0.942
+Inverse Probability Weighting 1 0.969 0.980 0.011 0.111 0.112 0.013 0.944
+Inverse Probability Weighting 2 0.969 0.976 0.008 0.118 0.123 0.015 0.934
+Augmented Inverse Probability Weighting 0.969 0.970 0.001 0.098 0.102 0.010 0.942
+Setting 5: Selection Bias Expected ($\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=1$)
+Oracle Estimator 0.971 0.976 0.006 0.133 0.137 0.019 0.946
+Simple Reg. 0.971 1.187 0.222 0.136 0.127 0.063 0.650
+Multiple Reg. 0.971 1.182 0.217 0.074 0.074 0.050 0.194
+IPTW Estimator 0.971 1.195 0.231 0.075 0.279 0.128 0.280
+Survey-Weighted Multiple Reg. 0.971 0.973 0.002 0.163 0.214 0.046 0.858
+IPTW Multiple Reg. 0.971 1.182 0.218 0.074 0.074 0.050 0.194
+IPTW + Survey-Weighted Multiple Reg. 0.971 0.972 0.001 0.163 0.216 0.046 0.866
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.971 0.978 0.008 0.159 0.207 0.043 0.858
+Outcome Modeling and Direct Standardization 0.971 0.979 0.008 0.154 0.137 0.019 0.974
+Inverse Probability Weighting 1 0.971 0.971 0.000 0.197 0.211 0.044 0.942
+Inverse Probability Weighting 2 0.971 0.944 -0.027 0.327 0.402 0.162 0.954
+Augmented Inverse Probability Weighting 0.971 0.971 0.000 0.209 0.227 0.051 0.946
+Setting 6: Both Confounding and Selection Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$)
+Oracle Estimator 1.002 1.020 0.018 0.116 0.118 0.014 0.946
+Simple Reg. 1.002 2.201 1.198 0.128 0.121 1.454 0.000
+Multiple Reg. 1.002 1.225 0.223 0.079 0.078 0.056 0.192
+IPTW Estimator 1.002 1.205 0.203 0.133 0.336 0.154 0.446
+Survey-Weighted Multiple Reg. 1.002 1.015 0.013 0.144 0.168 0.028 0.898
+IPTW Multiple Reg. 1.002 1.230 0.228 0.080 0.081 0.059 0.202
+IPTW + Survey-Weighted Multiple Reg. 1.002 1.018 0.016 0.147 0.187 0.035 0.874
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.002 1.023 0.021 0.143 0.176 0.031 0.868
+Outcome Modeling and Direct Standardization 1.002 1.021 0.019 0.136 0.119 0.015 0.978
+Inverse Probability Weighting 1 1.002 1.003 0.002 0.187 0.183 0.033 0.956
+Inverse Probability Weighting 2 1.002 1.026 0.025 0.261 0.307 0.095 0.982
+Augmented Inverse Probability Weighting 1.002 1.022 0.020 0.199 0.244 0.060 0.960
+Setting 7: Selection Bias Expected ($\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=1$)
+Oracle Estimator 0.997 0.991 -0.007 0.094 0.091 0.008 0.956
+Simple Reg. 0.997 1.170 0.173 0.097 0.093 0.038 0.562
+Multiple Reg. 0.997 1.205 0.208 0.053 0.051 0.046 0.030
+IPTW Estimator 0.997 4.548 3.560 0.095 0.171 12.637 0.000
+Survey-Weighted Multiple Reg. 0.997 1.000 0.002 0.122 0.137 0.019 0.902
+IPTW Multiple Reg. 0.997 1.205 0.208 0.053 0.051 0.046 0.030
+IPTW + Survey-Weighted Multiple Reg. 0.997 0.998 0.001 0.122 0.140 0.020 0.898
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.997 1.003 0.005 0.120 0.134 0.018 0.902
+Outcome Modeling and Direct Standardization 0.997 0.992 -0.005 0.110 0.092 0.008 0.978
+Inverse Probability Weighting 1 0.997 1.002 0.005 0.148 0.140 0.020 0.954
+Inverse Probability Weighting 2 0.997 1.026 0.029 0.262 0.332 0.111 0.934
+Augmented Inverse Probability Weighting 0.997 1.001 0.003 0.158 0.153 0.023 0.968
+Setting 8: Both Confounding and Selection Bias Expected ($\tau_{X}=1$, $\beta_{A}=1$, $\beta_{X}=1$)
+Oracle Estimator 1.001 0.991 -0.010 0.097 0.093 0.009 0.956
+Simple Reg. 1.001 2.210 1.207 0.117 0.110 1.474 0.000
+Multiple Reg. 1.001 1.232 0.230 0.073 0.079 0.059 0.126
+IPTW Estimator 1.001 3.877 2.872 0.173 0.172 8.301 0.000
+Survey-Weighted Multiple Reg. 1.001 0.994 -0.007 0.118 0.134 0.018 0.920
+IPTW Multiple Reg. 1.001 1.233 0.231 0.074 0.081 0.060 0.144
+IPTW + Survey-Weighted Multiple Reg. 1.001 1.016 0.015 0.109 0.123 0.015 0.912
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.001 0.998 -0.004 0.122 0.145 0.021 0.912
+Outcome Modeling and Direct Standardization 1.001 0.993 -0.008 0.114 0.095 0.009 0.982
+Inverse Probability Weighting 1 1.001 0.996 -0.005 0.157 0.155 0.024 0.954
+Inverse Probability Weighting 2 1.001 1.017 0.016 0.226 0.264 0.070 0.968
+Augmented Inverse Probability Weighting 1.001 0.992 -0.010 0.154 0.158 0.025 0.946
+4.3 Sensitivity Analyses
+We carry out five sensitivity analyses to better understand the robustness of our approach to the following various real-world situations.
+In the following, we briefly describe the setup of each sensitivity analysis and our key takeaways.
+Full sensitivity analysis results can be found in Appendix B.
+4.3.1 Single-Stage Unequal Probability Sampling
+We compare our estimators to the current standard methods, but with the data drawn from a single-stage unequal probability sample, to examine how the performances of the methods change under different sampling designs.
+In this sensitivity analysis, we set the number of strata and clusters to be $n_{strata}=n_{cluster}=1$ and retain all other parameter settings from our main simulation.
+When the data are generated under a single-stage unequal probability sample, all estimators show reduced Monte Carlo variability compared to our main, stratified clustered complex design setting, but their relative performances remain the same.
+The unweighted and survey-weighted regression methods are unbiased across all eight scenarios, with analytic SEs matching MCSEs and coverage near 95% in the earlier settings, with performance decreasing under settings with confounding and/or selection bias.
+In contrast, the four proposed methods have low bias, within $\pm$ 1% of the truth, ASEs closely aligned with the MCSEs, and coverages of 94-98% across all settings.
+Outcome Modeling often achieves the lowest ASE and MCSE, while IPW1 and IPW2 remain robust with minimal bias and nominal coverage, and AIPW combines these results to offer reliable inference even when either the propensity or outcome model alone is correctly specified.
+Full results are given in Supplemental Table B.1.
+4.3.2 Propensity/Outcome Model Misspecification
+In a second sensitivity analysis, we study our proposed approaches under increasing misspecification of either/both the propensity and outcome models.
+We consider nine scenarios by crossing three levels of propensity score model misspecification (via the interaction effect, $\tau_{X12}\in\{0,0.1,0.5\}$) with three levels of outcome model misspecification (via the group heterogeneity effect, $\alpha_{AX}\in\{0,0.1,0.5\}$).
+In each case, we estimate only main effects models, thereby omitting the true interaction terms when they are non-zero.
+The results of this sensitivity analysis show that outcome modeling alone is vulnerable to both mild and moderate outcome model misspecification, as expected.
+Further, IPW1 is fairly robust to moderate propensity score misspecification but loses efficiency, while IPW2 is unstable under propensity score misspecification.
+The AIPW estimator remains the only estimator with low bias in most settings, reasonable variance, and coverage closest to 95% across the fully factorial design.
+Specifically, when neither model is misspecified, all four proposed methods match the oracle’s performance (bias $\approx 0$, ASE $\approx$ MCSE, and coverage at the nominal level), with outcome modeling/direct standardization being most efficient, and AIPW and IPW1 being slightly more efficient than IPW2.
+Under mild outcome model misspecification, outcome modeling incurs 17% bias and has only 68% coverage, while IPW1, IPW2, and AIPW are unbiased and having nominal coverage.
+These results are magnified in the moderate outcome misspecification setting.
+When only the propensity model is mildly or moderately misspecified, outcome modeling and AIPW remain unbiased, and IPW1 is fairly robust, but IPW2 shows large bias under moderate misspecification (-20%) despite nominal coverage (94%).
+In scenarios where both models are mildly or mixed misspecified, AIPW consistently has bias near 2-3% and coverage $\geq$ 88%, IPW1 maintains < 3% bias and $\geq$ 91% coverage, whereas outcome modeling and IPW2 break down under moderate misspecification.
+Finally, when both models are moderately misspecified, AIPW’s bias is 7.9% (coverage 83%) and IPW1’s bias is 6.1% (coverage 85%), while outcome modeling and IPW2 have bias > 20%.
+Overall, the doubly robust approach offers the best balance of bias control and inference validity across all misspecification patterns, as expected.
+Full results are given in Supplemental Table B.2.
+4.3.3 Limited Overlap
+In practice, the positivity assumption may be violated due to limited overlap in propensity scores between groups.
+To study the impact of this violation, we conduct a sensitivity analysis exploring scenarios with different degrees of overlap.
+Following Zeng et al.
+(2025), we tune the parameters in the propensity model generating mechanism to simulate propensity scores with good versus poor overlap.
+Specifically, we vary $\tau_{A}\in\{0.5,2.0\}$, in the true propensity model to emulate settings with good versus poor overlap in the propensity scores, respectively.
+Figure 2 illustrates the distributions of the propensity scores under these scenarios.
+Under good overlap ($\tau_{A}=0.5$), all four proposed methods yield unbiased ACD estimates and nominal coverage.
+Outcome modeling has the smallest bias (-0.2%) and ASE (0.113), while IPW1 and IPW2 have modestly higher bias (0.7% and 2.8%) and (0.153, 0.213).
+AIPW balances these outcomes, with 0.8% bias, ASE = 0.150, and 96.8% coverage.
+When overlap is poor ($\tau_{A}=2.0$), outcome modeling remains unbiased (1.4%), has low ASE (0.129) that aligns with its MCSE (0.109), and maintains nominal coverage (97.2%).
+Similarly, AIPW maintains low bias (-1.0%) and nominal coverage (96.2%).
+In contrast, IPW1 and IPW2 have higher bias (-9.5% and 2.5%), underestimated standard errors (ASE of 0.542 and 0.352 versus and MCSE of 0.966 and 0.714, respectively), reducing their coverage to 91.8% and 90.6%.
+These results suggest that while the IPW-based estimators are affected by poorer overlap, outcome modeling and AIPW are robust and provide more stable performance in terms of bias, variability, and coverage.
+For the full simulation results, see Supplemental Table B.3.
+Figure 2: Distribution of estimated propensity scores by group membership under two overlap scenarios.
+A.
+Good Overlap $(\tau_{A}=0.5)$.
+Propensity scores for the comparison group (purple, $A=0$) and group of interest (teal, $A=1$) show good overlap around 0.5, supporting stable inverse-probability weighting.
+B.
+Poor Overlap $(\tau_{A}=2.0)$.
+Propensity scores concentrate near 0 for the comparison group and near 1 for the group of interest, indicating limited common support.
+4.3.4 Estimated Selection Mechanism
+As discussed in Section 3.4.2, if sample selection mechanism is not known, we need to model $\Pr(S\mid A,\boldsymbol{X})$.
+In this supplementary analysis, we treat the probability of selection given our variable of interest and our confounders as unknown and estimate $\Pr(S\mid A,\boldsymbol{X})$ via a beta regression, as recommended by Elliot (2009).
+Note that this deviates from our motivating example, but is relevant for applications with non-probability samples.
+Here, we replicate all the settings of our main simulation, but for our proposed methods we treat the selection probabilities as unknown and estimate them.
+These results show that our proposed approaches still have low bias and again are the only methods to achieve nominal coverage in our setting of interest.
+Specifically, outcome modeling remains nearly unbiased (relative bias within $\pm$ 1%) with ASEs that align with the Monte Carlo SEs and 94-96% coverage.
+Both IPW1 and IPW2 also maintain minimal bias ($\leq$ 1%) and nominal coverage (92-96%), although their Monte Carlo variability is occasionally slightly larger than their ASEs in more complex settings.
+AIPW delivers the most consistent performance, combining low bias ($\pm$ 1%) with ASEs within 10% of the MCSEs and coverage around 95% in every scenario.
+Moreover, while the standard survey-weighted regressions also achieve low bias, they all have coverage $\leq$ 90.4% in our setting of interest.
+Full results are given in Supplemental Table B.4.
+4.3.5 Sampling Covariate Omission
+Lastly, we extend our main simulation by introducing two additional, unobserved covariates, $U_{1}$ and $U_{2}$, that vary by stratum and cluster and drive the selection mechanism, but do not enter the true propensity or outcome models.
+In generating each sample, we first simulate $U_{1}$ and $U_{2}$ with stratum- and cluster-specific means and variances, then let the probability of inclusion depend on $A$, $X_{1}$, $X_{2}$, $U_{1}$, and $U_{2}$, with coefficients $\beta_{U_{1}}=\beta_{U_{2}}=1$ across all settings.
+All other parameter settings followed from the main simulation.
+In our analysis, however, we treat the selection mechanism as unknown and estimate $\Pr(S=1\mid A,\boldsymbol{X})$, omitting $U_{1}$ and $U_{2}$ to understand if this induces bias.
+We then apply our four proposed estimators using the estimated selection model, and compare their performance to the oracle estimator and the existing methods.
+Despite this misspecification, all four proposed methods retain low bias and nominal coverage in most settings.
+Outcome modeling incurs a modest negative bias (ranging from –3.3% to –5.7% across all scenarios), but it remains efficient (MCSE from 0.06 to 0.12) and maintains coverage at the nominal level.
+IPW1, which uses a single estimated selection model, shows slightly larger bias (-1.7% to -10.6%) with coverage generally between 86% and 96%.
+IPW2, which estimates more models, exhibits bias within $\pm$ 3.6% in most scenarios and preserves coverage around 95-97%, though its variance inflation is substantial in some settings.
+Lastly, AIPW strikes the best balance, with bias remaining under 8.1% (-8.1% to 6.7%), ASEs aligning closely with the MCSEs, and coverage between 90.8% and 95.4%.
+Overall, even when key design covariates are omitted from the estimated selection model, the proposed approaches have reasonable performance, with the AIPW estimator being particularly robust to this form of misspecification.
+Full results are given in Supplemental Table B.5.
 4.4 Summary Takeaways
-In summary, the simulation settings described above are designed to vary the associations between $Y$, $A$, $X$, and $S$ with increasing complexity, in order to introduce either/both confounding and/or selection bias into the data generating mechanism.
-In settings with neither confounding nor selection bias, all existing and proposed methods perform comparably well, estimating the ACD with low bias and MSE, and nominal coverage.
-In settings of confounding bias (without selection bias), the methods which account for both $A$ and $X$, namely all of the multiple regression and IPTW-based estimators, tend to perform comparatively well, whereas methods that only model $A$ suffer from high bias and negligible coverage.
-When $X$ is introduced into the generation of either/both $S$ and $A$, that is, when selection or selection + confounding bias exists, we see that standard analytic approaches lead to biased estimates and do not attain nominal coverage, whereas the proposed methods perform empirically well, supporting theoretical results derived in the previous section.
-Lastly, our sensitivity analysis reveals that the proposed methods which rely on contrasts and inverse probability weighting, rather than outcome modeling and direct standardization, are more robust to misspecification, particularly as the degree of effect heterogeneity between the groups of interest increases.
+In summary, the simulation and sensitivity analyses described above vary the associations between $Y$, AA, $\boldsymbol{X}$, $\boldsymbol{U}$, and $S$ to introduce varying degrees of confounding and/or selection bias into the data generating mechanism.
+In our main simulation, across all eight simulation scenarios, the drivers of performance are (1) whether the estimator properly adjusts for the source(s) of bias (confounding, selection, or both), and (2) whether its variance calculation accounts for the stratified, clustered survey design and any weight variability.
+In settings without confounding or selection bias, most existing and proposed methods perform similarly well, providing estimates of the ACD with low bias, low MSE, and nominal coverage.
+In settings with confounding bias (but no selection bias), methods that account for both $A$ and $\boldsymbol{X}$, such as multiple regression and IPTW-based estimators, tend to perform well.
+Methods that only model $A$ often show high bias and negligible coverage.
+When $\boldsymbol{X}$ is introduced into the generation of either/both $S$ and $A$, that is, when selection or selection and confounding bias exists, we see that standard analytic approaches lead to biased estimates and do not attain nominal coverage, whereas the proposed methods perform empirically well, supporting theoretical results derived in the previous section.
+Our sensitivity analyses reveal several key points that have implications in practice.
+First, in settings of good propensity score overlap, all proposed methods are unbiased with nominal coverage, outcome modeling is most efficient, IPW2 most variable, and AIPW intermediate.
+In poor overlap scenarios, outcome modeling and AIPW remain relatively stable, while IPW1 and IPW2 have larger bias or variance.
+Researchers should examine the degree of overlap and possible violations of the positivity assumption when conducting analyses.
+Second, when either the outcome or propensity model alone is mildly or moderately misspecified, AIPW and IPW1/IPW2 maintain low bias and nominal coverage.
+In contrast, outcome modeling incurs bias and low coverage under outcome model misspecification, and IPW2 can be unstable under propensity model misspecification.
+When both models are misspecified, AIPW is the most robust.
+Third, when the selection mechanism is unknown, estimating $\Pr(S=1\mid A,\boldsymbol{X})$ may lead to slight increases in bias and under coverage for IPW1/IPW2, but AIPW remains robust.
+Lastly, even when key sampling covariates are omitted from the estimated selection model, IPW1, IPW2, and AIPW maintains lower bias and good coverage, while outcome modeling shows modest negative bias but retains high coverage due to conservative SEs.
+Thus, in real-world survey analyses, where both confounding and selection mechanisms may be complex or partially unknown, overlap can be limited, and model misspecification is a concern, the doubly robust augmented IPW estimator provides the best balance of bias protection, valid variance estimation under complex sampling, and efficiency.
+We therefore recommend our proposed AIPW estimator as the default analytic approach for estimating controlled differences or causal contrasts in complex survey settings such as the one described in our motivating example.
 5 NHANES: Race, SES, and Telomere Length
-We now turn our attention to our motivating example, which comes from the National Health and Nutrition Examination Survey (NHANES).
+We now present our motivating example, which comes from NHANES.
 NHANES is a nationally representative survey sample collected by the Centers for Disease Control, which contains a wealth of demographic, dietary, physical examination, laboratory, and questionnaire data on individuals.
-Our variable of interest is the survey participant’s self-reported race.
+Our group variable of interest is the survey participant’s self-reported race.
 Racial groups are categories of people treated as a social entity by virtue of physical characteristics.
@@
 Telomeres are regions of DNA at the ends of chromosomes that protect against cellular aging and senescence (Lin et al., 2016).
-Telomere shortening is associated with disease incidence and a variety of health outcomes (Shammas, 2011).
-Variation in telomere lengths is attributable to age and shortening during mitosis, in addition to genetics, social history, exposure to environmental contaminants, and psycho-social stressors (Jiang et al., 2007; Blackburn et al., 2015).
+Telomere shortening is associated with disease incidence and many health outcomes (Shammas, 2011).
+Variation in telomere lengths is attributable to age and shortening during mitosis, in addition to genetics, social history, exposure to environmental contaminants, and psychosocial stressors (Jiang et al., 2007; Blackburn et al., 2015).
 Paradoxically, Black individuals typically have longer telomeres than White individuals (Needham et al., 2013; Venkateswaran et al., 2024), with plausible explanations citing leukocyte cell composition, population stratification, or genetic differences (Freedman et al., 1997; Lin et al., 2016; Hansen et al., 2016).
 Others posit that environmental exposures associated with longer telomere length underlie this so-called ‘race effect,’ as these exposures differ significantly between Blacks and Whites (Needham et al., 2020; Roberts et al., 2022).
-Further, studies have found telomere length to be comparable between Black and White individuals in socio-economically homogeneous populations (Geronimus et al., 2015), calling into question purely genetic explanations.
-Our motivating question is this: if we could hypothetically balance socioeconomic status (SES; $X$) between Black and White individuals ($A$) in a nationally representative sample ($S$), would we still see significant Black/White differences in telomere length ($Y$)?
-Figure 2: Simulation setup and results.
-Columns denote each of eight simulation settings.
-The top row gives the assumed relationships between variables for each setting.
-The second row gives the parameter settings for each data generating mechanism.
-The third row gives the average percent bias and Monte Carlo standard errors for each method across 200 simulation replicates.
-The fourth row gives the coverage probabilities for each method.
+Further, studies have found telomere length to be comparable between Black and White individuals in socioeconomically homogeneous populations (Geronimus et al., 2015), calling into question purely genetic explanations.
+Our motivating question is this: if we could hypothetically balance SES ($\boldsymbol{X}$) between Black and White individuals ($A$) in a nationally representative sample ($S$), would we still see significant Black/White differences in telomere length ($Y$)?
 5.2 NHANES Sampling Design
@@
 Counties are stratified and selected with probability proportional to their weighted average population counts, where the weights give a higher probability of selection to counties with a higher proportion of underrepresented subgroups (e.g., certain age groups, marginalized racial groups, lower-income households).
-In the second stage, the PSUs are further divided (e.g., into city blocks) and sub-sampled, and in the third stage, households are sampled from these sample segments.
+In the second stage, the PSUs are further divided (e.g., into city blocks) and subsampled, and in the third stage, households are sampled from these sample segments.
 Again, households within segments that have a greater proportion of underrepresented age, ethnic, or income groups are given additional weight for being selected.
 Lastly, individuals within households are sampled according to designated age-sex-race/ethnicity screening subdomains.
-This complex sampling scheme motivates our proposed methods, as sample selection depends on self-reported race and potential confounders (i.e., socioeconomic status).
+This complex sampling scheme motivates our proposed methods, as sample selection depends on self-reported race and potential confounders (i.e., SES).
 5.3 Study Sample
@@
 As discussed previously, race was our variable of interest.
-We further adjusted our models for study participant age, sex, and blood cell composition, including complete blood count and five part differential consisting of the proportion of neutrophils, eosinophils, basophils, lymphocytes, and monocytes, to account for known differences in these factors (Freedman et al., 1997), as well as twelve indicators of SES, collectively representing our confounding variables $X$.
+We further adjusted our models for study participant age, sex, and blood cell composition, including complete blood count and five part differential consisting of the proportion of neutrophils, eosinophils, basophils, lymphocytes, and monocytes, to account for known differences in these factors (Freedman et al., 1997), as well as twelve indicators of SES, collectively representing our confounding variables $\boldsymbol{X}$.
 Ten of these, namely marital status, education level, household income, insurance status, Special Supplemental Nutrition Program for Women, Infants, and Children (WIC) usage, household size, home ownership, home type, food security status, and an individual’s poverty income ratio (PIR), were extracted directly from the NHANES demographic and occupation questionnaires.
-Occupation category was constructed by mapping occupation group codes in the NHANES 1999-2002 occupation questionnaire to the national statistics socioeconomic job classifications, and employment status was derived from three occupational measures: type of work done last week, hours worked last week at all jobs, and main reason for not working last week (Rehkopf et al., 2008; Rose et al., 2005).
+Occupation category was constructed by mapping occupation group codes in the NHANES occupation questionnaire to the national statistics socioeconomic job classifications, and employment status was derived from three occupational measures: type of work done last week, hours worked last week at all jobs, and main reason for not working last week (Rehkopf et al., 2008; Rose et al., 2005).
 In later modeling, we perform complete case analysis on $n=5,270$ participants after removing 28 individuals without measured blood composition.
-Descriptive statistics for our analytic sample are presented in Table LABEL:tab:nhanesdesc, both overall and stratified by self-reported race.
-As shown, non-Hispanic Black participants had a median telomere length (mean T/S ratio) of 1.05 (i.e., 0.05 times longer than the standard reference DNA), as compared to non-Hispanic White participants with a median telomere length of 0.98.
+Full descriptive statistics for our analytic sample are presented in Supplemental Table C1, both overall and stratified by self-reported race.
+Non-Hispanic Black participants had a median telomere length (mean T/S ratio) of 1.05 (i.e., 0.05 times longer than the standard reference DNA), as compared to non-Hispanic White participants with a median telomere length of 0.98.
 With respect to the identified socioeconomic characteristics, non-Hispanic Black versus White participants differed significantly across all 12 SES indicators, with Black participants having lower educational attainment, lower household income, and lower home ownership rates, as well as higher food insecurity and WIC utilization rates, larger household sizes, and a higher proportion uninsured, unemployed, and below the federal poverty line.
-Further, a higher proportion of Black participants reported not working or working in low blue collar occupations, a higher proportion reported being widowed, divorced, or separated, and a higher proportion reported residing in apartments, all of which suggest lower socio-economic status (Rose et al., 2005).
+Further, a higher proportion of Black participants reported not working or working in low blue collar occupations, a higher proportion reported being widowed, divorced, or separated, and a higher proportion reported residing in apartments, all of which suggest lower socioeconomic status (Rose et al., 2005).
 Based on these factors, we then calculated propensity scores with race as the outcome using unweighted and survey-weighted logistic regressions.
-The results from these models fit are given in supplementary Table LABEL:tab:nhanespropscore.
+The results from these models fit are given in Supplemental Table C2, and plots of the resulting covariate balance statistics are given in Supplemental Figures C1 and C2.
+As shown, the SES covariate balance was improved across all variables after adjustment by both the unweighted and survey-weighted propensity scores, bringing all below a threshold of 0.1 for both the absolute standardized mean differences and Kolmogorov-Smirnov statistics.
 The propensity scores from these models were then carried forward to our later outcome models in order to estimate the ACD of telomere length by race.
@@
 We compared our proposed approaches to six methods that one might consider in practice, each of which successively increase in their respective adjustments for potential confounding and selection bias.
-Namely, we fit a multiple linear regression (Approach 2 in Table 1) of log-transformed telomere length on race, age, sex, the blood composition variables, and the socioeconomic indicators, not accounting for the potential confounding or selection bias in the study.
-We then fit a survey-weighted multiple regression ((Approach 4), and inverse probability of treatment weighted (IPTW) multiple regression (Approach 5), a regression weighted by the product of the within sample propensity score weights and the survey weights (Approach 6), and a regression weighted by the product of the survey-weighted propensity score weights and the survey weights (Approach 7), not unlike our simulation study.
-We omitted the oracle regression, simple linear regression, and the naive g-computation methods, as they were not applicable (as with the oracle regression) or inadvisable (as with the simple linear regression and naive g-computation).
-For the proposed methods, standard errors for the ACD were calculated via Equation (9), whereas for the existing methods, we calculate either model-based or design-based standard errors where appropriate.
+Namely, we fit a multiple linear regression (Approach 2 in Table 1) of log-transformed telomere length on race, age, sex, the blood composition variables, and the socioeconomic indicators.
+We then fit a survey-weighted multiple regression (Approach 4), and inverse probability of treatment weighted (IPTW) multiple regression (Approach 5), a regression weighted by the product of the within sample propensity score weights and the survey weights (Approach 6), and a regression weighted by the product of the survey-weighted propensity score weights and the survey weights (Approach 7), not unlike our simulation study.
+We omitted the oracle regression and simple linear regression, as they were not applicable (as with the oracle regression) or inadvisable (as with the simple linear regression).
+We calculated standard errors for the ACD with the proposed methods via Equation (9), whereas for the existing methods, we calculate model- versus design-based standard errors where appropriate.
 ACD estimates and corresponding 95% confidence intervals across the various analytic approaches are presented in Table 3 below.
@@
 For example, the multiple linear regression had an ACD estimate of 0.0265, corresponding to approximately a 2.7% longer telomere length, on average, among non-Hispanic Black individuals as compared to non-Hispanic White individuals, adjusting for all other factors (95% CI: 0.0106-0.0423).
-In contrast the ACD point estimates from our proposed approaches varied from 0.0132 (IPW2) to 0.0176 (OM), though these average controlled differences were not statistically different from zero.
+In contrast, the ACD point estimates from our proposed approaches varied from 0.0122 (AIPW) to 0.0176 (OM), though these ACD estimates were not statistically different from zero.
 These results suggest that there is a strong confounding relationship between SES and race.
-Moreover, the methods which properly incorporated the NHANES stratified, clustered design tended to have more conservative standard error estimates than those which did not, as evidenced by the wider confidence intervals (Table 3).
-These results align with our analytic and simulation results in two important ways.
+Moreover, the methods which incorporated the NHANES stratified, clustered design tended to have more conservative standard error estimates, as evidenced by the wider confidence intervals (Table 3).
+Table 3: Average controlled difference (ACD) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of race (non-Hispanic Black versus non-Hispanic White participants) on log-transformed telomere length among $n=5,270$ participants in the National Health and Nutrition Examination Survey (NHANES).
+IPTW: Inverse Probability of Treatment Weighting.
+Method ACD Estimate 95% Confidence Interval
+Multiple Regression 0.0265 0.0106, 0.0424
+IPTW Estimator 0.0262 0.0079, 0.0445
+Survey-Weighted Multiple Regression 0.0298 -0.0010, 0.0606
+IPTW Multiple Regression 0.0183 0.0053, 0.0312
+IPTW + Survey-Weighted Multiple Regression 0.0220 -0.0088, 0.0527
+Weighted IPTW + Survey-Weighted Multiple Regression 0.0186 -0.0126, 0.0498
+Proposed Outcome Modeling and Direct Standardization 0.0176 -0.0029, 0.0381
+Proposed Inverse Probability Weighting 1 0.0141 -0.0134, 0.0416
+Proposed Inverse Probability Weighting 2 0.0131 -0.0081, 0.0342
+Proposed Augmented Inverse Probability Weighting 0.0122 -0.0077, 0.0321
+These results align with our analytic and simulation results in two ways.
 Firstly, we show analytically that if our conceptual diagram holds, the proposed approaches will lead to consistent estimates of the ACD, whereas methods which do not account for confounding and/or the sampling design appropriately will incur bias in estimating the ACD.
@@
 Secondly, the methods which respect the complex survey design tend to have wider confidence intervals, appropriately, as individuals within sampling units are correlated.
-5.5 Supplementary Analysis: Lead Exposure, SES and Telomere Length in NHANES
-In a supplementary analysis, we present a different example from NHANES which may be amenable to a causal interpretation (in the sense that counterfactual/potential outcomes and plausibly changing exposures are more meaningful).
+Lastly, in a supplementary analysis, we present a different example from NHANES which may be amenable to a causal interpretation (in the sense that counterfactual/potential outcomes and plausibly changing exposures are more meaningful).
 To illustrate how the proposed methods can, under stronger assumptions, identify the average difference in potential outcome means (population average treatment effect; PATE), we study the effect of exposure to a particular environmental contaminant, lead, on telomere length (Zota et al., 2015).
-Lead exposure is associated with negative health outcomes, such as cardiovascular and kidney diseases (Navas-Acien et al., 2007, 2009), and biological markers of oxidative stress and inflammation (Lin et al., 2016).
-Recent works suggest that lead exposure may be associated with telomere shortening (Lin et al., 2016; Wu et al., 2012; Zota et al., 2015), however, evidence from these studies is mixed and depends on the analytic strategy.
-Thus, further exploring the estimated effect of lead exposure on telomere length with our proposed approaches may be of interest.
-Blood lead levels were measured at the CDC’s National Center for Environmental Health (Atlanta, Georgia) using a simultaneous multi-element atomic absorption spectrometer (SIMAA 6000; PerkinElmer, Norwalk, Connecticut) with Zeeman background correction (Centers for Disease Control and Prevention, 2001a, b).
-The limits of detection (LOD) were 0.3 $\mu$g/dL for blood lead, with < 1% of study participants having blood lead concentrations below the LOD.
-Among these patients, we imputed their blood lead concentrations by LOD divided by the square root of 2 (Hornung and Reed, 1990).
-To define our exposure, we dichotomized LOD-imputed blood lead concentrations at the median $\mu$g/dL to study the difference in potential outcome means between ‘higher’ and ‘lower’ exposure groups.
-We repeated our analysis above, treating the dichotomized lead variable as our exposure $(A)$.
-As before, we fit a propensity score for lead exposure conditional on the twelve socioeconomic indicators $(X)$ described above.
-We further included self-reported race using two strategies – first as a precision covariate along with age and blood composition (i.e., not in $X$) and second, in a sensitivity analysis, as a part of our confounders, $X$.
-The results of this analysis are presented in Supplementary Table C2.
-While the point estimates and findings differed across methods, we found a significant exposure effect for both the traditional and proposed IPW estimators, with an estimated 3.3-5.0% reduction in telomere length.
-For the remaining methods, including the proposed outcome modeling estimator, we failed to reject the null hypothesis.
-Further, there were no appreciable differences in these estimates or the corresponding conclusions regarding significance when comparing our results with versus without including self-reported race in the propensity model (where applicable).
-Table 2: Characteristics of the $n=5,270$ non-Hispanic Black and non-Hispanic White National Health and Nutrition Examination Survey (NHANES) participants in our analytic sample; overall and stratified by self-reported race.
+Please see Supplement C for full details.
+6 Discussion
+There is a growing literature on the use of propensity score weighting techniques in complex survey design settings (Zanutto, 2006; DuGoff et al., 2014; Ridgeway et al., 2015; Austin et al., 2018; Dong et al., 2020; Lenis et al., 2018; Yang et al., 2024; McCaffrey et al., 2024).
+However, a consensus has not been reached for the setting we consider in this manuscript.
+Notably, propensity score methods have been developed and used primarily for simple random samples.
+When analyzing results from more complicated sampling schemes, such as the NHANES clustered, stratified multi-stage sample, one must make the distinction between the population and the sample as the target of inference (Heckman and Vytlacil, 2001).
+This distinction between the target populations and target estimands is crucial, as researchers must be clear in their inferential statements at both the propensity score model and outcome model stages in their analysis.
+This article examines a particular consideration for drawing inference in which sample selection depends on the group variable of interest, motivated from a social epidemiology question regarding racial disparity and socioeconomic confounding in health outcomes.
+This work advances current literature by deriving new identification formulas and conducting new simulations for analyzing the structure of our motivating data.
+The previous works of Ridgeway et al.
+(2015) and Yang et al.
+(2024) studied the question of when to incorporate survey weights in multi-stage, propensity score-based modeling for continuous and binary outcomes, respectively.
+Empirically, the authors show that a two-stage modeling approaches where the survey weights are included in the calculation of the PATE, but not in the outcome model, yielded the best performance compared to other potential analysis strategies.
+However, among these two-stage approaches, including the survey weights in the propensity model versus not yielded comparable empirical performance.
+We note that these conclusions are consistent with other recent papers that study propensity score methods in complex survey designs, and this ambiguity in the context of our specific case is clarified by the factorization of the joint probability of being selected and belonging to the treatment/exposure group.
+Further, our results dovetail with McCaffrey et al.
+(2024) on how and when to incorporate sampling weights into propensity score-based estimators.
+Like them, we show that the decision to weight the propensity score model hinges on the underlying ignorability assumptions, specifically, which covariates drive both treatment (or group membership) and selection.
+In our setting, with discrete comparison groups that influence sample selection, we further demonstrate two valid factorizations of the joint probability of group membership and selection.
+In the first, one estimates a survey weighted propensity score and directly uses the selection weight conditional on $A$ and $\boldsymbol{X}$.
+In the second, one uses an unweighted propensity model and must then marginalize the sample selection probability over $A$.
+This explicit marginalization is critical for identification of the marginal ACD.
+Further, our analytic and numeric results suggest that, in settings where our assumed structure among the variables holds, our proposed methods minimize the bias in estimating the average controlled difference and achieve nominal coverage rates, as compared to standard analytic methods.
+This represents a significant step forward in the capacity to infer disparities, as it explicitly incorporates the complex relationships between race, SES, and sample selection bias, providing a rigorous approach for studying health outcomes in diverse populations.
+Further, we note that targeting the ACD as our estimand is particularly useful in contexts where the ‘exposure’ group membership (in our context, race), is not manipulable.
+Unlike causal targets such as the ATE or ATT, which are rooted in the assumption that an intervention could, even if only hypothetically, be implemented (i.e., ‘no causation without manipulation,’ see Holland, 1986), the ACD is meant to provide a descriptive comparison in observational studies, where the propensity score is used as a covariate balancing tool to facilitate a controlled comparison between exposure groups.
+Mechanically, we note that our estimators take the same form whether the target is the ACD or ATE, with the difference lying in the stronger assumptions needed to claim a causal target.
+These include the stable unit treatment value assumption, which underpins the existence of counterfactual potential outcomes, the addition of a (weak) treatment exchangeability assumption and a modification of the weak selection exchangeability assumption in the context of the potential outcomes.
+The ACD requires only the positivity assumptions and weak selection exchangeability.
+This distinction has been previously articulated in Li and Li (2023), which emphasizes that ACD estimation requires fewer assumptions, even though its formula matches that of the ATE estimator.
+Li and Li (2023) discussed the ACD in the context of studying racial differences in healthcare expenditure.
+We took inspiration from this, and subsequent works, which have developed consensus around this distinction.
+We believe that the context of studying racial disparities presents these particularities in such a way that should be rigorously studied for best practice recommendations, as racial disparities are of increasing interest and relevance in scientific research.
+Moreover, the assumed notion of ‘race’ itself warrants discussion.
+If the scientific question under study seeks to understand the biological underpinnings of a so-called ‘race effect,’ say due to differences in genetic ancestry, then we cannot posit a causal counterfactual framework, as ‘race’ cannot be manipulated or thought of as a ‘cause.’ However, if race is measured as a social construct, we can then discuss altering racialization as a lived experience (Marcellesi, 2013).
+In our motivating example, we are interested in studying differences in telomere length between non-Hispanic Black and non-Hispanic White individuals.
+We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for available socioeconomic status variables in NHANES and after utilizing appropriate propensity score and survey weighting techniques.
+Among our proposed methods, we estimated the ACD in telomere length to be between 1.22% and 1.76%, though we failed to detect a statistically significant difference.
+This contrasts with the results from standard analytic approaches, which varied from 1.81% to 3.0% in this study.
+Needham et al.
+(2013) estimated 4.8% longer telomeres in Black individuals, on average, adjusting for education, income, health behavior, and body mass index.
+Needham et al.
+(2020) and Roberts et al.
+(2022) also suggested evidence of social and environmental factors playing a role in racial disparities in telomere length.
+Taken together, our proposed method contributes to the collection of strategies that can investigate the role of social determinants of health on disparities, and in particular through the use of nationally representative survey data.
+Though we focus on a particular research question within a particular sampling design, these concepts readily extend to other settings where observational data are collected via an underlying sample selection mechanism.
+In our experience, the design does not fundamentally change the results patterns or our conclusions, however, a more systematic/comprehensive exploration of how our results may vary under a wide range of multi-stage sampling designs is important future work.
+More broadly, studying this type of problem is crucial to prevent research from worsening health inequities by inadequately considering confounding factors and biases arising from non-representative selection in databases.
+In particular, other large-scale databases, such as electronic medical records, provide valuable insights into real-world data but may be subject to their own selection biases.
+Areas of interest for future work include electronic health records data with unknown sample selection probabilities and inferential targets with a time-to-event outcome.
+Several conceptual questions about the in-sample or population generalizable propensity score have not yet been reconciled in the literature.
+For example, while this does not appear to be an issue in our data, another consideration may be practical violations of the positivity assumption.
+This may be more likely to happen under the motivating data structure, as we are estimating the joint distribution of ‘treatment’ and selection.
+As we develop this method further in different contexts, we can consider calibration weighting methods for better balance in situations of non-positivity, which should be checked in practice.
+Further, causal principles related to this work may be extended to more complex relational diagrams or sampling designs.
+For instance, as certain biomedical data become inexpensive and easy to collect, but other endpoints are not, a natural question is whether this framework can be extended to two-stage sampling or sequential designs.
+Lastly, we can expand on the utility of the proposed method by studying deviations from the intended sampling design, such as non-response and dropout.
+In conclusion, this work outlines approaches and novel estimands for using propensity score and survey weights in tandem to study controlled outcome differences when the groups under comparison influence sample selection.
+Though the conceptual diagram for this study is rather specific, this setup is common to many settings of health disparities research.
+That is, our proposed methods and investigation to estimate the ACD follow a motivating health disparities research question in which the sampling scheme and corresponding survey weights depend on the target of the propensity score model.
+We fill this gap in the literature by providing recommendation on the form of the appropriate estimators in this setting and provide analytic and numeric evidence of their properties.
+We have shown that when sample selection depends on the comparison groups of interest and/or potential confounders, we must use the direct standardization, propensity weighting, or a doubly robust combination of these strategies, as outlined here, rather than standard analytic approaches.
+It is our hope that this work will not only further our understanding of appropriate survey design considerations in this framework, but also make the task of analyzing such data in similar settings more accessible to researchers.
+Acknowledgments
+We thank Dr.
+Yajuan Si for helpful discussions on this topic.
+This work was supported in part by NIH/NIGMS R35 GM144128 and the Fred Hutchinson Cancer Center J.
+Orin Edson Foundation Endowed Chair (SS), NIH/NIMH DP2MH122405, R01HD107015, and P2C HD042828 (THM), NSF DMS 1712933 and NIH/NCI UG3-CA267907 (BM), and NIH/NIGMS R01GM139926 (XS).
+Appendix A Additional Analytic Results
+A.1 Identification of the Average Controlled Difference
+A.1.1 Setup and Assumptions
+We work under the sample law $\mathcal{Q}=\Pr(\ \cdot\mid S=1)$, rather than the population law, $\mathcal{P}$, as $(Y,A,\boldsymbol{X})$ are observed only when $S=1$.
+We write $e^{\mathcal{Q}}_{a}(\boldsymbol{X})=\mathcal{Q}(A=a\mid\boldsymbol{X})$, $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}[Y\mid A=a,\boldsymbol{X}]$, and let $\Pr(S=1\mid\boldsymbol{X})$ denote the known selection probability marginalized over $A$.
+For $a\in\{0,1\}$, the averaged controlled difference (ACD) is ${\rm ACD}=\mu(1)-\mu(0)$ with
+$\mu(a)=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right],$
+which we will show is equal to $\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]$, with known $w(\boldsymbol{X})=\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}$.
+The assumptions required to estimate $\mu(a)$, and consequently the ACD, are relatively weaker than those required to identify the causal potential outcome means (see below).
+In particular, the following assumptions are sufficient for showing that our functionals, Equations (1), (3), and (4) in the main text, estimate the $\mu(a)$:
+  1. (i)
+    Positivity: $e^{\mathcal{Q}}_{a}(\boldsymbol{X}):=\mathcal{Q}(A=a\mid\boldsymbol{X})>0$ $\mathcal{Q}$-a.s. $\forall a\in\mathcal{A}$
+  2. (ii)
+    Selection Positivity: $\pi_{a}(\boldsymbol{X}):=\Pr(S=1\mid A=a,\boldsymbol{X})>0$ on the support of $A,\boldsymbol{X}$
+  3. (iii)
+    Weak Selection Exchangeability: $\mathbb{E}_{\mathcal{P}}(Y\mid A=a,\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})=g_{a}(\boldsymbol{X})$.
+A.1.2 Change of Measure Under the Sample Law
+For any integrable $h(\boldsymbol{X})$,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})h(\boldsymbol{X})\right]$ $=\int w(\boldsymbol{x})h(\boldsymbol{x})f_{\mathcal{Q}}(\boldsymbol{x})d\boldsymbol{x};\quad\text{(definition of expectation)}$
+$=\int\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{x})}h(\boldsymbol{x})\frac{\Pr(S=1\mid\boldsymbol{x})f_{\mathcal{P}}(\boldsymbol{x})}{\Pr(S=1)}d\boldsymbol{x};\quad\text{(Bayes: }f_{\mathcal{Q}}(\boldsymbol{x})=f_{\mathcal{P}}(\boldsymbol{x}\mid S=1)\text{)}$
+$=\int h(\boldsymbol{x})f_{\mathcal{P}}(\boldsymbol{X})d\boldsymbol{x}$
+$=\mathbb{E}_{\mathcal{P}}\left[h(\boldsymbol{X})\right].$
+Hence $\mathbb{E}_{\mathcal{P}}\left[h(\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})h(\boldsymbol{X})\right]$.
+A.1.3 Outcome Modeling and Direct Standardization
+We first show that our Outcome Modeling and Direct Standardization estimator (OM) identifies $\mu(a)$.
+Assumptions (i) and (ii) ensure that $\mu(a)$ is well-defined.
+Focusing on $\mu(a)=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]$, we have that
+$\mu(a)$ $=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right];\quad$ (definition of $\mu(a)$)
+$=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right];\quad$ (assumption (iii))
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ (change of measure under $\mathcal{Q}$).
+This shows the equivalence between Main Eq. (1) and $\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]$.
+A.1.4 Inverse Probability Weighting 2
+We next show that our Inverse Probability Weighting 2 (IPW2) estimator identifies $\mu(a)$, as this follows from the above.
+Again, assumptions (i) and (ii) ensure that $\mu(a)$ is well-defined.
+Consider $\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right]$.
+By iterated expectation,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\mathbb{E}_{\mathcal{Q}}\left\{\left.\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right|\boldsymbol{X}\right\}\right];\quad$ (law of iterated expectations)
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)Y\right|\boldsymbol{X}\right\}\right];\quad$ $\text{(pulling out }e^{\mathcal{Q}}_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)\mathbb{E}_{\mathcal{Q}}[Y\mid A,\boldsymbol{X}]\right|\boldsymbol{X}\right\}\right];\quad$ (tower property)
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}e^{\mathcal{Q}}_{a}(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ $\text{(since }\mathbb{E}_{\mathcal{Q}}[\mathbb{I}(A=a)\mid\boldsymbol{X}]=e^{\mathcal{Q}}_{a}(\boldsymbol{X}))$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ $\text{(=}\mu(a)\text{ from above).}$
+This shows the equivalence between Main Eq. (4) and $\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]$.
+A.1.5 Inverse Probability Weighting 1
+We now come back to the Inverse Probability Weighting 1 estimator (IPW1).
+We define $e^{\mathcal{P}}_{a}(\boldsymbol{X})=\mathcal{P}(A=a\mid\boldsymbol{X})$, where $\mathcal{P}$ denotes the population law.
+Further, $\pi_{a}(\boldsymbol{X})=\Pr(S=1\mid A=a,\boldsymbol{X})$, so that $\pi(\boldsymbol{X})=\Pr(S=1\mid\boldsymbol{X})=\sum_{a}e^{\mathcal{P}}_{a}(\boldsymbol{X})\pi_{a}(\boldsymbol{X})$.
+As before, $w(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\pi(\boldsymbol{X})}$, and we now also define $w_{a}(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\pi_{a}(\boldsymbol{X})}$.
+By Bayes’ rule,
+$e^{\mathcal{Q}}_{a}(\boldsymbol{X})=\frac{\Pr(S=1\mid A=a,\boldsymbol{X})\mathcal{P}(A=a\mid\boldsymbol{X})}{\Pr(S=1\mid\boldsymbol{X})}=\frac{\pi_{a}(\boldsymbol{X})e^{\mathcal{P}}_{a}(\boldsymbol{X})}{\pi(\boldsymbol{X})}.$
+Consequently,
+$\frac{w(\boldsymbol{X})}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}=\frac{\Pr(S=1)/\pi(\boldsymbol{X})}{\pi_{a}(\boldsymbol{X})e^{\mathcal{P}}_{a}(\boldsymbol{X})/\pi(\boldsymbol{X})}=\frac{\Pr(S=1)}{\pi_{a}(\boldsymbol{X})e^{\mathcal{P}}_{a}(\boldsymbol{X})}=\frac{w_{a}(\boldsymbol{X})}{e^{\mathcal{P}}_{a}(\boldsymbol{X})}.$
+Again, assumptions (i) and (ii) ensure that $\mu(a)$ is well-defined.
+Consider $\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{P}}_{a}(\boldsymbol{X})}Y\right]$.
+By the identity above,
+$\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{P}}_{a}(\boldsymbol{X})}Y\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right];\quad$ (algebraically equivalent)
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ $\text{(=}\mu(a)\text{ from above).}$
+This shows the equivalence between Main Eq. (3) and $\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]$.
+A.2 Identification of the Potential Outcome Means
+Though not applicable to our motivating research question, in many scenarios the group membership variable of interest is a ‘treatment’ or ‘exposure’ in the more traditional sense, and we seek to target a causal quantity.
+To that end, we can adopt a counterfactual framework and, under stronger assumptions, show that the proposed estimators equivalently identify such a parameter.
+We will introduce some additional notation.
+Take $Y^{a}$ to be the random variable representing the potential counterfactual outcome within each group (Rubin, 2005).
+That is, despite only observing one outcome per individual, we know that each study participant had two potential outcomes – the outcome that would be observed if the individual was in comparison group $a=0$, denoted by $Y^{0}$, and the potential outcome that would be observed if the participant instead was in comparison group $a=1$, denoted by $Y^{1}$.
+By considering an individual’s potential outcome values based on their observed and counterfactual group membership, an inherent missing data problem arises.
+Importantly, there are two layers to the missingness mechanism in our setup.
+Firstly, we observe no data when $S=0$.
+Secondly, due to the missingness of counterfactual values, we cannot fully observe all potential outcomes, $Y^{a}$.
+Further, as before, $\mathbb{E}_{\mathcal{Q}}\left[Y^{a}\right]\neq\mathbb{E}_{\mathcal{P}}\left[Y^{a}\right]$ in general, as the sampled study participants may not be representative of the larger target population.
+We can, however, make assumptions to estimate the unobserved, counterfactual outcome in the population.
+With the assumptions of the previous section, the following additional conditions are sufficient to identify the population potential outcome means, $\mathbb{E}_{\mathcal{P}}[Y^{a}]\ \forall a\in\mathcal{A}$:
+  1. (iii’)
+    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]=\mathbb{E}_{\mathcal{Q}}[Y^{a}\mid A=a,\boldsymbol{X}]$, that is, selection bias comes from $A,\boldsymbol{X}$.
+This is implied by, but does not imply, strong generalizability (in distribution), $Y^{a}\perp S\mid A,\boldsymbol{X}$.
+  2. (iv)
+    Stable Unit Treatment Value (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
+  3. (v)
+    Mean (Weak) Exchangeability: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid\boldsymbol{X}]=\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]$, that is, confounding bias comes from $\boldsymbol{X}$.
+This is implied by, but does not imply, strong exchangeability, i.e., $Y^{a}\perp A\mid\boldsymbol{X}$.
+This is thematically similar to Dahabreh et al.
+(2020), who study the generalization of causal inferences from individuals in randomized trials.
+The authors also differentiate between conditional exchangeability in the mean over $A$ in a randomized population and conditional exchangeability in the mean over $S$ in the context of randomized trials.
+A.2.1 Outcome Modeling and Direct Standardization
+As before, $\mu(a)$ is well-defined under (i) and (ii).
+SUTVA underpins the existence of the counterfactual outcomes, $Y^{a}$, and (iv) captures two aspects of SUTVA: ‘no interference,’ which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and ‘no hidden variation,’ which states that the potential outcome is well-defined for each treatment level.
+By (iv), $\mathbb{E}_{\mathcal{Q}}[Y^{a}\mid A=a,\boldsymbol{X}]=g_{a}(\boldsymbol{X})$.
+Further, by (iii’), $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]=g_{a}(\boldsymbol{X})$.
+To identify $\mathbb{E}_{\mathcal{P}}\left[Y^{a}\right]$,
+$\mathbb{E}_{\mathcal{P}}[Y^{a}]$ $=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid\boldsymbol{X})\right];\quad$ (iterated expectation)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid A=a,\boldsymbol{X})\right];\quad$ (v)
+$=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right];\quad$ (iii’)
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ (change of measure)
+This shows the equivalence between Main Eq. (1) and $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+A.2.2 Inverse Probability Weighting 2
+Under assumptions (i) - (v), we further show that the observed data functional for our inverse probability weighted 2 estimator also identifies the potential outcome mean in the target population, $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+As before, $\mu(a)$ is well-defined under (i) and (ii).
+Using the tower property under $\mathcal{Q}$,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\mathbb{E}_{\mathcal{Q}}\left\{\left.\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}Y\right|\boldsymbol{X}\right\}\right];\quad$ $\text{(law of iterated expectations under }\mathcal{Q}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)Y\right|\boldsymbol{X}\right\}\right];\quad$ $\text{(pulling out }e^{\mathcal{Q}}_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)\mathbb{E}_{\mathcal{Q}}[Y\mid A,\boldsymbol{X}]\right|\boldsymbol{X}\right\}\right];\quad$ (tower property)
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\underbrace{\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)\right|\boldsymbol{X}\right\}}_{=e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\underbrace{\mathbb{E}_{\mathcal{Q}}[Y\mid A=a,\boldsymbol{X}]}_{=g_{a}(\boldsymbol{X})}\right];\quad$ $\text{(definitions of }e^{\mathcal{Q}}_{a}\text{ and }g_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ $\text{(algebra; }e^{\mathcal{Q}}_{a}>0\text{ by positivity (i))}$
+$=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right];\quad$ (change of measure)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{Q}}[Y\mid A=a,\boldsymbol{X}]\right];\quad$ $\text{(definition of }g_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{Q}}[Y^{a}\mid A=a,\boldsymbol{X}]\right];\quad$ (iv)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]\right];\quad$ (iii’)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid\boldsymbol{X})\right];\quad$ (v)
+$=\mathbb{E}_{\mathcal{P}}[Y^{a}];\quad$ $\text{(law of iterated expectations under }\mathcal{P}\text{)}.$
+This shows the equivalence between Main Eq. (4) and $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+A.2.3 Inverse Probability Weighting 1
+Lastly, we show that under assumptions (i) - (v), the observed data functional for our inverse probability weighted 1 estimator also identifies the potential outcome mean in the target population, $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+As before, $\mu(a)$ is well-defined under (i) and (ii).
+Using the tower property under $\mathcal{Q}$,
+$\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}Y\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\mathbb{E}_{\mathcal{Q}}\left\{\left.\frac{\mathbb{I}(A=a)}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}Y\right|\boldsymbol{X}\right\}\right];\quad$ $\text{(law of iterated expectations under }\mathcal{Q}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{1}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\mathbb{I}(A=a)Y\mid\boldsymbol{X}\right\}\right];\quad$ $\text{(pulling out }e_{a}^{\mathcal{P}}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{1}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}\mathbb{E}_{\mathcal{Q}}\left\{\mathbb{I}(A=a)g_{a}(\boldsymbol{X})\mid\boldsymbol{X}\right\}\right];\quad$ $\text{(tower property; definition of }g_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathcal{Q}(A=a\mid\boldsymbol{X})}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}g_{a}(\boldsymbol{X})\right];\quad$ $\text{(definition of }\mathcal{Q}(A=a\mid\boldsymbol{X})\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\pi_{a}(\boldsymbol{X})e_{a}^{\mathcal{P}}(\boldsymbol{X})}{\pi(\boldsymbol{X})e_{a}^{\mathcal{P}}(\boldsymbol{X})}g_{a}(\boldsymbol{X})\right];\quad$ $\text{(Bayes under }\mathcal{P}:e_{a}^{Q}=\pi_{a}e_{a}^{\mathcal{P}}/\pi\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[\frac{\Pr(S=1)}{\pi_{a}(\boldsymbol{X})}\cdot\frac{\pi_{a}(\boldsymbol{X})}{\pi(\boldsymbol{X})}g_{a}(\boldsymbol{X})\right];\quad$ $\text{(substituting }w_{a}=\Pr(S=1)/\pi_{a}\text{)}$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right];\quad$ $\text{(since }w=\Pr(S=1)/\pi\text{)}$
+$=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right];$ (change of measure)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{Q}}(Y^{a}\mid A=a,\boldsymbol{X})\right];\quad$ (iv)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid A=a,\boldsymbol{X})\right];\quad$ (iii’)
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid\boldsymbol{X})\right];$ (v)
+$=\mathbb{E}_{\mathcal{P}}[Y^{a}];\quad$ $\text{(law of iterated expectations under }\mathcal{P}\text{).}$
+This shows the equivalence between Main Eq. (3) and $\mathbb{E}_{\mathcal{P}}[Y^{a}]$.
+A.3 Augmented Inverse Probability Weighting
+A.3.1 Efficient Influence Function
+Data Structure: Recall that we observe i.i.d. draws, $\mathcal{O}=(\boldsymbol{X},A,Y)$, from a joint law, $\mathcal{Q}$, where
+$\mathcal{Q}(\cdot)=\Pr(\ \cdot\mid S=1).$
+We will derive the efficient influence function (EIF) for our setting under this conditional (sample) law, $\mathcal{Q}$, which restricts to $S=1$, where $S\in\{0,1\}$ is a selection/observation indicator, as this is where $\boldsymbol{X},A,Y$ are observable.
+Throughout, expectations, $\mathbb{E}(\cdot)$, are taken under $\mathcal{Q}$, unless stated otherwise.
+Nuisance Functions: We further define the following nuisances for the propensity model, $e^{\mathcal{Q}}_{a}(\cdot)$, and the outcome model, $g_{a}(\cdot)$ under $\mathcal{Q}$, as
+$e^{\mathcal{Q}}_{a}(\boldsymbol{x})=\mathcal{Q}(A=a\mid\boldsymbol{X}=\boldsymbol{x}),\quad g_{a}(\boldsymbol{x})=\mathbb{E}_{\mathcal{Q}}[Y\mid A=a,\boldsymbol{X}=\boldsymbol{x}],$
+and we denote the design (i.e., selection) weights by
+$w(\boldsymbol{x})=\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X}=\boldsymbol{x})}=\frac{\Pr(S=1)}{\pi(\boldsymbol{x})}.$
+In practice, $w(\cdot)$ is either known from the design (preferred, as $\boldsymbol{X}$ is not observed for $S=0$) or estimated from external information.
+In our setting, and throughout, we treat $w(\cdot)$ as a known nuisance function.
+Target Parameter: Under the assumption of weak selection exchangeability in the main text,
+$\mu(a)=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right],$
+where $\mathcal{P}(\cdot)=\Pr(\cdot)$ is the ‘full law,’ and the above reflects the Bayes change of measure,
+$f_{\boldsymbol{X}}(\boldsymbol{x})=f_{\boldsymbol{X}\mid S=1}(\boldsymbol{x})\frac{\Pr(S=1)}{\pi(\boldsymbol{x})}.$
+Tangent Space Under $\mathcal{Q}$: Any regular parametric submodel, $Q_{\varepsilon}$, of $\mathcal{Q}$ has score
+$s(\mathcal{O})=s_{\boldsymbol{X}}(\boldsymbol{X})+s_{A\mid\boldsymbol{X}}(A\mid\boldsymbol{X})+s_{Y\mid A,\boldsymbol{X}}(Y\mid A,\boldsymbol{X}),$
+with the usual conditional mean-zero properties:
+$\mathbb{E}_{\mathcal{Q}}[s_{\boldsymbol{X}}(\boldsymbol{X})]=0,\quad\mathbb{E}_{\mathcal{Q}}[s_{A\mid\boldsymbol{X}}(A\mid\boldsymbol{X})\mid\boldsymbol{X}]=0,\quad\mathbb{E}_{\mathcal{Q}}[s_{Y\mid A,\boldsymbol{X}}(Y\mid A,\boldsymbol{X})\mid A,\boldsymbol{X}]=0.$
+Lastly, let $\mathcal{T}_{\mathcal{Q}}$ denote the tangent space (closure in $L^{2}_{0}(Q)$ of finite linear combinations of such scores).
+Pathwise Derivative of $\mu(a)=\mathbb{E}_{\mathcal{Q}}[w(\boldsymbol{X})g_{a}(\boldsymbol{X})]$: Let $Q_{\varepsilon}$ be a submodel with score $s$. We write
+$g_{a,\varepsilon}(\boldsymbol{x})=\mathbb{E}_{\mathcal{Q}_{\varepsilon}}[Y\mid A=a,\boldsymbol{X}=\boldsymbol{x}].$
+Since $w(\cdot)$ is treated as fixed (i.e., design known), the map $\Psi(\mathcal{Q})=\mathbb{E}_{\mathcal{Q}}[w(\boldsymbol{X})g_{a}(\boldsymbol{X})]$ is smooth and
+$\left.\frac{\partial}{\partial\varepsilon}\Psi(\mathcal{Q}_{\varepsilon})\right|_{\varepsilon=0}=\underbrace{\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\dot{g}_{a}(\boldsymbol{X})\right]}_{\text{perturb }p_{Y\mid A,\boldsymbol{X}}}+\underbrace{\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})s_{\boldsymbol{X}}(\boldsymbol{X})\right]}_{\text{perturb }p_{\boldsymbol{X}}},$
+where $\dot{g}_{a}(\boldsymbol{x})=\left.\tfrac{\partial}{\partial\varepsilon}g_{a,\varepsilon}(\boldsymbol{x})\right|_{0}$.
+Contribution from $p_{Y\mid A,\boldsymbol{X}}$: For fixed $\boldsymbol{x}$, the standard identity for differentiating a conditional mean gives
+$\dot{g}_{a}(\boldsymbol{x})=\mathbb{E}_{\mathcal{Q}}\left\{[Y-g_{a}(\boldsymbol{x})]s_{Y\mid A,\boldsymbol{X}}(Y\mid A=a,\boldsymbol{X}=\boldsymbol{x})\mid A=a,\boldsymbol{X}=\boldsymbol{x}\right].$
+Multiplying by $w(\boldsymbol{x})$, integrating over $\boldsymbol{x}$, and using iterated expectations,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\dot{g}_{a}(\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left\{w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{x})}[Y-g_{a}(\boldsymbol{X})]s_{Y\mid A,A}(Y\mid A,\boldsymbol{X})\right\}.$
+Hence the Riesz representer for this component is
+$\phi_{Y}(\mathcal{O})=w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{x})}[Y-g_{a}(\boldsymbol{X})].$
+Contribution from $p_{\boldsymbol{X}}$: The second term is already in the inner product form,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})s_{\boldsymbol{X}}(\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left\{[w(\boldsymbol{X})g_{a}(\boldsymbol{X})-\mu(a)]s_{\boldsymbol{X}}(\boldsymbol{X})\right\},$
+since $\mathbb{E}_{\mathcal{Q}}[s_{\boldsymbol{X}}(\boldsymbol{X})]=0$. Thus the Riesz representer for perturbations of $p_{\boldsymbol{X}}$ is
+$\phi_{\boldsymbol{X}}(\mathcal{O})=w(\boldsymbol{X})g_{a}(\boldsymbol{X})-\mu(a).$
+Orthogonality to $s_{A\mid\boldsymbol{X}}$: We also need $\mathbb{E}_{\mathcal{Q}}[\phi s_{A\mid\boldsymbol{X}}(A\mid\boldsymbol{X})]=0$.
+Using $\mathbb{E}_{\mathcal{Q}}[Y-g_{a}(\boldsymbol{X})\mid A,\boldsymbol{X}]=0$,
+$\mathbb{E}_{\mathcal{Q}}\left\{w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{x})}[Y-g_{a}(\boldsymbol{X})]s_{A\mid\boldsymbol{X}}\right\}=0,$
+and $\mathbb{E}_{\mathcal{Q}}\left\{[w(\boldsymbol{X})g_{a}(\boldsymbol{X})-\mu(a)]s_{A\mid\boldsymbol{X}}\right\}=0$ by $\mathbb{E}_{\mathcal{Q}}[s_{A\mid\boldsymbol{X}}\mid\boldsymbol{X}]=0$.
+Hence the sum is orthogonal to $s_{A\mid\boldsymbol{X}}$.
+EIF under $\mathcal{Q}$: Combining the two components, the efficient influence function under the sample law, $\mathcal{Q}$, is
+$\phi_{a}^{\mathcal{Q}}(\mathcal{O})$ $=\underbrace{w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{x})}[Y-g_{a}(\boldsymbol{X})]}_{\text{conditional\ outcome perturbation}}+\underbrace{\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})-\mu(a)\right]}_{\text{marginal }\boldsymbol{X}\text{ perturbation}}$
+$=w(\boldsymbol{X})\left\{\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{x})}[Y-g_{a}(\boldsymbol{X})]+g_{a}(\boldsymbol{X})\right\}-\mu(a).$
+This $\phi_{a}^{\mathcal{Q}}\in\mathcal{T}_{\mathcal{Q}}$ satisfies $\left.\frac{\partial}{\partial\varepsilon}\Psi(\mathcal{Q}_{\varepsilon})\right|_{0}=\mathbb{E}_{\mathcal{Q}}[\phi_{a}^{\mathcal{Q}}(\mathcal{O})s(\mathcal{O})]$ for every regular submodel $\mathcal{Q}_{\varepsilon}$, and it is the unique element of $\mathcal{T}_{\mathcal{Q}}$ with minimal variance (canonical gradient).
+Relation to a $\mathcal{P}$-law EIF: As $A,\boldsymbol{X},Y$ are only observed when $S=1$, an EIF under the ‘full law,’ $\mathcal{P}$, necessarily lives on the larger observed-data space $\mathcal{D}=(S,S\boldsymbol{X},SA,SY)$.
+Deriving that EIF requires specifying whether $\pi(\boldsymbol{x})=\Pr(S=1\mid\boldsymbol{X}=\boldsymbol{x})$, or $\pi_{a}(\boldsymbol{x})=\Pr(S=1\mid A=a,\boldsymbol{X}=\boldsymbol{x})$, is:
+  • Design-Known (e.g., known survey weights): Here, we can embed the $\mathcal{Q}$-law result in $\mathcal{P}$ via the identity $\mu(a)=\mathbb{E}_{\mathcal{P}}\left[Sw(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]$ and obtain an EIF in $\mathcal{P}$ that reduces, among $S=1$ units, to $\phi_{a}^{\mathcal{Q}}$.
+  • Unknown and Unconstrained: With $A,\boldsymbol{X}$ unobserved when $S=0$, $\pi(\cdot)$ is not estimable from the internal data alone, so the canonical gradient under $\mathcal{P}$ includes the selection tangent component and is therefore model-dependent.
+In this setting, it is both standard and cleaner to work under $\mathcal{Q}$ (where $A,\boldsymbol{X},Y$ are observed) and treat $w(\cdot)$ as external/known.
+Given that we only observe $A,\boldsymbol{X},Y$ when $S=1$ in our design, we base estimation and inference on the $\mathcal{Q}$-law EIF above, which emits the usual doubly-robust estimator and influence function-based standard errors.
+A.3.2 Doubly-Robust Estimator
+Given this form of the EIF, an augmented inverse probability weighted (AIPW; i.e., doubly-robust) estimator is given by
+$\hat{\mu}_{\rm DR}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\left\{w(\boldsymbol{X}_{i})\hat{g}_{a}(\boldsymbol{X}_{i})+w(\boldsymbol{X}_{i})\frac{\mathbb{I}(A_{i}=a)}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}[Y_{i}-\hat{g}_{a}(\boldsymbol{X}_{i})]\right\}$
+$=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\left\{\hat{g}_{a}(\boldsymbol{X}_{i})+\frac{\mathbb{I}(A_{i}=a)}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}[Y_{i}-\hat{g}_{a}(\boldsymbol{X}_{i})]\right\}.$
+Under standard rate/positivity conditions (and with $w$ treated as fixed/known), we have that
+$\sqrt{n}[\hat{\mu}_{\rm DR}(a)-\mu(a)]\overset{d}{\to}\mathcal{N}(0,\mathbb{V}_{\mathcal{Q}}[\phi_{a}^{\mathcal{Q}}]),$
+[…]
+A.3.3 Proof of Double Robustness
+From above, define the doubly-robust estimator under the sample law $\mathcal{Q}=\Pr(\ \cdot\mid S=1)$ as
+$\hat{\mu}_{\mathrm{DR}}(a)=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\left\{\hat{g}_{a}(\boldsymbol{X}_{i})+\frac{\mathbb{I}(A_{i}=a)}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}\left(Y_{i}-\hat{g}_{a}(\boldsymbol{X}_{i})\right)\right\};\quad w(\boldsymbol{X})=\frac{\Pr(S=1)}{\pi(\boldsymbol{X})},$
+where $\hat{g}_{a}$ estimates $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})$, $\hat{e}^{\mathcal{Q}}_{a}$ estimates $e^{\mathcal{Q}}_{a}(\boldsymbol{X})=\mathcal{Q}(A=a\mid\boldsymbol{X})$, and $\pi(\boldsymbol{X})=\Pr(S=1\mid\boldsymbol{X})$ and $\Pr(S=1)$ are known from the design.
+The estimator is doubly robust in the sense that it is consistent for $\mu(a)$ if either (1) the outcome regression, $g_{a}$, is correctly specified, regardless of $e^{\mathcal{Q}}_{a}$, or (2) the sample propensity, $e^{\mathcal{Q}}_{a}$, is correctly specified, regardless of $g_{a}$.
+We assume positivity, so all ratios are well-defined: $e^{\mathcal{Q}}_{a}(\boldsymbol{X})>0$ and $\pi(\boldsymbol{X})>0$ on the support of $(A,\boldsymbol{X})$.
+Case 1. Correct Outcome Model:
+Assume $\hat{g}_{a}(\boldsymbol{X})\xrightarrow{p}g_{a}(\boldsymbol{X})$ and $w(\boldsymbol{X})$ known.
+Allow $\hat{e}^{\mathcal{Q}}_{a}$ to converge to an arbitrary (possibly misspecified) limit, $e^{\mathcal{Q},*}_{a}$.
+Consider the $\mathcal{Q}$-expectation of the summand:
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\left\{g_{a}(\boldsymbol{X})+\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q},*}_{a}(\boldsymbol{X})}\left(Y-g_{a}(\boldsymbol{X})\right)\right\}\right].$
+By the law of iterated expectations under $\mathcal{Q}$, and pulling out functions of $\boldsymbol{X}$,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\left\{g_{a}(\boldsymbol{X})+\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q},*}_{a}(\boldsymbol{X})}\left(Y-g_{a}(\boldsymbol{X})\right)\right\}\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]+\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q},*}_{a}(\boldsymbol{X})}\left(Y-g_{a}(\boldsymbol{X})\right)\right]$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]+\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{1}{e^{\mathcal{Q},*}_{a}(\boldsymbol{X})}\underbrace{\mathbb{E}_{\mathcal{Q}}\left\{\left.\mathbb{I}(A=a)\left(Y-g_{a}(\boldsymbol{X})\right)\right|\boldsymbol{X}\right\}}_{e^{\mathcal{Q}}_{a}(\boldsymbol{X})g_{a}(\boldsymbol{X})-e^{\mathcal{Q}}_{a}(\boldsymbol{X})g_{a}(\boldsymbol{X})=0}\right]$
+$=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right].$
+Using the change-of-measure identity $\mathbb{E}_{\mathcal{Q}}[w(\boldsymbol{X})h(\boldsymbol{X})]=\mathbb{E}_{\mathcal{P}}[h(\boldsymbol{X})]$, we have that
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right].$
+For a causal target, invoke SUTVA (iv) and mean selection exchangeability for potential outcomes (iii’) to equate $g_{a}(\boldsymbol{X})$ with $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]$, and then mean exchangeability (v), so that
+$\mathbb{E}_{\mathcal{P}}[g_{a}(\boldsymbol{X})]$ $=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid A=a,\boldsymbol{X})\right]$
+$=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}_{\mathcal{P}}(Y^{a}\mid\boldsymbol{X})\right]$
+$=\mathbb{E}_{\mathcal{P}}[Y^{a}]$
+$=\mu(a).$
+Thus, with $g_{a}$ correctly specified, $\hat{\mu}_{\mathrm{DR}}(a)$ is consistent, regardless of $e^{\mathcal{Q}}_{a}$.
+Case 2. Correct Propensity (and Selection) Model:
+Now assume $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})\xrightarrow{p}e^{\mathcal{Q}}_{a}(\boldsymbol{X})$ and $\pi(\boldsymbol{X})$ known.
+Let $g_{a}^{*}(\boldsymbol{X})$ denote an arbitrary (possibly misspecified) limit of $\hat{g}_{a}$.
+Then, conditioning on $\boldsymbol{X}$,
+$\mathbb{E}_{\mathcal{Q}}\left[\left.\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left(Y-g_{a}^{*}(\boldsymbol{X})\right)\right|\boldsymbol{X}\right]=\underbrace{\mathbb{E}_{\mathcal{Q}}[Y\mid A=a,\boldsymbol{X}]}_{=g_{a}(\boldsymbol{X})}-g_{a}^{*}(\boldsymbol{X}).$
+Therefore,
+$\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\left\{g_{a}^{*}(\boldsymbol{X})+\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left(Y-g_{a}^{*}(\boldsymbol{X})\right)\right\}\right]$ $=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right]$
+$=\mathbb{E}_{\mathcal{P}}\left[g_{a}(\boldsymbol{X})\right]$
+$=\mu(a),$
+where the last equalities use the same change-of-measure (and, for the causal target, assumptions (iii’)–(v)) as above.
+Hence, with $e^{\mathcal{Q}}_{a}$ correctly specified, $\hat{\mu}_{\mathrm{DR}}(a)$ is consistent, regardless of $g_{a}$.
+Appendix B Additional Simulation Results
+B.1 Sensitivity Analysis: Single-Stage Unequal Probability Sampling
+Table B1: Sensitivity analysis comparing the proposed methods to the oracle estimator and standard analytic methods for calculating the population average controlled difference (ACD) in $Y$ across $A$, generating population-level data from a single stratum and single cluster.
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
+Setting 1: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 1.027 1.027 0.000 0.109 0.106 0.011 0.956
+Simple Reg. 1.027 1.023 -0.005 0.196 0.193 0.037 0.942
+Multiple Reg. 1.027 1.028 0.000 0.109 0.107 0.011 0.952
+IPTW Estimator 1.027 1.030 0.002 0.110 0.275 0.076 0.568
+Survey-Weighted Multiple Reg. 1.027 1.028 0.000 0.109 0.107 0.011 0.962
+IPTW Multiple Reg. 1.027 1.028 0.000 0.109 0.107 0.011 0.950
+IPTW + Survey-Weighted Multiple Reg. 1.027 1.028 0.000 0.109 0.107 0.011 0.962
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.027 1.028 0.000 0.109 0.107 0.011 0.962
+Outcome Modeling and Direct Standardization 1.027 1.028 0.000 0.110 0.107 0.011 0.952
+Inverse Probability Weighting 1 1.027 1.027 0.000 0.110 0.107 0.011 0.964
+Inverse Probability Weighting 2 1.027 1.030 0.003 0.113 0.109 0.012 0.960
+Augmented Inverse Probability Weighting 1.027 1.028 0.000 0.110 0.107 0.011 0.964
+Setting 2: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 0.913 0.919 0.006 0.117 0.115 0.013 0.958
+Simple Reg. 0.913 1.873 1.051 0.192 0.197 0.960 0.002
+Multiple Reg. 0.913 0.919 0.006 0.117 0.115 0.013 0.960
+IPTW Estimator 0.913 0.918 0.005 0.133 0.223 0.050 0.746
+Survey-Weighted Multiple Reg. 0.913 0.918 0.005 0.118 0.116 0.013 0.962
+IPTW Multiple Reg. 0.913 0.917 0.005 0.118 0.118 0.014 0.952
+IPTW + Survey-Weighted Multiple Reg. 0.913 0.917 0.004 0.118 0.118 0.014 0.952
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.913 0.917 0.004 0.118 0.118 0.014 0.952
+Outcome Modeling and Direct Standardization 0.913 0.919 0.006 0.118 0.115 0.013 0.964
+Inverse Probability Weighting 1 0.913 0.917 0.004 0.133 0.132 0.017 0.962
+Inverse Probability Weighting 2 0.913 0.918 0.005 0.134 0.134 0.018 0.950
+Augmented Inverse Probability Weighting 0.913 0.917 0.004 0.120 0.119 0.014 0.954
+Setting 3: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 0.500 0.505 0.011 0.090 0.096 0.009 0.918
+Simple Reg. 0.500 0.498 -0.004 0.165 0.164 0.027 0.950
+Multiple Reg. 0.500 0.506 0.012 0.090 0.096 0.009 0.922
+IPTW Estimator 0.500 -2.991 -6.984 0.168 0.322 12.288 0.000
+Survey-Weighted Multiple Reg. 0.500 0.505 0.010 0.090 0.096 0.009 0.920
+IPTW Multiple Reg. 0.500 0.506 0.012 0.090 0.096 0.009 0.922
+IPTW + Survey-Weighted Multiple Reg. 0.500 0.505 0.010 0.091 0.096 0.009 0.916
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.500 0.505 0.010 0.090 0.096 0.009 0.918
+Outcome Modeling and Direct Standardization 0.500 0.506 0.012 0.091 0.096 0.009 0.922
+Inverse Probability Weighting 1 0.500 0.505 0.010 0.092 0.096 0.009 0.926
+Inverse Probability Weighting 2 0.500 0.505 0.011 0.100 0.101 0.010 0.930
+Augmented Inverse Probability Weighting 0.500 0.505 0.010 0.092 0.096 0.009 0.928
+Setting 4: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.068 1.060 -0.008 0.100 0.100 0.010 0.956
+Simple Reg. 1.068 2.004 0.877 0.167 0.154 0.900 0.000
+Multiple Reg. 1.068 1.079 0.011 0.104 0.104 0.011 0.952
+IPTW Estimator 1.068 2.909 1.725 0.182 0.192 3.428 0.000
+Survey-Weighted Multiple Reg. 1.068 1.061 -0.006 0.101 0.101 0.010 0.956
+IPTW Multiple Reg. 1.068 1.076 0.008 0.105 0.108 0.012 0.946
+IPTW + Survey-Weighted Multiple Reg. 1.068 1.076 0.008 0.105 0.108 0.012 0.944
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.068 1.058 -0.009 0.101 0.103 0.011 0.948
+Outcome Modeling and Direct Standardization 1.068 1.061 -0.007 0.103 0.100 0.010 0.958
+Inverse Probability Weighting 1 1.068 1.064 -0.004 0.127 0.127 0.016 0.946
+Inverse Probability Weighting 2 1.068 1.059 -0.008 0.142 0.143 0.020 0.948
+Augmented Inverse Probability Weighting 1.068 1.058 -0.009 0.105 0.102 0.011 0.954
+Setting 5: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.118 1.115 -0.002 0.064 0.061 0.004 0.968
+Simple Reg. 1.118 1.320 0.181 0.066 0.067 0.046 0.154
+Multiple Reg. 1.118 1.319 0.180 0.037 0.035 0.042 0.000
+IPTW Estimator 1.118 1.312 0.174 0.037 0.188 0.073 0.204
+Survey-Weighted Multiple Reg. 1.118 1.118 0.000 0.087 0.092 0.008 0.948
+IPTW Multiple Reg. 1.118 1.319 0.180 0.037 0.035 0.042 0.000
+IPTW + Survey-Weighted Multiple Reg. 1.118 1.118 0.000 0.087 0.092 0.008 0.948
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.118 1.120 0.002 0.087 0.091 0.008 0.946
+Outcome Modeling and Direct Standardization 1.118 1.116 -0.001 0.076 0.062 0.004 0.990
+Inverse Probability Weighting 1 1.118 1.120 0.002 0.100 0.091 0.008 0.976
+Inverse Probability Weighting 2 1.118 1.117 -0.001 0.125 0.129 0.017 0.960
+Augmented Inverse Probability Weighting 1.118 1.118 0.000 0.103 0.095 0.009 0.976
+Setting 6: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.069 1.052 -0.016 0.085 0.085 0.007 0.938
+Simple Reg. 1.069 2.201 1.059 0.099 0.097 1.292 0.000
+Multiple Reg. 1.069 1.277 0.195 0.061 0.061 0.047 0.080
+IPTW Estimator 1.069 1.221 0.142 0.119 0.304 0.116 0.512
+Survey-Weighted Multiple Reg. 1.069 1.067 -0.001 0.105 0.116 0.013 0.924
+IPTW Multiple Reg. 1.069 1.280 0.198 0.062 0.063 0.049 0.080
+IPTW + Survey-Weighted Multiple Reg. 1.069 1.070 0.001 0.109 0.128 0.016 0.902
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.069 1.073 0.004 0.107 0.125 0.015 0.904
+Outcome Modeling and Direct Standardization 1.069 1.053 -0.015 0.097 0.086 0.008 0.970
+Inverse Probability Weighting 1 1.069 1.046 -0.022 0.138 0.142 0.021 0.940
+Inverse Probability Weighting 2 1.069 1.070 0.001 0.156 0.166 0.027 0.978
+Augmented Inverse Probability Weighting 1.069 1.071 0.002 0.134 0.133 0.018 0.962
+Setting 7: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.187 1.177 -0.009 0.038 0.038 0.002 0.936
+Simple Reg. 1.187 1.303 0.097 0.039 0.037 0.015 0.156
+Multiple Reg. 1.187 1.379 0.161 0.022 0.022 0.037 0.000
+IPTW Estimator 1.187 6.149 4.179 0.052 0.103 24.626 0.000
+Survey-Weighted Multiple Reg. 1.187 1.188 0.000 0.054 0.059 0.003 0.942
+IPTW Multiple Reg. 1.187 1.379 0.161 0.022 0.022 0.037 0.000
+IPTW + Survey-Weighted Multiple Reg. 1.187 1.184 -0.003 0.055 0.061 0.004 0.940
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.187 1.188 0.001 0.053 0.058 0.003 0.940
+Outcome Modeling and Direct Standardization 1.187 1.178 -0.008 0.046 0.039 0.002 0.968
+Inverse Probability Weighting 1 1.187 1.188 0.001 0.061 0.059 0.004 0.962
+Inverse Probability Weighting 2 1.187 1.128 -0.050 0.079 0.082 0.010 0.860
+Augmented Inverse Probability Weighting 1.187 1.188 0.000 0.065 0.063 0.004 0.962
+Setting 8: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.109 1.108 -0.001 0.064 0.062 0.004 0.952
+Simple Reg. 1.109 2.173 0.960 0.080 0.071 1.137 0.000
+Multiple Reg. 1.109 1.305 0.177 0.052 0.053 0.041 0.038
+IPTW Estimator 1.109 4.305 2.883 0.123 0.129 10.231 0.000
+Survey-Weighted Multiple Reg. 1.109 1.108 -0.001 0.074 0.078 0.006 0.946
+IPTW Multiple Reg. 1.109 1.304 0.176 0.053 0.055 0.041 0.048
+IPTW + Survey-Weighted Multiple Reg. 1.109 1.123 0.013 0.070 0.074 0.006 0.932
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.109 1.109 0.000 0.079 0.085 0.007 0.934
+Outcome Modeling and Direct Standardization 1.109 1.108 -0.001 0.074 0.062 0.004 0.974
+Inverse Probability Weighting 1 1.109 1.090 -0.017 0.105 0.108 0.012 0.944
+Inverse Probability Weighting 2 1.109 1.104 -0.004 0.110 0.110 0.012 0.964
+Augmented Inverse Probability Weighting 1.109 1.107 -0.001 0.093 0.088 0.008 0.972
+Reg.: Regression
+B.2 Sensitivity Analysis: Propensity/Outcome Model Misspecification
+Table B2: Sensitivity analysis comparing the proposed methods to the oracle estimator for calculating the population average controlled difference (ACD) in $Y$ across AA for varying degrees of propensity and/or outcome model misspecification.
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Coverage
+Neither Misspecified
+Oracle Estimator 0.997 0.995 -0.002 0.096 0.090 0.008 0.952
+Outcome Modeling and Direct Standardization 0.997 0.991 -0.006 0.100 0.059 0.004 0.990
+Inverse Probability Weighting 1 0.997 0.994 -0.003 0.162 0.144 0.021 0.964
+Inverse Probability Weighting 2 0.997 1.007 0.010 0.219 0.218 0.048 0.968
+Augmented Inverse Probability Weighting 0.997 0.996 -0.001 0.158 0.135 0.018 0.962
+Outcome Model Mildly Misspecified
+Oracle Estimator 1.016 1.024 0.007 0.092 0.095 0.009 0.938
+Outcome Modeling and Direct Standardization 1.016 1.189 0.170 0.104 0.066 0.034 0.678
+Inverse Probability Weighting 1 1.016 1.025 0.008 0.148 0.152 0.023 0.948
+Inverse Probability Weighting 2 1.016 1.048 0.031 0.213 0.228 0.053 0.970
+Augmented Inverse Probability Weighting 1.016 1.029 0.012 0.148 0.154 0.024 0.956
+Outcome Model Moderately Misspecified
+Oracle Estimator 0.875 0.857 -0.020 0.109 0.111 0.013 0.940
+Outcome Modeling and Direct Standardization 0.875 1.783 1.038 0.163 0.085 0.832 0.004
+Inverse Probability Weighting 1 0.875 0.959 0.096 0.171 0.195 0.045 0.886
+Inverse Probability Weighting 2 0.875 0.975 0.114 0.308 0.388 0.161 0.928
+Augmented Inverse Probability Weighting 0.875 0.896 0.024 0.185 0.239 0.057 0.884
+Propensity Model Mildly Misspecified
+Oracle Estimator 1.000 0.990 -0.010 0.102 0.101 0.010 0.960
+Outcome Modeling and Direct Standardization 1.000 1.003 0.003 0.106 0.066 0.004 0.996
+Inverse Probability Weighting 1 1.000 1.013 0.013 0.170 0.167 0.028 0.958
+Inverse Probability Weighting 2 1.000 0.951 -0.049 0.252 0.595 0.355 0.966
+Augmented Inverse Probability Weighting 1.000 0.999 -0.001 0.171 0.197 0.039 0.960
+Propensity Model Moderately Misspecified
+Oracle Estimator 0.999 0.985 -0.014 0.097 0.098 0.010 0.934
+Outcome Modeling and Direct Standardization 0.999 0.999 0.000 0.102 0.070 0.005 0.986
+Inverse Probability Weighting 1 0.999 1.027 0.028 0.160 0.156 0.025 0.966
+Inverse Probability Weighting 2 0.999 0.804 -0.196 0.284 1.039 1.115 0.944
+Augmented Inverse Probability Weighting 0.999 0.975 -0.024 0.192 0.392 0.154 0.962
+Both Mildly Misspecified
+Oracle Estimator 0.998 1.014 0.016 0.100 0.103 0.011 0.936
+Outcome Modeling and Direct Standardization 0.998 1.155 0.158 0.112 0.070 0.030 0.796
+Inverse Probability Weighting 1 0.998 1.022 0.025 0.162 0.164 0.027 0.960
+Inverse Probability Weighting 2 0.998 0.976 -0.021 0.232 0.284 0.081 0.968
+Augmented Inverse Probability Weighting 0.998 1.014 0.017 0.160 0.167 0.028 0.970
+Propensity Model Mildly Misspecified, Outcome Model Moderately Misspecified
+Oracle Estimator 1.099 1.092 -0.006 0.092 0.091 0.008 0.956
+Outcome Modeling and Direct Standardization 1.099 1.832 0.667 0.132 0.073 0.542 0.002
+Inverse Probability Weighting 1 1.099 1.130 0.028 0.139 0.162 0.027 0.918
+Inverse Probability Weighting 2 1.099 1.069 -0.027 0.231 0.344 0.119 0.960
+Augmented Inverse Probability Weighting 1.099 1.127 0.026 0.150 0.197 0.040 0.888
+Propensity Model Moderately Misspecified, Outcome Model Mildly Misspecified
+Oracle Estimator 0.970 0.984 0.015 0.111 0.116 0.014 0.950
+Outcome Modeling and Direct Standardization 0.970 1.117 0.152 0.121 0.077 0.028 0.854
+Inverse Probability Weighting 1 0.970 1.011 0.043 0.181 0.192 0.038 0.964
+Inverse Probability Weighting 2 0.970 0.764 -0.212 0.326 0.666 0.484 0.930
+Augmented Inverse Probability Weighting 0.970 0.988 0.019 0.199 0.279 0.078 0.956
+Both Moderately Misspecified
+Oracle Estimator 1.114 1.119 0.005 0.090 0.090 0.008 0.952
+Outcome Modeling and Direct Standardization 1.114 1.767 0.587 0.127 0.075 0.432 0.000
+Inverse Probability Weighting 1 1.114 1.182 0.061 0.137 0.164 0.031 0.850
+Inverse Probability Weighting 2 1.114 0.895 -0.196 0.305 0.712 0.553 0.942
+Augmented Inverse Probability Weighting 1.114 1.202 0.079 0.159 0.226 0.059 0.830
+B.3 Sensitivity Analysis: Limited Overlap
+Table B3: Sensitivity analysis comparing the proposed methods to the oracle estimator for calculating the population average controlled difference (ACD) in $Y$ across AA for varying degrees of propensity score overlap ($\tau_{A}=0.5$; good overlap versus 2.0; poor overlap).
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
+Setting 1: Good Overlap $(\tau_{A}=0.5)$
+Oracle Estimator 1.010 1.006 -0.004 0.095 0.094 0.009 0.954
+Outcome Modeling and Direct Standardization 1.010 1.008 -0.002 0.113 0.095 0.009 0.982
+Inverse Probability Weighting 1 1.010 1.017 0.007 0.153 0.148 0.022 0.960
+Inverse Probability Weighting 2 1.010 1.038 0.028 0.213 0.223 0.050 0.986
+Augmented Inverse Probability Weighting 1.010 1.018 0.008 0.150 0.144 0.021 0.968
+Setting 2: Poor Overlap $(\tau_{A}=2.0)$
+Oracle Estimator 0.999 1.013 0.013 0.108 0.106 0.011 0.952
+Outcome Modeling and Direct Standardization 0.999 1.014 0.014 0.129 0.109 0.012 0.972
+Inverse Probability Weighting 1 0.999 0.904 -0.095 0.542 0.966 0.940 0.918
+Inverse Probability Weighting 2 0.999 1.024 0.025 0.352 0.714 0.510 0.906
+Augmented Inverse Probability Weighting 0.999 0.990 -0.010 0.223 0.414 0.171 0.962
+B.4 Sensitivity Analysis: Estimated Selection Mechanism
+Table B4: Sensitivity analysis comparing the proposed methods to the oracle estimator and standard analytic methods for calculating the population average controlled difference (ACD) in $Y$ across $A$, treating the selection mechanism as unknown for the proposed methods.
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
+Setting 1: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 1.024 1.026 0.002 0.109 0.109 0.012 0.954
+Simple Reg. 1.024 1.018 -0.006 0.205 0.204 0.042 0.936
+Multiple Reg. 1.024 1.026 0.001 0.109 0.109 0.012 0.954
+IPTW Estimator 1.024 1.022 -0.003 0.111 0.281 0.079 0.576
+Survey-Weighted Multiple Reg. 1.024 1.025 0.001 0.110 0.109 0.012 0.956
+IPTW Multiple Reg. 1.024 1.026 0.001 0.109 0.109 0.012 0.954
+IPTW + Survey-Weighted Multiple Reg. 1.024 1.025 0.001 0.110 0.109 0.012 0.954
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.024 1.025 0.001 0.110 0.109 0.012 0.954
+Outcome Modeling and Direct Standardization 1.024 1.016 -0.009 0.109 0.108 0.012 0.952
+Inverse Probability Weighting 1 1.024 1.016 -0.009 0.110 0.109 0.012 0.948
+Inverse Probability Weighting 2 1.024 1.015 -0.009 0.109 0.108 0.012 0.956
+Augmented Inverse Probability Weighting 1.024 1.016 -0.009 0.109 0.108 0.012 0.952
+Setting 2: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 1.010 1.015 0.005 0.116 0.114 0.013 0.948
+Simple Reg. 1.010 2.114 1.092 0.200 0.184 1.252 0.000
+Multiple Reg. 1.010 1.015 0.005 0.116 0.114 0.013 0.950
+IPTW Estimator 1.010 1.019 0.009 0.137 0.292 0.085 0.676
+Survey-Weighted Multiple Reg. 1.010 1.015 0.005 0.116 0.115 0.013 0.956
+IPTW Multiple Reg. 1.010 1.016 0.006 0.116 0.117 0.014 0.942
+IPTW + Survey-Weighted Multiple Reg. 1.010 1.016 0.006 0.116 0.118 0.014 0.946
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.010 1.016 0.006 0.116 0.118 0.014 0.946
+Outcome Modeling and Direct Standardization 1.010 1.005 -0.005 0.115 0.113 0.013 0.958
+Inverse Probability Weighting 1 1.010 1.011 0.001 0.136 0.141 0.020 0.936
+Inverse Probability Weighting 2 1.010 1.010 0.000 0.135 0.140 0.020 0.946
+Augmented Inverse Probability Weighting 1.010 1.007 -0.004 0.116 0.116 0.013 0.946
+Setting 3: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 0.991 0.996 0.005 0.090 0.090 0.008 0.948
+Simple Reg. 0.991 0.984 -0.007 0.169 0.158 0.025 0.980
+Multiple Reg. 0.991 0.995 0.005 0.090 0.090 0.008 0.952
+IPTW Estimator 0.991 2.298 1.319 0.106 0.180 1.740 0.000
+Survey-Weighted Multiple Reg. 0.991 0.995 0.004 0.090 0.090 0.008 0.954
+IPTW Multiple Reg. 0.991 0.995 0.005 0.090 0.090 0.008 0.954
+IPTW + Survey-Weighted Multiple Reg. 0.991 0.994 0.004 0.091 0.090 0.008 0.956
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.991 0.995 0.004 0.090 0.090 0.008 0.954
+Outcome Modeling and Direct Standardization 0.991 0.985 -0.006 0.092 0.089 0.008 0.958
+Inverse Probability Weighting 1 0.991 0.994 0.003 0.093 0.090 0.008 0.954
+Inverse Probability Weighting 2 0.991 0.984 -0.006 0.092 0.089 0.008 0.958
+Augmented Inverse Probability Weighting 0.991 0.985 -0.006 0.092 0.089 0.008 0.958
+Setting 4: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 0.991 0.991 0.000 0.096 0.098 0.010 0.958
+Simple Reg. 0.991 2.046 1.065 0.164 0.154 1.138 0.000
+Multiple Reg. 0.991 1.015 0.025 0.099 0.101 0.011 0.948
+IPTW Estimator 0.991 2.325 1.347 0.154 0.185 1.814 0.000
+Survey-Weighted Multiple Reg. 0.991 0.991 0.001 0.096 0.099 0.010 0.952
+IPTW Multiple Reg. 0.991 1.013 0.022 0.100 0.103 0.011 0.946
+IPTW + Survey-Weighted Multiple Reg. 0.991 1.013 0.023 0.100 0.104 0.011 0.948
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.991 0.990 -0.001 0.097 0.100 0.010 0.952
+Outcome Modeling and Direct Standardization 0.991 0.980 -0.011 0.096 0.097 0.010 0.950
+Inverse Probability Weighting 1 0.991 0.985 -0.006 0.113 0.114 0.013 0.960
+Inverse Probability Weighting 2 0.991 0.974 -0.017 0.119 0.119 0.014 0.950
+Augmented Inverse Probability Weighting 0.991 0.979 -0.012 0.098 0.099 0.010 0.952
+Setting 5: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 0.981 0.995 0.014 0.121 0.122 0.015 0.952
+Simple Reg. 0.981 1.199 0.222 0.128 0.123 0.063 0.618
+Multiple Reg. 0.981 1.211 0.234 0.067 0.067 0.057 0.074
+IPTW Estimator 0.981 1.200 0.222 0.067 0.274 0.123 0.300
+Survey-Weighted Multiple Reg. 0.981 0.989 0.008 0.154 0.187 0.035 0.896
+IPTW Multiple Reg. 0.981 1.211 0.234 0.067 0.067 0.057 0.074
+IPTW + Survey-Weighted Multiple Reg. 0.981 0.989 0.007 0.153 0.187 0.035 0.896
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.981 0.994 0.013 0.150 0.182 0.033 0.898
+Outcome Modeling and Direct Standardization 0.981 0.964 -0.018 0.138 0.121 0.015 0.958
+Inverse Probability Weighting 1 0.981 0.965 -0.017 0.179 0.180 0.033 0.958
+Inverse Probability Weighting 2 0.981 0.968 -0.014 0.286 0.330 0.109 0.946
+Augmented Inverse Probability Weighting 0.981 0.955 -0.027 0.188 0.194 0.038 0.956
+Setting 6: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.001 0.994 -0.007 0.113 0.112 0.013 0.944
+Simple Reg. 1.001 2.227 1.225 0.126 0.126 1.519 0.000
+Multiple Reg. 1.001 1.221 0.220 0.077 0.078 0.055 0.210
+IPTW Estimator 1.001 1.275 0.274 0.127 0.344 0.193 0.388
+Survey-Weighted Multiple Reg. 1.001 0.998 -0.003 0.142 0.155 0.024 0.908
+IPTW Multiple Reg. 1.001 1.224 0.222 0.078 0.080 0.056 0.206
+IPTW + Survey-Weighted Multiple Reg. 1.001 0.999 -0.002 0.144 0.171 0.029 0.894
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.001 1.006 0.005 0.141 0.162 0.026 0.900
+Outcome Modeling and Direct Standardization 1.001 0.972 -0.029 0.129 0.112 0.013 0.958
+Inverse Probability Weighting 1 1.001 0.948 -0.053 0.179 0.183 0.036 0.920
+Inverse Probability Weighting 2 1.001 0.985 -0.016 0.246 0.327 0.107 0.974
+Augmented Inverse Probability Weighting 1.001 0.971 -0.030 0.179 0.192 0.038 0.950
+Setting 7: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 0.995 1.003 0.008 0.096 0.094 0.009 0.942
+Simple Reg. 0.995 1.179 0.185 0.099 0.095 0.043 0.550
+Multiple Reg. 0.995 1.202 0.208 0.054 0.053 0.045 0.026
+IPTW Estimator 0.995 4.551 3.572 0.097 0.168 12.669 0.000
+Survey-Weighted Multiple Reg. 0.995 0.996 0.001 0.125 0.145 0.021 0.898
+IPTW Multiple Reg. 0.995 1.202 0.208 0.054 0.053 0.045 0.026
+IPTW + Survey-Weighted Multiple Reg. 0.995 0.997 0.002 0.125 0.144 0.021 0.892
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.995 1.000 0.005 0.123 0.141 0.020 0.898
+Outcome Modeling and Direct Standardization 0.995 0.989 -0.007 0.111 0.094 0.009 0.966
+Inverse Probability Weighting 1 0.995 0.987 -0.008 0.146 0.148 0.022 0.952
+Inverse Probability Weighting 2 0.995 0.976 -0.019 0.241 0.273 0.075 0.928
+Augmented Inverse Probability Weighting 0.995 0.980 -0.016 0.154 0.150 0.023 0.954
+Setting 8: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.003 1.000 -0.002 0.097 0.099 0.010 0.930
+Simple Reg. 1.003 2.162 1.157 0.114 0.110 1.357 0.000
+Multiple Reg. 1.003 1.239 0.235 0.072 0.070 0.061 0.088
+IPTW Estimator 1.003 3.820 2.809 0.171 0.165 7.963 0.000
+Survey-Weighted Multiple Reg. 1.003 1.006 0.003 0.115 0.136 0.018 0.904
+IPTW Multiple Reg. 1.003 1.244 0.240 0.073 0.071 0.063 0.080
+IPTW + Survey-Weighted Multiple Reg. 1.003 1.027 0.024 0.106 0.125 0.016 0.900
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.003 1.009 0.006 0.118 0.147 0.022 0.892
+Outcome Modeling and Direct Standardization 1.003 0.970 -0.032 0.109 0.095 0.010 0.956
+Inverse Probability Weighting 1 1.003 0.993 -0.009 0.151 0.159 0.025 0.922
+Inverse Probability Weighting 2 1.003 0.993 -0.010 0.212 0.225 0.051 0.956
+Augmented Inverse Probability Weighting 1.003 0.973 -0.030 0.142 0.149 0.023 0.922
+Reg.: Regression
+B.5 Sensitivity Analysis: Sampling Covariate Omission
+Table B5: Sensitivity analysis comparing the proposed methods to the oracle estimator and standard analytic methods for calculating the population average controlled difference (ACD) in $Y$ across $A$, with the addition of two unmeasured covariates that affect sample selection.
+Results presented are the true ACD (ACD) estimated ACD (Est.
+ACD), relative bias (Rel.
+Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
+Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
+Setting 1: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 0.958 0.954 -0.005 0.060 0.060 0.004 0.966
+Simple Reg. 0.958 0.956 -0.003 0.109 0.107 0.011 0.952
+Multiple Reg. 0.958 0.954 -0.004 0.060 0.061 0.004 0.956
+IPTW Estimator 0.958 0.957 -0.002 0.060 0.121 0.015 0.692
+Survey-Weighted Multiple Reg. 0.958 0.972 0.014 0.139 0.168 0.028 0.896
+IPTW Multiple Reg. 0.958 0.954 -0.004 0.060 0.061 0.004 0.956
+IPTW + Survey-Weighted Multiple Reg. 0.958 0.972 0.014 0.139 0.168 0.028 0.896
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.958 0.972 0.014 0.139 0.167 0.028 0.900
+Outcome Modeling and Direct Standardization 0.958 0.922 -0.038 0.095 0.064 0.005 0.984
+Inverse Probability Weighting 1 0.958 0.942 -0.017 0.162 0.167 0.028 0.948
+Inverse Probability Weighting 2 0.958 0.925 -0.035 0.312 0.311 0.098 0.954
+Augmented Inverse Probability Weighting 0.958 0.939 -0.020 0.167 0.166 0.028 0.954
+Setting 2: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 1.014 1.011 -0.003 0.068 0.067 0.005 0.942
+Simple Reg. 1.014 2.089 1.061 0.117 0.117 1.170 0.000
+Multiple Reg. 1.014 1.025 0.011 0.068 0.067 0.005 0.944
+IPTW Estimator 1.014 1.027 0.013 0.085 0.186 0.035 0.620
+Survey-Weighted Multiple Reg. 1.014 1.012 -0.002 0.165 0.177 0.031 0.928
+IPTW Multiple Reg. 1.014 1.024 0.010 0.069 0.068 0.005 0.944
+IPTW + Survey-Weighted Multiple Reg. 1.014 1.009 -0.005 0.161 0.179 0.032 0.918
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.014 1.011 -0.003 0.161 0.176 0.031 0.922
+Outcome Modeling and Direct Standardization 1.014 0.982 -0.031 0.113 0.069 0.006 0.986
+Inverse Probability Weighting 1 1.014 0.969 -0.044 0.219 0.210 0.046 0.954
+Inverse Probability Weighting 2 1.014 0.987 -0.026 0.440 0.482 0.232 0.958
+Augmented Inverse Probability Weighting 1.014 0.977 -0.036 0.205 0.188 0.037 0.950
+Setting 3: No Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=0$)
+Oracle Estimator 1.024 1.025 0.001 0.044 0.045 0.002 0.946
+Simple Reg. 1.024 1.015 -0.009 0.082 0.077 0.006 0.962
+Multiple Reg. 1.024 1.025 0.001 0.044 0.045 0.002 0.950
+IPTW Estimator 1.024 2.598 1.536 0.054 0.093 2.486 0.000
+Survey-Weighted Multiple Reg. 1.024 1.027 0.002 0.112 0.132 0.017 0.900
+IPTW Multiple Reg. 1.024 1.025 0.001 0.044 0.045 0.002 0.950
+IPTW + Survey-Weighted Multiple Reg. 1.024 1.026 0.002 0.112 0.133 0.018 0.906
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.024 1.027 0.002 0.112 0.132 0.017 0.904
+Outcome Modeling and Direct Standardization 1.024 0.991 -0.033 0.077 0.047 0.003 0.984
+Inverse Probability Weighting 1 1.024 1.011 -0.013 0.132 0.129 0.017 0.960
+Inverse Probability Weighting 2 1.024 1.047 0.022 0.260 0.283 0.080 0.940
+Augmented Inverse Probability Weighting 1.024 0.993 -0.031 0.136 0.134 0.019 0.948
+Setting 4: Confounding Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.002 0.996 -0.006 0.053 0.053 0.003 0.950
+Simple Reg. 1.002 2.013 1.008 0.090 0.086 1.028 0.000
+Multiple Reg. 1.002 1.030 0.028 0.055 0.056 0.004 0.914
+IPTW Estimator 1.002 2.491 1.485 0.087 0.103 2.226 0.000
+Survey-Weighted Multiple Reg. 1.002 0.994 -0.009 0.132 0.157 0.025 0.930
+IPTW Multiple Reg. 1.002 1.030 0.028 0.056 0.058 0.004 0.924
+IPTW + Survey-Weighted Multiple Reg. 1.002 1.011 0.009 0.131 0.159 0.025 0.924
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.002 0.992 -0.010 0.130 0.156 0.024 0.916
+Outcome Modeling and Direct Standardization 1.002 0.959 -0.043 0.086 0.054 0.005 0.984
+Inverse Probability Weighting 1 1.002 0.971 -0.031 0.159 0.161 0.027 0.950
+Inverse Probability Weighting 2 1.002 0.963 -0.039 0.299 0.326 0.108 0.962
+Augmented Inverse Probability Weighting 1.002 0.957 -0.045 0.160 0.164 0.029 0.940
+Setting 5: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 0.978 0.989 0.011 0.072 0.071 0.005 0.956
+Simple Reg. 0.978 1.219 0.246 0.071 0.071 0.063 0.072
+Multiple Reg. 0.978 1.210 0.238 0.039 0.037 0.055 0.000
+IPTW Estimator 0.978 1.185 0.212 0.039 0.146 0.064 0.148
+Survey-Weighted Multiple Reg. 0.978 0.965 -0.013 0.176 0.257 0.066 0.846
+IPTW Multiple Reg. 0.978 1.210 0.238 0.039 0.037 0.055 0.000
+IPTW + Survey-Weighted Multiple Reg. 0.978 0.966 -0.013 0.176 0.257 0.066 0.842
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.978 0.972 -0.007 0.172 0.248 0.062 0.844
+Outcome Modeling and Direct Standardization 0.978 0.923 -0.057 0.120 0.074 0.008 0.982
+Inverse Probability Weighting 1 0.978 0.903 -0.077 0.210 0.230 0.059 0.878
+Inverse Probability Weighting 2 0.978 0.938 -0.041 0.371 0.422 0.179 0.948
+Augmented Inverse Probability Weighting 0.978 0.898 -0.081 0.225 0.255 0.071 0.888
+Setting 6: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=0$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.021 1.032 0.011 0.059 0.061 0.004 0.942
+Simple Reg. 1.021 2.178 1.134 0.064 0.062 1.343 0.000
+Multiple Reg. 1.021 1.231 0.206 0.040 0.037 0.046 0.000
+IPTW Estimator 1.021 1.153 0.129 0.066 0.179 0.049 0.438
+Survey-Weighted Multiple Reg. 1.021 1.010 -0.011 0.151 0.208 0.043 0.874
+IPTW Multiple Reg. 1.021 1.233 0.208 0.040 0.038 0.047 0.000
+IPTW + Survey-Weighted Multiple Reg. 1.021 1.011 -0.010 0.151 0.219 0.048 0.852
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.021 1.017 -0.003 0.148 0.207 0.043 0.860
+Outcome Modeling and Direct Standardization 1.021 0.979 -0.041 0.106 0.062 0.006 0.980
+Inverse Probability Weighting 1 1.021 0.912 -0.106 0.197 0.221 0.061 0.866
+Inverse Probability Weighting 2 1.021 0.983 -0.036 0.319 0.375 0.142 0.970
+Augmented Inverse Probability Weighting 1.021 0.952 -0.067 0.211 0.246 0.065 0.924
+Setting 7: Selection Bias Expected ($\tau_{\boldsymbol{X}}=0$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 0.986 0.991 0.005 0.057 0.057 0.003 0.958
+Simple Reg. 0.986 1.094 0.110 0.057 0.054 0.015 0.536
+Multiple Reg. 0.986 1.231 0.249 0.031 0.029 0.061 0.000
+IPTW Estimator 0.986 4.312 3.374 0.052 0.106 11.076 0.000
+Survey-Weighted Multiple Reg. 0.986 0.986 0.000 0.148 0.191 0.037 0.900
+IPTW Multiple Reg. 0.986 1.231 0.249 0.031 0.029 0.061 0.000
+IPTW + Survey-Weighted Multiple Reg. 0.986 0.980 -0.006 0.151 0.200 0.040 0.882
+Weighted IPTW + Survey-Weighted Multiple Reg. 0.986 0.991 0.005 0.146 0.184 0.034 0.898
+Outcome Modeling and Direct Standardization 0.986 0.951 -0.036 0.100 0.063 0.005 0.988
+Inverse Probability Weighting 1 0.986 0.948 -0.039 0.183 0.179 0.034 0.934
+Inverse Probability Weighting 2 0.986 1.052 0.067 0.368 0.484 0.239 0.966
+Augmented Inverse Probability Weighting 0.986 0.936 -0.051 0.219 0.239 0.060 0.952
+Setting 8: Both Confounding and Selection Bias Expected ($\tau_{\boldsymbol{X}}=1$, $\beta_{A}=1$, $\beta_{\boldsymbol{X}}=1$)
+Oracle Estimator 1.012 1.000 -0.012 0.058 0.057 0.003 0.950
+Simple Reg. 1.012 2.026 1.002 0.071 0.069 1.033 0.000
+Multiple Reg. 1.012 1.201 0.186 0.044 0.045 0.038 0.014
+IPTW Estimator 1.012 3.546 2.504 0.090 0.100 6.430 0.000
+Survey-Weighted Multiple Reg. 1.012 1.000 -0.012 0.145 0.191 0.037 0.890
+IPTW Multiple Reg. 1.012 1.201 0.186 0.045 0.045 0.038 0.014
+IPTW + Survey-Weighted Multiple Reg. 1.012 1.016 0.004 0.137 0.179 0.032 0.892
+Weighted IPTW + Survey-Weighted Multiple Reg. 1.012 1.000 -0.012 0.143 0.190 0.036 0.890
+Outcome Modeling and Direct Standardization 1.012 0.957 -0.055 0.102 0.061 0.007 0.978
+Inverse Probability Weighting 1 1.012 0.932 -0.079 0.183 0.206 0.049 0.914
+Inverse Probability Weighting 2 1.012 1.018 0.006 0.281 0.308 0.094 0.952
+Augmented Inverse Probability Weighting 1.012 0.946 -0.065 0.191 0.202 0.045 0.908
+Reg.: Regression
+Appendix C Additional Data Example Results
+C.1 Main Analysis of Race, SES, and Telomere Length
+Table C1: Characteristics of the $n=5,270$ non-Hispanic Black and non-Hispanic White National Health and Nutrition Examination Survey (NHANES) participants in our analytic sample; overall and stratified by self-reported race.
 Descriptive statistics are reported as median (IQR) for continuous measures or number (%) for categorical measures.
@@
 Characteristic Overall, N = 5,2981 Non-Hispanic White, N = 3,9651 Non-Hispanic Black, N = 1,3331 p-value2
-Telomere Length, Mean T/S 1.00 (0.84, 1.18) 0.98 (0.83, 1.16) 1.05 (0.88, 1.25) $<$0.001
-Age, Years 50 (35, 67) 52 (35, 70) 45 (34, 62) $<$0.001
+Telomere Length, Mean T/S 1.00 (0.84, 1.18) 0.98 (0.83, 1.16) 1.05 (0.88, 1.25) <<0.001
+Age, Years 50 (35, 67) 52 (35, 70) 45 (34, 62) <<0.001
 Sex 0.4
@@
    Refused/Unknown 7 (0.1%) 5 (0.1%) 2 (0.2%)
-Marital Status $<$0.001
+Marital Status <<0.001
    Never Married 754 (14%) 436 (11%) 318 (24%)
@@
    Refused/Unknown 243 (4.6%) 180 (4.5%) 63 (4.7%)
-Household Size $<$0.001
+Household Size <<0.001
    1 Person 882 (17%) 663 (17%) 219 (16%)
@@
    5+ People 741 (14%) 444 (11%) 297 (22%)
-Home Ownership $<$0.001
+Home Ownership <<0.001
    Does Not Own Home 1,551 (29%) 924 (23%) 627 (47%)
@@
    Refused/Unknown 41 (0.8%) 30 (0.8%) 11 (0.8%)
-Home Type $<$0.001
+Home Type <<0.001
    Family Home Detached 3,361 (63%) 2,681 (68%) 680 (51%)
@@
    Refused/Unknown 42 (0.8%) 30 (0.8%) 12 (0.9%)
-Household Income $<$0.001
+Household Income <<0.001
    \$0 - \$20,000 1,076 (20%) 715 (18%) 361 (27%)
@@
    Refused/Unknown 567 (11%) 360 (9.1%) 207 (16%)
-Poverty-Income Ratio $<$0.001
-   $<$ 1.3 1,055 (20%) 643 (16%) 412 (31%)
-   1.3 $\leq$ PIR $<$ 3.5 1,841 (35%) 1,348 (34%) 493 (37%)
+Poverty-Income Ratio <<0.001
+   << 1.3 1,055 (20%) 643 (16%) 412 (31%)
+   1.3 $\leq$ PIR << 3.5 1,841 (35%) 1,348 (34%) 493 (37%)
    $\geq$ 3.5 1,954 (37%) 1,672 (42%) 282 (21%)
@@
    Refused/Unknown 3,094 (58%) 2,335 (59%) 759 (57%)
-Occupation Category $<$0.001
+Occupation Category <<0.001
    No Work 53 (1.0%) 30 (0.8%) 23 (1.7%)
@@
    Refused/Unknown 3,171 (60%) 2,383 (60%) 788 (59%)
-Insurance Status $<$0.001
+Insurance Status <<0.001
    Not Insured 681 (13%) 430 (11%) 251 (19%)
@@
    Refused/Unknown 48 (0.9%) 32 (0.8%) 16 (1.2%)
-Food Security Status $<$0.001
+Food Security Status <<0.001
    Food Secure 4,397 (83%) 3,443 (87%) 954 (72%)
@@
    Refused/Unknown 201 (3.8%) 140 (3.5%) 61 (4.6%)
-WIC Utilization $<$0.001
+WIC Utilization <<0.001
    No WIC 4,919 (93%) 3,775 (95%) 1,144 (86%)
@@
    Refused/Unknown 13 (0.2%) 5 (0.1%) 8 (0.6%)
-White Blood Cell Count, SI 6.90 (5.60, 8.40) 7.10 (5.90, 8.60) 6.30 (5.10, 7.60) $<$0.001
+White Blood Cell Count, SI 6.90 (5.60, 8.40) 7.10 (5.90, 8.60) 6.30 (5.10, 7.60) <<0.001
    Unknown 2 1 1
-Lymphocytes, % 29 (23, 35) 28 (23, 33) 34 (28, 41) $<$0.001
+Lymphocytes, % 29 (23, 35) 28 (23, 33) 34 (28, 41) <<0.001
    Unknown 28 19 9
-Monocytes, % 8.10 (6.80, 9.58) 8.00 (6.80, 9.40) 8.40 (6.90, 9.90) $<$0.001
+Monocytes, % 8.10 (6.80, 9.58) 8.00 (6.80, 9.40) 8.40 (6.90, 9.90) <<0.001
    Unknown 28 19 9
-Segmented Neutrophils, % 59 (53, 65) 61 (55, 66) 54 (46, 61) $<$0.001
+Segmented Neutrophils, % 59 (53, 65) 61 (55, 66) 54 (46, 61) <<0.001
    Unknown 28 19 9
@@
 2Wilcoxon rank sum test; Pearson’s Chi-squared test; Fisher’s exact test
-Table 3: Average controlled difference (ACD) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of race (non-Hispanic Black versus non-Hispanic White participants) on log-transformed telomere length among $n=5,270$ participants in the National Health and Nutrition Examination Survey (NHANES).
-IPTW: Inverse Probability of Treatment Weighting.
-Method ACD Estimate 95% Confidence Interval
-Multiple Regression 0.0265 0.0106, 0.0423
-IPTW Estimator 0.0263 0.0080, 0.0446
-Survey-Weighted Multiple Regression 0.0298 -0.0012, 0.0607
-IPTW Multiple Regression 0.0181 0.0052, 0.0311
-IPTW + Survey-Weighted Multiple Regression 0.0219 -0.0090, 0.0527
-Weighted IPTW + Survey-Weighted Multiple Regression 0.0186 -0.0127, 0.0499
-Outcome Modeling and Direct Standardization 0.0176 -0.0030, 0.0381
-Inverse Probability Weighting 1 0.0150 -0.0151, 0.0451
-Inverse Probability Weighting 2 0.0132 -0.0081, 0.0345
-6 Discussion
-There is a growing literature considering the use of propensity score weighting techniques in complex survey design settings.
-However, a consensus has not been reached for the setting we consider in this manuscript.
-Notably, propensity score methods have been developed and used primarily for simple random samples.
-When analyzing results from more complicated sampling schemes, such as the NHANES clustered, stratified multi-stage sample, one must make the distinction between the population and the sample as the target of inference (Heckman and Vytlacil, 2001).
-This distinction between the target populations and target estimands are important as researchers must be clear in their inferential statements at both the propensity score model and outcome model stages in their analysis.
-Further, this article examines a particular consideration for drawing inference in which sample selection depends on the group variable of interest, motivated from a social epidemiology question regarding racial disparity and socioeconomic confounding in health outcomes.
-Our analytic and numeric results suggest that, in settings where our assumed structure among the variables holds, our proposed methods minimize the bias in estimating the target average controlled difference and achieve nominal coverage rates, as compared to standard analysis methods.
-We believe that the context of studying racial disparities presents these particularities in such a way that should be rigorously studied for best practice recommendations, as racial disparities are important and of increasing relevance in scientific research.
-Moreover the assumed notion of ‘race’ itself warrants discussion.
-If the scientific question under study seeks to understand the biological underpinnings of a so-called ‘race effect,’ say due to differences in genetic ancestry, then we cannot posit a causal counterfactual framework, as ‘race’ cannot be manipulated or thought of as a ‘cause.’ However, if race is measured as a social construct, we can then discuss altering racialization as a lived experience (Marcellesi, 2013).
-In our motivating example, we are interested in studying differences in telomere length between non-Hispanic Black and non-Hispanic White individuals.
-We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for available socioeconomic status variables in NHANES and after utilizing appropriate propensity score and survey weighting techniques.
-Among our proposed methods, we estimated the ACD in telomere length to be between 1.32% and 1.76%, though we failed to detect a statistically significant difference.
-This contrasts with the results from other potential approaches, which varied from 1.81% to 3.0% in this study.
-Needham et al.
-(2013) estimated 4.8% longer telomeres in Black individuals, on average, adjusting for education, income, health behavior, and body mass index.
-Needham et al.
-(2020) and Roberts et al.
-(2022) also suggested evidence of social and environmental factors playing a role in racial disparities in telomere length.
-Taken together, our proposed method contributes to the collection of strategies that can investigate the role of social determinants of health on disparities, and in particular through the use of nationally representative survey data.
-Though we focus on a particular research question and methods for propensity scores within the frame of complex survey designs, these concepts readily extend to other settings where observational data are collected via an underlying sample selection mechanism.
-More broadly, this type of problem is important to study to ensure research does not exacerbate health inequities through inadequate consideration of confounding and biases due to non-representative selection into databases.
-In particular, other large-scale databases, such as electronic medical records, provide valuable insights into real-world data but may be subject to their own selection biases.
-Areas of interest for future work include electronic health records data with unknown sample selection probabilities, non-Gaussian outcomes, and more complex questions that may include affect heterogeneity.
-Several conceptual questions about the in-sample or population generalizable propensity score have not yet been reconciled in the literature.
-Further, causal principles related to this work may be extended to more complex relational diagrams or sampling designs.
-For instance, as certain biomedical data become inexpensive and easy to collect, but other endpoints are not, a natural question is whether this framework can be extended to two-stage sampling or sequential designs.
-Lastly, we can formally expand on the utility of the proposed method by proposing a doubly robust, augmented inverse probability weighted estimator.
-In conclusion, this work outlines approaches and novel estimands for using propensity score and survey weights in tandem to study controlled outcome differences when the groups under comparison influence sample selection.
-Though the conceptual diagram for this study is rather specific, this setup is common to many settings of health disparities research.
-That is, our proposed methods and investigation to estimate the ACD follow a motivating health disparities research question in which the sampling scheme and corresponding survey weights depend on the target of the propensity score model.
-We fill this gap in the literature by providing recommendation on the form of the appropriate estimators in this setting and provide analytic and numeric evidence of their properties.
-We have shown that when sample selection depends on the comparison groups of interest and/or potential confounders, we must use the direct standardization or propensity weighting strategies outlined here, rather than standard analytic approaches.
-It is our hope that this work will not only further our understanding of appropriate survey design considerations in this framework, but also make the task of analyzing such data in similar settings more accessible to researchers.
-Acknowledgments
-We thank Dr.
-Yajuan Si for helpful discussions on this topic.
-This work was supported in part by NIH/NIGMS R35 GM144128 and the Fred Hutchinson Cancer Center J.
-Orin Edson Foundation Endowed Chair (SS), NIH/NIMH DP2MH122405, R01HD107015, and P2C HD042828 (THM), NSF DMS 1712933 and NIH/NCI UG3-CA267907 (BM), and NIH/NIGMS R01GM139926 (XS).
-Appendix A Additional Analytic Results
-A.1 Average Controlled Difference
-The assumptions required to estimate the averaged controlled difference (ACD) are relatively weaker than those required to identify the causal potential outcome means (see below).
-In particular, the following assumptions are sufficient for showing that our functionals (2), (3), and (4) in the main text estimate the ACD:
-  1. (i)
-    Positivity: $\Pr(A=a\mid X=x)>0\ \forall a\in\mathcal{A}$ and every $x$ such that $f_{X}(x)>0$
-  2. (ii)
-    Positivity of Sample Selection: $\Pr(S=1\mid A=a,X=x)>0$ for every $a,x$ such that $f_{A,X}(a,x)>0$
-  3. (iii)
-    Mean (Weak) Generalizability (Selection Exchangeability): $\mathbb{E}[Y\mid A=a,X]=\mathbb{E}[Y\mid A=a,S=1,X]$, that is, selection bias comes from $A,X$.
-This is implied by, but does not imply, strong generalizability (in distribution), $Y\perp S\mid A,X$.
-We proceed by deriving the estimators in Equations (2), (3), and (4) from the main text below. We define the ACD as
-${\rm ACD}=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=1,X\right]-\mathbb{E}\left[Y\mid A=0,X\right]\right].$
-Outcome Modeling and Direct Standardization
-Under assumptions (i) - (iii), we show that the observed data functional, $\mu(a)$, given by
-$\mu(a)=\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right],$
-can be used to estimate the ACD. Assumptions (i) and (ii) come in to ensure that $\mu(a)$ is well defined. Focusing on $\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$, we have that
-$\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$ =$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
-[…]
-[…]
-[…]
-[…]
-$=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right]$ (Main Eq. 2)
-Inverse Probability Weighting
-Under assumptions (i) - (iii), we further show that the observed data functionals for our inverse probability weighted estimators, given by
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$
-also estimate $\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$. Let $C=A\times S=\mathbb{I}(A=1,S=1)$, that is,
-$C=\begin{cases}1;&A=a,S=1\\ 0;&S=0\text{ or }A\neq a,S=1\end{cases}$
-so that
-$\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,X\right]\right]$ =$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
-$=\mathbb{E}_{X}\left[\frac{C\times Y}{\Pr(C=1\mid X)}\right];\quad\text{following the traditional IPW proof}$
-$=\mathbb{E}_{X}\left[\frac{A\times S\times Y}{\Pr(A=a,S=1\mid X)}\right];\quad\text{by definition of }C$
-$=\mathbb{E}_{X}\left[\frac{A\times 1\times Y}{\Pr(A=a,S=1\mid X)}\mid S=1\right]\times\Pr(S=1)+0$
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a,S=1\mid X)}\mid S=1\right]$
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$ (Main Eq. 3)
-or
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$ (Main Eq. 4)
-where again, the observed data functionals are well defined under (i) and (ii).
-A.2 Identification of the Potential Outcome Means
-Though not applicable to our motivating research question, in many scenarios the group membership variable of interest is a ‘treatment’ or ‘exposure’ in the more traditional sense, and we seek to target a causal quantity.
-To that end, we can adopt a counterfactual framework and, under stronger assumptions, show that the proposed estimators equivalently identify such a parameter.
-We introduce some additional notation below.
-Take $Y^{a}$ to be the random variable representing the potential counterfactual outcome within each group (Rubin, 2005).
-That is, despite only observing one outcome per individual, we know that each study participant had two potential outcomes – the outcome that would be observed if the individual was in comparison group $a=0$, denoted by $Y^{0}$, and the potential outcome that would be observed if the participant instead was in comparison group $a=1$, denoted by $Y^{1}$.
-By considering an individual’s potential outcome values based on their observed and counterfactual group membership, an inherent missing data problem arises.
-Importantly, there are two layers to the missingness mechanism in our setup.
-Firstly, we observe no data when $S=0$.
-Secondly, due to the missingness of counterfactual values, we cannot fully observe all potential outcomes, $Y^{a}$.
-Further, as before, $\mathbb{E}\left[Y^{a}\right]\neq\mathbb{E}\left[Y^{a}\mid S=1\right]$ in general, as the sampled study participants may not be representative of the larger target population.
-We can, however, make assumptions to estimate the unobserved, counterfactual outcome in the population.
-With the assumptions of the previous section, the following additional conditions are sufficient to identify the population potential outcome means, $\mathbb{E}[Y^{a}]\ \forall a\in\mathcal{A}$:
-  1. (iii*)
-    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}[Y^{a}\mid A=a,X]=\mathbb{E}[Y^{a}\mid A=a,S=1,X]$, that is, selection bias comes from $A,X$.
-This is implied by, but does not imply, strong generalizability (in distribution), $Y^{a}\perp S\mid A,X$.
-  2. (iv)
-    Stable Unit Treatment Value (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
-  3. (v)
-    Mean (Weak) Exchangeability: $\mathbb{E}[Y^{a}\mid X]=\mathbb{E}[Y^{a}\mid A=a,X]$, that is, confounding bias comes from $X$.
-This is implied by, but does not imply, strong exchangeability, i.e., $Y^{a}\perp A\mid X$.
-This is thematically similar to Dahabreh et al.
-(2020), who study the generalization of causal inferences from individuals in randomized trials.
-The authors also differentiate between conditional exchangeability in the mean over $A$ in a randomized population and conditional exchangeability in the mean over $S$ in the context of randomized trials.
-Outcome Modeling and Direct Standardization
-Under assumptions (i) - (v), we show that the observed data functional, $\mu(a)$, identifies the potential outcome mean in the target population, $\mathbb{E}[Y^{a}]$, where
-$\mu(a)=\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right].$
-As before, $\mu(a)$ is well defined under (i) and (ii).
-The SUTVA, or consistency, assumption underpins the existence of the counterfactual outcomes, $Y^{a}$.
-The compact representation in (iv) captures two aspects of SUTVA – ‘no interference,’ which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and ‘no hidden variation,’ which states that the potential outcome is well-defined for each treatment level, $a$.
-To identify $\mathbb{E}\left[Y^{a}\right]$, we have that
-$\mathbb{E}\left[Y^{a}\right]$ $=\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid X\right]\right];\quad\text{by law of total expectation}$
-=$\stackrel{{\scriptstyle(v)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,X\right]\right]$
-=$\stackrel{{\scriptstyle(iii*)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,S=1,X\right]\right]$
-=$\stackrel{{\scriptstyle(iv)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
-[…]
-[…]
-[…]
-[…]
-$=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right]$ (Main Eq. 2)
-Inverse Probability Weighting
-Under assumptions (i) - (v), we further show that the observed data functionals for our inverse probability weighted estimators, given by
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$
-$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$
-also identify the potential outcome mean in the target population, $\mathbb{E}[Y^{a}]$. Let $C=A\times S=\mathbb{I}(A=1,S=1)$, that is,
-$C=\begin{cases}1;&A=a,S=1\\ 0;&S=0\text{ or }A\neq a,S=1\end{cases}$
-so that
-$\mathbb{E}\left[Y^{a}\right]$ $=\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid X\right]\right];\quad\text{by law of total expectation}$
-=$\stackrel{{\scriptstyle(v)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,X\right]\right]$
-=$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y^{a}\mid A=a,S=1,X\right]\right]$
-=$\stackrel{{\scriptstyle(iv)}}{{=}}\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=a,S=1,X\right]\right]$
-$=\mathbb{E}_{X}\left[\frac{C\times Y}{\Pr(C=1\mid X)}\right];\quad\text{following the traditional IPW proof}$
-$=\mathbb{E}_{X}\left[\frac{A\times S\times Y}{\Pr(A=a,S=1\mid X)}\right];\quad\text{by definition of }C$
-$=\mathbb{E}_{X}\left[\frac{A\times 1\times Y}{\Pr(A=a,S=1\mid X)}\mid S=1\right]\times\Pr(S=1)+0$
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a,S=1\mid X)}\mid S=1\right]$
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right]$ (Main Eq. 3)
-or
-$=\mathbb{E}_{X}\left[\frac{AY\ \Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right],$ (Main Eq. 4)
-where again, the observed data functionals are well defined under (i) and (ii).
-Appendix B Additional Simulation Results
-Table B1: Simulation results comparing different analytic methods for calculating the population average controlled difference (ACD) for $A$ on $Y$ across different data generating mechanism settings.
-Results presented are the estimated ACD, percent bias, analytic standard error (SE), Monte Carlo standard error, mean squared error (MSE) and coverage probability over 200 subsamples for each setting
-Method True ACD Estimated ACD Percent Bias Analytic SE Monte Carlo SE MSE Coverage
-Setting 1: $\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=0$
-Oracle Estimator 1.096 1.101 0.471 0.068 0.069 0.477 0.925
-Simple Regression 1.096 1.098 0.166 0.097 0.105 1.103 0.930
-Multiple Regression 1.096 1.101 0.450 0.068 0.070 0.484 0.915
-IPTW Estimator 1.096 1.101 0.475 0.068 0.070 0.484 0.920
-Survey-Weighted Multiple Regression 1.096 1.102 0.536 0.068 0.069 0.481 0.930
-IPTW Multiple Regression 1.096 1.101 0.453 0.060 0.070 0.484 0.880
-IPTW + Survey-Weighted Multiple Regression 1.096 1.102 0.538 0.060 0.069 0.482 0.875
-Weighted IPTW + Survey-Weighted Multiple Regression 1.096 1.102 0.538 0.060 0.069 0.482 0.875
-Naive G-Computation 1.096 0.992 -9.478 0.030 0.102 2.111 0.290
-Outcome Modeling and Direct Standardization 1.096 1.090 -0.523 0.067 0.069 0.476 0.920
-Inverse Probability Weighting 1 1.096 1.089 -0.647 0.068 0.069 0.480 0.925
-Inverse Probability Weighting 2 1.096 1.091 -0.499 0.067 0.069 0.475 0.920
-Setting 2: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=0$
-Oracle Estimator 1.090 1.079 -0.989 0.066 0.067 0.452 0.930
-Simple Regression 1.090 1.957 79.480 0.083 0.076 75.638 0.000
-Multiple Regression 1.090 1.079 -0.977 0.066 0.067 0.455 0.925
-IPTW Estimator 1.090 1.071 -1.793 0.094 0.107 1.176 0.930
-Survey-Weighted Multiple Regression 1.090 1.080 -0.886 0.066 0.068 0.468 0.925
-IPTW Multiple Regression 1.090 1.077 -1.246 0.060 0.069 0.490 0.885
-IPTW + Survey-Weighted Multiple Regression 1.090 1.078 -1.131 0.060 0.070 0.506 0.890
-Weighted IPTW + Survey-Weighted Multiple Regression 1.090 1.078 -1.132 0.060 0.070 0.506 0.885
-Naive G-Computation 1.090 0.969 -11.125 0.029 0.089 2.268 0.180
-Outcome Modeling and Direct Standardization 1.090 1.069 -1.935 0.065 0.066 0.480 0.930
-Inverse Probability Weighting 1 1.090 1.060 -2.778 0.093 0.106 1.220 0.925
-Inverse Probability Weighting 2 1.090 1.060 -2.742 0.093 0.106 1.204 0.925
-Setting 3: $\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=0$
-Oracle Estimator 1.100 1.096 -0.422 0.050 0.047 0.220 0.960
-Simple Regression 1.100 1.099 -0.138 0.073 0.074 0.547 0.945
-Multiple Regression 1.100 1.096 -0.418 0.050 0.047 0.223 0.965
-IPTW Estimator 1.100 1.096 -0.421 0.050 0.047 0.223 0.965
-Survey-Weighted Multiple Regression 1.100 1.096 -0.417 0.056 0.047 0.222 0.990
-IPTW Multiple Regression 1.100 1.096 -0.417 0.050 0.047 0.223 0.965
-IPTW + Survey-Weighted Multiple Regression 1.100 1.096 -0.416 0.056 0.047 0.222 0.990
-Weighted IPTW + Survey-Weighted Multiple Regression 1.100 1.096 -0.417 0.050 0.047 0.222 0.970
-Naive G-Computation 1.100 0.998 -9.341 0.025 0.073 1.589 0.230
-Outcome Modeling and Direct Standardization 1.100 1.084 -1.495 0.051 0.047 0.242 0.955
-Inverse Probability Weighting 1 1.100 1.098 -0.217 0.051 0.047 0.219 0.965
-Inverse Probability Weighting 2 1.100 1.084 -1.497 0.051 0.047 0.243 0.955
-Setting 4: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$
-Oracle Estimator 1.106 1.102 -0.419 0.055 0.058 0.342 0.905
-Simple Regression 1.106 1.972 78.249 0.070 0.069 75.408 0.000
-Multiple Regression 1.106 1.121 1.377 0.058 0.062 0.404 0.915
-IPTW Estimator 1.106 1.123 1.487 0.103 0.105 1.129 0.940
-Survey-Weighted Multiple Regression 1.106 1.123 1.499 0.050 0.062 0.411 0.890
-IPTW Multiple Regression 1.106 1.116 0.923 0.044 0.065 0.426 0.830
-IPTW + Survey-Weighted Multiple Regression 1.106 1.118 1.045 0.050 0.065 0.430 0.870
-Weighted IPTW + Survey-Weighted Multiple Regression 1.106 1.119 1.116 0.045 0.065 0.434 0.820
-Naive G-Computation 1.106 0.997 -9.867 0.022 0.070 1.680 0.175
-Outcome Modeling and Direct Standardization 1.106 1.090 -1.442 0.055 0.058 0.362 0.925
-Inverse Probability Weighting 1 1.106 1.104 -0.210 0.076 0.081 0.648 0.940
-Inverse Probability Weighting 2 1.106 1.089 -1.526 0.081 0.083 0.706 0.945
-Setting 5: $\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=1$
-Oracle Estimator 1.104 1.108 0.389 0.046 0.048 0.227 0.940
-Simple Regression 1.104 1.209 9.544 0.048 0.045 1.311 0.365
-Multiple Regression 1.104 1.205 9.214 0.034 0.032 1.135 0.110
-IPTW Estimator 1.104 1.205 9.222 0.034 0.032 1.136 0.110
-Survey-Weighted Multiple Regression 1.104 1.205 9.236 0.046 0.045 1.239 0.355
-IPTW Multiple Regression 1.104 1.205 9.214 0.030 0.032 1.135 0.090
-IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.209 0.041 0.045 1.235 0.270
-Weighted IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.182 0.040 0.045 1.225 0.265
-Naive G-Computation 1.104 1.003 -9.119 0.015 0.074 1.559 0.135
-Outcome Modeling and Direct Standardization 1.104 1.090 -1.252 0.048 0.047 0.236 0.935
-Inverse Probability Weighting 1 1.104 1.100 -0.302 0.057 0.054 0.291 0.945
-Inverse Probability Weighting 2 1.104 1.099 -0.448 0.058 0.057 0.325 0.955
-Setting 6: $\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$
-Oracle Estimator 1.096 1.096 -0.018 0.041 0.042 0.176 0.935
-Simple Regression 1.096 2.004 82.816 0.044 0.043 82.577 0.000
-Multiple Regression 1.096 1.184 8.045 0.036 0.036 0.904 0.310
-IPTW Estimator 1.096 1.222 11.486 0.058 0.054 1.874 0.440
-Survey-Weighted Multiple Regression 1.096 1.189 8.485 0.044 0.059 1.207 0.425
-IPTW Multiple Regression 1.096 1.185 8.115 0.030 0.037 0.927 0.220
-IPTW + Survey-Weighted Multiple Regression 1.096 1.182 7.881 0.040 0.048 0.972 0.440
-Weighted IPTW + Survey-Weighted Multiple Regression 1.096 1.182 7.847 0.040 0.047 0.958 0.435
-Naive G-Computation 1.096 0.999 -8.879 0.015 0.065 1.370 0.130
-Outcome Modeling and Direct Standardization 1.096 1.079 -1.514 0.044 0.041 0.197 0.950
-Inverse Probability Weighting 1 1.096 1.083 -1.209 0.066 0.072 0.540 0.940
-Inverse Probability Weighting 2 1.096 1.089 -0.633 0.066 0.070 0.490 0.935
-Setting 7: $\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=1$
-Oracle Estimator 1.102 1.102 0.073 0.035 0.034 0.115 0.975
-Simple Regression 1.102 1.058 -3.992 0.036 0.034 0.309 0.805
-Multiple Regression 1.102 1.175 6.701 0.026 0.025 0.605 0.190
-IPTW Estimator 1.102 1.174 6.580 0.026 0.025 0.587 0.210
-Survey-Weighted Multiple Regression 1.102 1.177 6.809 0.038 0.031 0.661 0.510
-IPTW Multiple Regression 1.102 1.175 6.698 0.026 0.025 0.605 0.190
-IPTW + Survey-Weighted Multiple Regression 1.102 1.176 6.793 0.037 0.032 0.663 0.455
-Weighted IPTW + Survey-Weighted Multiple Regression 1.102 1.177 6.808 0.034 0.031 0.658 0.385
-Naive G-Computation 1.102 1.017 -7.688 0.013 0.054 1.009 0.135
-Outcome Modeling and Direct Standardization 1.102 1.086 -1.405 0.038 0.033 0.135 0.970
-Inverse Probability Weighting 1 1.102 1.100 -0.172 0.043 0.042 0.178 0.970
-Inverse Probability Weighting 2 1.102 1.076 -2.322 0.044 0.045 0.265 0.910
-Setting 8: $\tau_{X}=1$, $\beta_{A}=1$, $\beta_{X}=1$
-Oracle Estimator 1.104 1.101 -0.209 0.034 0.033 0.106 0.970
-Simple Regression 1.104 1.919 73.914 0.040 0.037 66.681 0.000
-Multiple Regression 1.104 1.210 9.640 0.034 0.032 1.234 0.130
-IPTW Estimator 1.104 1.172 6.226 0.070 0.066 0.904 0.835
-Survey-Weighted Multiple Regression 1.104 1.200 8.727 0.033 0.063 1.321 0.300
-IPTW Multiple Regression 1.104 1.212 9.830 0.022 0.031 1.273 0.015
-IPTW + Survey-Weighted Multiple Regression 1.104 1.207 9.378 0.031 0.034 1.188 0.130
-Weighted IPTW + Survey-Weighted Multiple Regression 1.104 1.205 9.214 0.030 0.041 1.204 0.160
-Naive G-Computation 1.104 0.986 -10.660 0.011 0.053 1.669 0.040
-Outcome Modeling and Direct Standardization 1.104 1.087 -1.531 0.038 0.032 0.131 0.970
-Inverse Probability Weighting 1 1.104 1.115 1.015 0.050 0.051 0.274 0.950
-Inverse Probability Weighting 2 1.104 1.083 -1.907 0.047 0.047 0.267 0.935
-Table B2: Sensitivity analysis simulation results comparing the proposed methods to the oracle estimator in terms of bias, mean squared error, and coverage across different treatment heterogeneity levels (size of $\gamma_{AX}$ in the true data generating mechanism.
-Results are averaged over 200 independent sub-samplings from a super population of 100,000 individuals where all other simulation settings match that of Setting 8 in the main text.
-Results for OM are presented where the outcome model is correctly specified versus misspecified.
-ACD: Average Controlled Difference; SE: Standard Error; MSE: Mean Squared Error.
-Method True ACD Estimated ACD Bias Analytic SE Monte Carlo SE MSE Coverage
-Setting 1: $\gamma_{AX}=0$
-Oracle Estimator $0.996$ $0.985$ $-0.011$ $0.034$ $0.035$ $0.001$ $0.920$
-Outcome Modeling and Direct Standardization $0.996$ $0.971$ $-0.026$ $0.038$ $0.035$ $0.002$ $0.885$
-Outcome Modeling and Direct Standardization, Misspecified $0.996$ $0.968$ $-0.028$ $0.034$ $0.030$ $0.002$ $0.900$
-Inverse Probability Weighting 1 $0.996$ $0.983$ $-0.014$ $0.051$ $0.052$ $0.003$ $0.950$
-Inverse Probability Weighting 2 $0.996$ $0.956$ $-0.041$ $0.046$ $0.045$ $0.004$ $0.835$
-Setting 2: $\gamma_{AX}=0.01$
-Oracle Estimator $1.000$ $0.991$ $-0.009$ $0.034$ $0.033$ $0.001$ $0.960$
-Outcome Modeling and Direct Standardization $1.000$ $0.976$ $-0.024$ $0.038$ $0.032$ $0.002$ $0.940$
-Outcome Modeling and Direct Standardization, Misspecified $1.000$ $0.983$ $-0.017$ $0.034$ $0.028$ $0.001$ $0.970$
-Inverse Probability Weighting 1 $1.000$ $0.992$ $-0.008$ $0.050$ $0.052$ $0.003$ $0.935$
-Inverse Probability Weighting 2 $1.000$ $0.955$ $-0.045$ $0.047$ $0.050$ $0.004$ $0.825$
-Setting 3: $\gamma_{AX}=0.05$
-Oracle Estimator $1.050$ $1.051$ $0.001$ $0.034$ $0.030$ $0.001$ $0.970$
-Outcome Modeling and Direct Standardization $1.050$ $1.035$ $-0.015$ $0.038$ $0.029$ $0.001$ $0.965$
-Outcome Modeling and Direct Standardization, Misspecified $1.050$ $1.059$ $0.008$ $0.035$ $0.026$ $0.001$ $0.985$
-Inverse Probability Weighting 1 $1.050$ $1.041$ $-0.009$ $0.051$ $0.048$ $0.002$ $0.955$
-Inverse Probability Weighting 2 $1.050$ $1.012$ $-0.037$ $0.046$ $0.043$ $0.003$ $0.870$
-Setting 4: $\gamma_{AX}=0.1$
-Oracle Estimator $1.106$ $1.098$ $-0.008$ $0.034$ $0.034$ $0.001$ $0.950$
-Outcome Modeling and Direct Standardization $1.106$ $1.082$ $-0.022$ $0.038$ $0.033$ $0.002$ $0.925$
-Outcome Modeling and Direct Standardization, Misspecified $1.106$ $1.130$ $0.021$ $0.036$ $0.030$ $0.001$ $0.930$
-Inverse Probability Weighting 1 $1.106$ $1.102$ $-0.003$ $0.050$ $0.047$ $0.002$ $0.965$
-Inverse Probability Weighting 2 $1.106$ $1.070$ $-0.033$ $0.047$ $0.044$ $0.003$ $0.880$
-Setting 5: $\gamma_{AX}=0.5$
-Oracle Estimator $1.501$ $1.489$ $-0.008$ $0.035$ $0.031$ $0.001$ $0.960$
-Outcome Modeling and Direct Standardization $1.501$ $1.470$ $-0.020$ $0.037$ $0.034$ $0.002$ $0.895$
-Outcome Modeling and Direct Standardization, Misspecified $1.501$ $1.720$ $0.146$ $0.041$ $0.029$ $0.049$ $0.000$
-Inverse Probability Weighting 1 $1.501$ $1.497$ $-0.002$ $0.050$ $0.053$ $0.003$ $0.925$
-Inverse Probability Weighting 2 $1.501$ $1.467$ $-0.023$ $0.048$ $0.047$ $0.003$ $0.900$
-Appendix C Additional Data Example Results
-C.1 Main Analysis of Race, SES, and Telomere Length
 Table C2: Results from our (a) unweighted and (b) survey-weighted ‘propensity for race’ models.
@@
    High School or GED
-   Some College 0.75 0.64, 0.89 $<$0.001 0.80 0.67, 0.96 0.016
-   College Graduate 0.49 0.39, 0.61 $<$0.001 0.49 0.38, 0.62 $<$0.001
-   Refused/Unknown 0.96 0.12, 5.42 $>$0.9 0.29 0.02, 3.60 0.3
+   Some College 0.75 0.64, 0.89 <<0.001 0.80 0.67, 0.96 0.016
+   College Graduate 0.49 0.39, 0.61 <<0.001 0.49 0.38, 0.62 <<0.001
+   Refused/Unknown 0.96 0.12, 5.42 >>0.9 0.29 0.02, 3.60 0.3
 Marital Status
@@
    Widowed/Divorced/Separated 0.93 0.75, 1.16 0.5 0.88 0.69, 1.12 0.3
-   Married/Living with Partner 0.43 0.35, 0.52 $<$0.001 0.41 0.33, 0.50 $<$0.001
+   Married/Living with Partner 0.43 0.35, 0.52 <<0.001 0.41 0.33, 0.50 <<0.001
    Refused/Unknown 0.70 0.49, 1.01 0.057 0.82 0.54, 1.23 0.3
@@
    2 People 1.41 1.11, 1.80 0.006 1.33 1.00, 1.77 0.049
-   3 People 2.55 1.95, 3.33 $<$0.001 2.27 1.66, 3.10 $<$0.001
-   4 People 2.51 1.88, 3.36 $<$0.001 2.18 1.57, 3.03 $<$0.001
-   5+ People 3.90 2.87, 5.29 $<$0.001 3.51 2.49, 4.94 $<$0.001
+   3 People 2.55 1.95, 3.33 <<0.001 2.27 1.66, 3.10 <<0.001
+   4 People 2.51 1.88, 3.36 <<0.001 2.18 1.57, 3.03 <<0.001
+   5+ People 3.90 2.87, 5.29 <<0.001 3.51 2.49, 4.94 <<0.001
 Household Income
@@
 Poverty-Income Ratio Category
-   $<$ 1.3
-   1.3 $\leq$ PIR $<$ 3.5 0.96 0.76, 1.21 0.7 0.99 0.75, 1.30 $>$0.9
+   << 1.3
+   1.3 $\leq$ PIR << 3.5 0.96 0.76, 1.21 0.7 0.99 0.75, 1.30 >>0.9
    $\geq$ 3.5 0.90 0.64, 1.26 0.5 0.83 0.56, 1.24 0.4
@@
    Not Working 0.84 0.62, 1.13 0.3 0.90 0.64, 1.28 0.6
-   Refused/Unknown 0.68 0.45, 1.04 0.070 1.01 0.63, 1.62 $>$0.9
+   Refused/Unknown 0.68 0.45, 1.04 0.070 1.01 0.63, 1.62 >>0.9
 Occupation Category
@@
    Low White Collar 0.87 0.44, 1.76 0.7 1.29 0.61, 2.74 0.5
-   High White Collar 0.73 0.36, 1.49 0.4 1.04 0.48, 2.26 $>$0.9
+   High White Collar 0.73 0.36, 1.49 0.4 1.04 0.48, 2.26 >>0.9
    Refused/Unknown 1.12 0.53, 2.40 0.8 1.26 0.55, 2.86 0.6
@@
    Food Secure
-   Marginally Food Secure 1.80 1.36, 2.38 $<$0.001 2.18 1.59, 2.99 $<$0.001
+   Marginally Food Secure 1.80 1.36, 2.38 <<0.001 2.18 1.59, 2.99 <<0.001
    Food Insecure 1.40 1.10, 1.79 0.006 1.46 1.12, 1.92 0.006
@@
    Family Home Attached 1.46 1.14, 1.86 0.002 1.55 1.19, 2.01 0.001
-   Apartment 2.23 1.78, 2.81 $<$0.001 2.35 1.82, 3.04 $<$0.001
+   Apartment 2.23 1.78, 2.81 <<0.001 2.35 1.82, 3.04 <<0.001
    Other 0.68 0.53, 0.88 0.004 0.85 0.65, 1.11 0.2
@@
 1OR = Odds Ratio, CI = Confidence Interval
+Figure C1: Absolute standardized mean differences and Kolmogorov-Smirnov statistics for the distribution of socioeconomic status factors before and after adjustment by an unweighted propensity score constructed via a generalized linear model on self-reported race.
+Dark blue circles represent pre-adjustment statistics, while light blue triangles represent post-adjustment differences.
+The solid vertical line represents a reference value of 0.0, while the dotted vertical line represents a chosen threshold value of 0.1.
+Figure C2: Absolute standardized mean differences and Kolmogorov-Smirnov statistics for the distribution of socioeconomic status factors before and after adjustment by a survey-weighted propensity score constructed via a generalized linear model on self-reported race.
+Dark blue circles represent pre-adjustment statistics, while light blue triangles represent post-adjustment differences.
+The solid vertical line represents a reference value of 0.0, while the dotted vertical line represents a chosen threshold value of 0.1.
 C.2 Supplemental Analysis of Lead Exposure, SES, and Telomere Length
-Table C2: Population Average Treatment Effect (PATE) estimates and corresponding 95% confidence intervals across various analytic approaches for the comparison of effect of lead exposure (high versus low exposure, dichotomized at median $\mu$g/dL) on log-transformed telomere length among $n=5,267$ participants in the National Health and Nutrition Examination Survey (NHANES).
-IPTW: Inverse Probability of Treatment Weighting.
-Method PATE Estimate 95% Confidence Interval
-Multiple Regression $0.0002$ $-0.0139,0.0142$
-IPTW Estimator $-0.0478$ $-0.0800,-0.0156$
-IPTW Estimator w/ Race in Propensity Model $-0.0504$ $-0.0809,-0.0199$
-Survey-Weighted Multiple Regression $-0.0032$ $-0.0229,0.0165$
-IPTW Multiple Regression $-0.0024$ $-0.0157,0.0109$
-IPTW Multiple Regression w/ Race in Propensity Model $-0.0029$ $-0.0161,0.0103$
-IPTW + Survey-Weighted Multiple Regression $-0.0051$ $-0.0250,0.0149$
-IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0053$ $-0.0255,0.0149$
-Weighted IPTW + Survey-Weighted Multiple Regression $-0.0049$ $-0.0253,0.0155$
-Weighted IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model $-0.0052$ $-0.0257,0.0153$
-Proposed OM $0.0001$ $-0.0121,0.0123$
-Proposed OM w/ Race in Propensity Model $0.0001$ $-0.0121,0.0123$
-Proposed IPW1 $-0.0358$ $-0.0469,-0.0248$
-Proposed IPW1 w/ Race in Propensity Model $-0.0368$ $-0.0479,-0.0257$
-Proposed IPW2 $-0.0328$ $-0.0433,-0.0223$
-Proposed IPW2 w/ Race in Propensity Model $-0.0332$ $-0.0440,-0.0224$
+In a supplementary analysis, we present a different example from NHANES which may be amenable to a causal interpretation (in the sense that counterfactual/potential outcomes and plausibly changing exposures are more meaningful).
+To illustrate how the proposed methods can, under stronger assumptions, identify the average difference in potential outcome means (population average treatment effect; PATE), we study the effect of exposure to a particular environmental contaminant, lead, on telomere length (Zota et al., 2015).
+Lead exposure is associated with negative health outcomes, such as cardiovascular and kidney diseases (Navas-Acien et al., 2007, 2009), and biological markers of oxidative stress and inflammation (Lin et al., 2016).
+Recent works suggest that lead exposure may be associated with telomere shortening (Lin et al., 2016; Wu et al., 2012; Zota et al., 2015), however, evidence from these studies is mixed and depends on the analytic strategy.
+Thus, further exploring the estimated effect of lead exposure on telomere length with our proposed approaches may be of interest.
+[…]
+We repeated our the setup of our main analysis, treating the dichotomized lead variable as our exposure $(A)$.
+As before, we fit a propensity score for lead exposure conditional on the twelve socioeconomic indicators $(\boldsymbol{X})$ described in the main text.
+We further included self-reported race using two strategies – first as a precision covariate along with age and blood composition (i.e., not in $\boldsymbol{X}$) and second, in a sensitivity analysis, as a part of our confounders, $\boldsymbol{X}$.
+The results of this analysis are presented in Supplemental Table C3.
+While the point estimates and findings differed across methods, we found a significant exposure effect for both the traditional and proposed IPW estimators, with an estimated 3.3-5.0% reduction in telomere length.
+For the remaining methods, including the proposed outcome modeling and AIPW estimators, we failed to reject the null hypothesis.
+Further, there were no appreciable differences in these estimates or the corresponding conclusions regarding significance when comparing our results with versus without including self-reported race in the propensity model (where applicable).
+[…]
+Method PATE 95% CI
+Multiple Regression 0.0002 -0.0139 0.0142
+IPTW Estimator -0.0478 -0.0800 -0.0156
+IPTW Estimator w/ Race in Propensity Model -0.0504 -0.0809 -0.0199
+Survey-Weighted Multiple Regression -0.0032 -0.0229 0.0165
+IPTW Multiple Regression -0.0024 -0.0157 0.0109
+IPTW Multiple Regression w/ Race in Propensity Model -0.0029 -0.0161 0.0103
+IPTW + Survey-Weighted Multiple Regression -0.0051 -0.0250 0.0149
+IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model -0.0053 -0.0255 0.0149
+Weighted IPTW + Survey-Weighted Multiple Regression -0.0049 -0.0253 0.0155
+Weighted IPTW + Survey-Weighted Multiple Regression w/ Race in Propensity Model -0.0052 -0.0257 0.0153
+Outcome Modeling and Direct Standardization 0.0001 -0.0121 0.0123
+Outcome Modeling and Direct Standardization w/ Race in Propensity Model 0.0001 -0.0121 0.0123
+Inverse Probability Weighting 1 -0.0358 -0.0469 -0.0248
+Inverse Probability Weighting 1 w/ Race in Propensity Model -0.0368 -0.0479 -0.0257
+Inverse Probability Weighting 2 -0.0328 -0.0433 -0.0223
+Inverse Probability Weighting 2 w/ Race in Propensity Model -0.0332 -0.0440 -0.0224
+Augmented Inverse Probability Weighting -0.0007 -0.0121 0.0106
+Augmented Inverse Probability Weighting w/ Race in Propensity Model -0.0009 -0.0124 0.0107